### v1

# Bayesian Synthetic Control with a Soft Simplex Constraint

Abstract
Whether the synthetic control method should be implemented with the simplex constraint and how to implement it in a high-dimensional setting have been widely discussed. To address both issues simultaneously, we propose a novel Bayesian synthetic control method that integrates a soft simplex constraint with spike-and-slab variable selection. Our model is featured by a hierarchical prior capturing how well the data aligns with the simplex assumption, which enables our method to efficiently adapt to the structure and information contained in the data by utilizing the constraint in a more flexible and data-driven manner. A unique computational challenge posed by our model is that conventional Markov chain Monte Carlo sampling algorithms for Bayesian variable selection are no longer applicable, since the soft simplex constraint results in an intractable marginal likelihood. To tackle this challenge, we propose to update the regression coefficients of two predictors simultaneously from their full conditional posterior distribution, which has an explicit but highly complicated characterization. This novel Gibbs updating scheme leads to an efficient Metropolis-within-Gibbs sampler that enables effective posterior sampling from our model and accurate estimation of the average treatment effect. Simulation studies demonstrate that our method performs well across a wide range of settings, in terms of both variable selection and treatment effect estimation, even when the true data-generating process does not adhere to the simplex constraint. Finally, application of our method to two empirical examples in the economic literature yields interesting insights into the impact of economic policies.
Keywords: average treatment effect; constrained linear regression; Gibbs sampling; Metropolis-within-Gibbs sampler; spike-and-slab prior.
1 Introduction
1.1 Background on Synthetic Control Methods
In many fields of social science, researchers need to identify the causal effect of an intervention or treatment of interest using only observational data, since conducting case-control experiments is either infeasible or improper. This is a common situation in economics, where interventions are implemented at an aggregate level affecting one single large unit (e.g., a country) on some aggregate outcome of interest (e.g., GDP). A classical example, which was studied in the seminal work of Abadie and Gardeazabal (2003), is how terrorist conflicts between the 1970s and 1990s in the Basque Country affected the local economy, such as the per capita GDP. Estimating how the economy would have evolved in the absence of political conflicts poses several challenges. On one hand, one should not simply compare the Basque Country with its past, as it would be confounded by other factors happening at the same time of the terrorism, such as the broader economic downturn experienced by Spain. On the other hand, the Basque Country differed from other Spanish regions in several key characteristics, and thus a straightforward comparison between the Basque Country and the rest of Spain would fail to separate the impact of terrorism from other variables affecting economic growth. To estimate the terrorism impact, Abadie and Gardeazabal (2003) proposed constructing a weighted average using the other regions in Spain to reflect the relevant economic characteristics of the Basque Country, and then this synthetic counterfactual was compared to the real GDP data of the Basque Country to assess the economic effect of terrorism. This approach is called the synthetic control method (SCM). Since its introduction, SCM has gained widespread adoption in empirical research due to its interpretability, transparency, and plausible identification assumptions. Athey and Imbens (2017) described SCM as “arguably the most important innovation in the policy evaluation literature in the last 15 years.”
SCM can be formulated as follows. Consider a panel with $T=T_{0}+T_{1}$ total observations over time for $N+1$ units, where only the first unit receives the treatment after period $T_{0}$. The remaining $N$ units form the control group, which are unaffected by the intervention. Let $Z_{i,t}$ denote the outcome of interest for unit $i$ at time $t$, where $i=1,2,...,N+1$ and $t=1,2,...,T$. The objective of SCM is to estimate the average treatment effect on the treated (ATT) by estimating the counterfactual outcome $Z_{1,t}^{(0)}$ for the treated unit at times $t=T_{0}+1,\dots,T$, which represents what the outcome of unit 1 would have been in the absence of the treatment. The standard approach is to define $Z_{1,t}^{(0)}$ as a weighted combination of the outcomes of the $N$ untreated units:
$Z_{1,t}^{(0)}(w)=\sum_{i=2}^{N+1}w_{i-1}Z_{i,t},$ (1)
where the weights $(w_{i})_{i=1}^{N}$ are time-independent, non-negative and sum to one. A larger $w_{i}$ indicates that the $i$-th control unit better approximates the treated unit’s pre-intervention characteristics. To find the optimal set of weights $(w_{i})_{i=1}^{N}$ such that the synthetic control approximates the pre-treatment outcome of the treated unit as closely as possible, a common strategy is to minimize the mean squared error between the treated unit and the synthetic control in the pre-treatment period $t=1,\dots,T_{0}$:
$\hat{w}=\operatorname*{argmin}_{w\in\Delta^{N-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\left(Z_{1,t}-\sum_{i=2}^{N+1}w_{i-1}Z_{i,t}\right)^{2}$ (2)
where
$\Delta^{N-1}=\left\{u\in\mathbb{R}^{N}\colon u_{i}\geq 0\text{ for each }i,\text{ and }\sum\nolimits_{i=1}^{N}u_{i}=1\right\}$ (3)
denotes the $(N-1)$-simplex. Note that for ease of presentation, we have not considered time-invariant characteristics in the above formulation, and readers are referred to Abadie (2021) for a full description of SCM. The average difference between the observed outcome $Z_{1,t}$ and the synthetic control $Z_{1,t}^{(0)}(\hat{w})$ serves as an estimate of the ATT:
$\widehat{\mathrm{ATT}}=\frac{1}{T_{1}}\sum_{t=T_{0}+1}^{T}\left(Z_{1,t}-\sum_{i=2}^{N+1}\hat{w}_{i-1}Z_{i,t}\right),$ (4)
where $T_{1}=T-T_{0}$. A large portion of the literature on SCM methodology focuses on how to move beyond the simplex constraint and extend SCM to cases where $N$ can be very large, possibly greater than $T_{0}$, and below we provide a brief review.
Some researchers advocated for using regression-based methods to estimate $w$ without making the simplex assumption (Doudchenko and Imbens, 2016), since they argued that the simplex constraint is not likely to hold in practice and negative weights may also be desirable in some scenarios. One influential work in this vein was Hsiao et al. (2012), who assumed the cross-sectional dependence is driven by a common factor model and estimated the counterfactual via a regression model without identifying the unobserved factors. They computed the weight $w$ in (1) as an OLS-type estimator (OLS stands for Ordinary Least Squares) without the simplex constraint, and thus the synthetic control $Z_{1,t}^{(0)}$ was no longer a convex combination of $(Z_{i,t})_{i=2}^{N+1}$. By applying the synthetic control to post-period data, they obtained counterfactual and ATT estimates for the treated unit. Xu (2017) further generalized the method of Hsiao et al. (2012) by estimating the unknown factors, and the resulting method could be applied to multiple treated units or data sets with large $N$. Note that since dimension reduction is inherent in factor model, their method does not require selecting control units when $N$ is large. Meanwhile, Abadie (2021) noted that such regression-based methods may increase the chance of extrapolation, leading to inconsistent ATT estimation, and they used a geometric argument to demonstrate that the simplex constraint has the desirable sparisty-inducing property; a formal proof was given in Goh and Yu (2022). Another theoretical justification for the simplex constraint was derived in the recent work of Martinez and Vives-i-Bastida (2022). Aiming to combine the strengths of both approaches, Li (2020) proposed to relax the simplex constraint by only assuming the weights are non-negative (but may not sum to one). Another point of contention is about the inclusion of the intercept term. Some researchers recommend incorporating the intercept term to allow systematic difference between the treated and control unit (Doudchenko and Imbens, 2016, Goh and Yu, 2022).
Application of SCM to high-dimensional data has also gained increasing attention, since in fields such as economics and political science, the number of observations can often be small due to the low frequency of data collection (often on a quarterly or even yearly basis). In certain applications, the number of pre-period observations, $T_{0}$, can be close to or even smaller than the number of control units (e.g., states, countries). When $N>T_{0}$, the optimization problem (2) is often ill-defined. In particular, OLS estimators cannot be computed since they require inverting a singular empirical covariance matrix. Similarly, though the simplex constraint significantly reduces the effective dimension of the parameter space, perfect fitting (i.e., the mean squared loss in (2) equal to zero) often occurs when $N$ is much larger than $T_{0}$. Consequently, methods proposed by Abadie and Gardeazabal (2003) and Hsiao et al. (2012) are considered infeasible when $N\gg T_{0}$, and variable selection techniques were introduced into SCM which assume that only a small number of control units should be used to construct the synthetic control. Specifically, variable selection methods that have been used in the SCM literature include elastic net (Doudchenko and Imbens, 2016), Lasso (Carvalho et al., 2018, Hollingsworth and Wing, 2020), and forward selection (Shi and Huang, 2023), among others.
Though not the focus of the present work, we note that there is a substantial body of literature on the asymptotic theory and inference for frequentist synthetic control methods. See, for example, Abadie et al. (2010) and Firpo and Possebom (2018) for the use of permutation methods (known as “placebo tests”) in ATT estimation, and Carvalho et al. (2018), Li (2020), Chernozhukov et al. (2021) for the inference with SCM under high-dimensional settings.
In this work, we adopt a Bayesian approach to constructing the synthetic control, a strategy that has become popular in the recent SCM literature for its flexibility in integrating with conventional SCM, regression-based methods and high-dimensional techniques. Another important advantage of the Bayesian approach is that it provides posterior distributions for both ATT and the counterfactual outcome of the treated unit, which are more informative than point estimates. By computing posterior mean estimates, Bayesian methods make statistical inference by averaging over the posterior distributions, a feature known as model averaging (Kass and Raftery, 1995), which contrasts with penalized regression methods like Lasso that focus on identifying a single best model to minimize the loss function. A potential limitation, however, is that the posterior fitting often requires the use of Markov chain Monte Carlo (MCMC) sampling, which can be very time-consuming compared to optimization-based methods. One recent work by Goh and Yu (2022) considered Bayesian linear regression where the prior distribution of the regression coefficients is uniform over a convex hull, corresponding to the simplex constraint proposed by Abadie and Gardeazabal (2003). They also discussed the Bayesian formulation of the method of Li (2020), which involves using a prior over the shifted conical hull. A similar model was studied in Martinez and Vives-i-Bastida (2022), and a Bernstein-von Mises result for the Bayes estimator was proved. Kim et al. (2020) proposed using either the horseshoe prior or the spike-and-slab prior, two well-known Bayesian approaches to variable selection, to construct synthetic control with selected control units in high-dimensional settings; the simplex constraint was dropped in their method. It was demonstrated in their paper that their method outperforms those frequentist SCM methods based on variable selection. Nevertheless, to our knowledge, there is no existing Bayesian variable selection method for constructing synthetic control that also incorporates the simplex constraint.
1.2 Overview of this Work
We take a neutral stance regarding the use of the simplex constraint. On one hand, this constraint makes the model easier to interpret and has demonstrated success in various real data sets. On the other hand, we agree that it can be overly restrictive in many cases. Hence, we propose a balanced approach based on a novel Bayesian model called BVS-SS (Bayesian Variable Selection with Soft Simplex constraint), which allows the data to determine whether the simplex constraint is appropriate on a case-by-case basis. BVS-SS aims to leverage the strengths of both simplex-constrained SCM and unconstrained regression-based methods, which we believe serves as an effective solution to the debate surrounding the use of this constraint. There is no similar method in the frequentist SCM literature to our knowledge.
BVS-SS utilizes a standard spike-and-slab-type prior for selecting which control units should be used for constructing the synthetic control, a technique that has gained significant success and widespread popularity in the Bayesian community (George and McCulloch, 1993, 1997, Brown et al., 1998). Different from existing models, BVS-SS assumes that there is an underlying sparse “mean weight vector” $\mu$ that satisfies the simplex constraint, but the actual weights $w_{1},\dots,w_{N}$ involved in (1) may deviate from $\mu$. The key innovation of BVS-SS lies in the introduction of a variance parameter $\tau$ quantifying the amount of this deviation. By imposing a non-informative prior on $\tau$, we let BVS-SS learn its value from the data, revealing whether the data actually agrees with the simplex assumption. To generate samples from the posterior distribution of BVS-SS, we propose an original Metropolis-within-Gibbs sampler, where $\mu$ and error variance are updated by Gibbs schemes and $\tau$ is updated by Metropolis-Hastings steps. Our Gibbs-type updating for $\mu$ is different from existing Gibbs schemes for similar problems, as those methods typically only update one coordinate at a time, which is not feasible in our context due to the simplex constraint (i.e., given $\mu_{-j}$, we have $\mu_{j}=1-\sum_{i\neq j}\mu_{i}$ almost surely). We overcome this difficulty by theoretically finding the joint conditional posterior distribution of two coordinates $(\mu_{i},\mu_{j})$ given all the other parameters, which enables us to update two coordinates simultaneously. This full conditional posterior distribution is highly complicated due to the spike-and-slab prior we use for variable selection, but fortunately sampling from this distribution can be performed straightforwardly and efficiently. The ATT estimation for SCM contains two stages. First, we compute the posterior distribution of the weights $(w_{i})_{i=1}^{N}$ using the pre-treatment data. Next, we estimate ATT by constructing a synthetic control counterfactual estimate using the post-treatment data. Since we have samples of weights generated from the proposed MCMC sampler, we can compute both the posterior mean estimate of ATT and its credible interval without extra computational cost.
It should be noted that BVS-SS is intrinsically different from other hierarchical spike-and-slab variable selection models, such as those used by Guan and Stephens (2011), Zhou and Guan (2019) for heritability estimation in biology, since the marginal likelihood for given model and variance parameters in BVS-SS does not have a closed-form expression due to the simplex constraint (i.e., the vector $\mu$ cannot be integrated out). This is why we cannot apply the more common approach in the Bayesian variable selection literature, where one directly updates an indicator vector representing the regression model by add/delete/swap proposals (see George and McCulloch (1997), Guan and Stephens (2011), Chang and Zhou (2024), among many others). While updating the regression coefficients instead of the indicator vector may seem less efficient, this offers an elegant solution to the posterior computation of BVS-SS and yields favorable results in all of our numerical experiments. Besides, the simplex constraint encoded by BVS-SS also differs fundamentally from other constraints considered in the variable selection literature, such as the grouping, hierarchical and anti-hierarchical constraints studied by Choi et al. (2010), Farcomeni (2010) and the network constraint utilized in Li and Li (2008).
In addition to simulated data sets, we apply our method to two empirical examples in the frequentist SCM literature. We first revisit the data set considered in Carvalho et al. (2018) to study how an anti-tax evasion program would affect inflation in Brazil. The control group consists of eight large metropolitan areas. While the method of Carvalho et al. (2018) selected the whole control group, BVS-SS selects a significantly smaller number of control units to construct the synthetic control. Then we revisit the data set studied in Shi and Huang (2023) to estimate the impact of anti-corruption policy on luxury watch imports in China. The control group contains 87 commodity categories and is larger than the number of observed treatment periods, making conventional synthetic control methods inapplicable. Our ATT estimation provides strong evidence supporting that China’s anti-corruption policy is effective, which is consistent with the finding of Shi and Huang (2023), but our estimated effect is slightly smaller than that of Shi and Huang (2023).
The paper proceeds as follows. Section 2 introduces the BVS-SS model and outlines how to perform ATT estimation. Section 3 details our MCMC sampler for posterior computation with BVS-SS. Section 4 presents simulation studies which illustrate the advantages of BVS-SS compared to other competing methods. Section 5 discusses ATT estimation via BVS-SS for the two empirical examples. Section 6 concludes the paper with some discussion.
2 Bayesian Variable Selection with Soft Simplex Constraint
2.1 Model and Prior
For generality and ease of notation, henceforth we use $Y\in\mathbb{R}^{M}$ to denote a response vector of $M$ observations and $X\in\mathbb{R}^{M\times N}$ to denote an arbitrary design matrix with $N$ explanatory variables. Assume that
$Y=Xw+\epsilon,\quad\epsilon\sim N(0,\phi^{-1}I),$ (5)
where $w\in\mathbb{R}^{N}$ is the vector of unknown regression coefficients, $\epsilon$ is the vector of i.i.d. Gaussian noise with variance $\phi^{-1}$, and $I$ denotes the identity matrix. For the synthetic control problem described in Section 1.1, we have $M=T_{0}$, $Y_{t}=Z_{1,t}$ and $X_{tj}=Z_{j+1,t}$ for $t=1,\dots,T_{0}$ and $j=1,\dots,N$.
We propose a hierarchical prior distribution on $w$ which induces sparsity and incorporates the simplex constraint. To describe it, introduce an indicator vector $\gamma\in\{0,1\}^{N}$ such that $\gamma_{i}=1$ if and only if $w_{i}\neq 0$. Let $|\gamma|=\sum_{j=1}^{N}\gamma_{j}$ denote the number of selected predictors. Given a vector $w$, we use $w_{\gamma}$ to denote the subvector of $w$ with entries indexed by $\{i\colon\gamma_{i}=1\}$. We consider the following prior on $(w,\phi)$,
ϕsimilar-toitalic-$\phi\sim\;$ $\mathrm{Gamma}(\kappa_{1}/2,\kappa_{2}/2),$ (6)
$\gamma_{i}\overset{\mathrm{i.i.d.}}{\sim}\;$ $\mathrm{Bernoulli}(\theta),$ (7)
$\tau\sim\;$ $\mathrm{Gamma}(a_{1},a_{2}),$ (8)
[…]
$\mu_{i}\mid\gamma_{i}=0\sim\;$ $\delta_{0},$ (10)
$w_{\gamma}\mid\gamma,\mu_{\gamma},\tau,\phi\sim\;$ $N(\mu_{\gamma},\,(\tau/\phi)I),$ (11)
$w_{i}\mid\gamma_{i}=0\sim\;$ $\delta_{0},$ (12)
[…]
In plain words, we assume that given $\gamma$ and $\phi$, $w_{\gamma}$ follows a normal distribution with mean vector $\mu_{\gamma}\in\Delta^{|\gamma|-1}$ and prior variance depending on the parameter $\tau$. This conditional prior for $w$ can be viewed as an interpolation between the simplex constraint and the unconstrained setting. If the data suggests that $w$ is likely to satisfy the simplex constraint, then the posterior distribution of $\tau$ should concentrate around zero, while if the simplex constraint is significantly violated, $\tau$ should stay away from zero in the posterior. As will be demonstrated in our simulation study, $\tau$ can be reasonably estimated even with a relatively small sample size, which tells the user if the simplex constraint is appropriate for the data set being analyzed. This approach effectively reconciles the different views on the use of the simplex constraint, and thus we call our model BVS-SS (Bayesian Variable Selection with Soft Simplex constraint). We note that our model is similar to the one studied in Martinez and Vives-i-Bastida (2022); however, they did not consider variable selection, and in their implementation they assumed that $w$ satisfies the hard simplex constraint (though the soft constraint was used in their theory). Regarding the hyperparameters, $a_{1},a_{2}$ should be chosen to reflect one’s prior belief on whether simplex constraint is likely to hold, and a larger value of $\alpha$ encourages a more even distribution of the weights $\{\mu_{i}\colon\gamma_{i}=1\}$. In practice, one often has limited information about the true weight vector $w$, but the effect of these hyperparameters quickly becomes negligible as sample size increases.
The other elements of our model are standard. The prior for $\phi$ is conjugate, and one can choose small values for $\kappa_{1},\kappa_{2}$ as a noninformative prior. The hyperparameter $\theta$ is the marginal prior probability of an explanatory variable being included in the regression model, and a smaller value of $\theta$ represents a heavier penalty on the model complexity.
2.2 Posterior Distribution and ATT Estimation
Let $X_{\gamma}$ denote the submatrix of $X$ with columns indexed by $\{i\colon\gamma_{i}=1\}$. Under the prior distribution of BVS-SS, we have
$Y\mid\gamma,\mu_{\gamma},\tau,\phi\sim N\left(X_{\gamma}\mu_{\gamma},\,\phi^{-1}(\tau X_{\gamma}X_{\gamma}^{\top}+I)\right).$ (13)
Since $\gamma$ can be viewed as a function of $\mu$ with $\gamma_{i}=\mathbbm{1}_{\{\mu_{i}\neq 0\}}$, conditioning on $(\gamma,\mu_{\gamma},\tau,\phi)$ is equivalent to conditioning on $(\mu,\tau,\phi)$. So henceforth we will denote the conditional distribution in (13) by $p(y\mid\mu,\tau,\phi)$, and $\gamma$ should be understood as $\gamma=\gamma(\mu)$. A routine calculation using Woodbury identity yields the marginal likelihood of $(\mu,\tau,\phi)$:
$p(y\mid\mu,\tau,\phi)\propto\phi^{M/2}\tau^{-|\gamma|/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\exp\left\{-\frac{\phi}{2}(y-X_{\gamma}\mu_{\gamma})^{\top}\Sigma_{\gamma,\tau}(y-X_{\gamma}\mu_{\gamma})\right\},$ (14)
where
$V_{\gamma,\tau}=X_{\gamma}^{\top}X_{\gamma}+\tau^{-1}I,\quad\Sigma_{\gamma,\tau}=I-X_{\gamma}V_{\gamma,\tau}^{-1}X_{\gamma}^{\top}.$ (15)
Further, using the conjugacy of normal-gamma prior, we find that the conditional posterior distributions of $w_{\gamma}$ and $\phi$ are given by
$\phi\mid y,\mu,\tau\sim\;$ $\mathrm{Gamma}\left(\frac{M+\kappa_{1}}{2},\,\frac{\kappa_{2}+(y-X_{\gamma}\mu_{\gamma})^{\top}\Sigma_{\gamma,\tau}(y-X_{\gamma}\mu_{\gamma})}{2}\right),$ (16)
$w\mid y,\mu,\tau,\phi\sim\;$ $N\left(V_{\gamma,\tau}^{-1}(X^{\top}_{\gamma}y+\tau^{-1}\mu_{\gamma}),\,\phi^{-1}V_{\gamma,\tau}^{-1}\right).$ (17)
The joint posterior distribution of $(\mu,\tau,\phi)$ can be computed by
$p(\mu,\tau,\phi\mid y)\propto p(y\mid\mu,\tau,\phi)p(\mu_{\gamma}\mid\gamma)p(\gamma)p(\tau)p(\phi)$ (18)
where we have used the prior independence between $\mu,\tau,\phi$. In Section 3 we will propose an efficient MCMC sampling algorithm targeting $p(\mu,\tau,\phi\mid y)$.
Prediction with BVS-SS can be conducted as follows. Let $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ denote the design matrix of another $\tilde{M}$ observations, and suppose we are interested in estimating $\tilde{X}w$. Letting $\mathbb{E}_{y}$ denote the posterior distribution given $Y=y$ and using (17), we obtain that
$\mathbb{E}_{y}[\tilde{X}w]=\mathbb{E}_{y}\left[\mathbb{E}_{y}(\tilde{X}w\mid\mu,\tau,\phi)\right]=\mathbb{E}_{y}\left[\tilde{X}V_{\gamma,\tau}^{-1}(X^{\top}_{\gamma}y+\tau^{-1}\mu_{\gamma})\right].$ (19)
Given a sequence of $n$ MCMC samples $(\mu^{(k)},\tau^{(k)},\phi^{(k)})_{k=1}^{n}$, we can approximate the above expectation using the sample average, which yields the posterior predictive mean of $\tilde{X}w$. The covariance matrix of the posterior predictive distribution of $\tilde{X}w$ can be calculated similarly.
For applications like synthetic control, the goal is to estimate ATT among the $\tilde{M}$ observations. Denote their response under the treatment by $\tilde{Y}^{(1)}$, which is assumed observed, and denote the counterfactual response without treatment by $\tilde{Y}^{(0)}$. Let $\delta=\tilde{Y}^{(1)}-\tilde{Y}^{(0)}$, and define ATT by
$\mathrm{ATT}=\frac{1}{\tilde{M}}\sum_{i=1}^{\tilde{M}}\delta_{i}=\frac{1}{\tilde{M}}\sum_{i=1}^{\tilde{M}}\left(\tilde{Y}^{(1)}_{i}-\tilde{Y}^{(0)}_{i}\right),$ (20)
which is the main parameter of interest. By replacing $\tilde{Y}^{(0)}$ using the synthetic control $\mathbb{E}_{y}[\tilde{X}w]$, we obtain an estimator for $\bar{\delta}$. If we have samples $(w^{(k)})_{k=1}^{n}$ approximating the marginal posterior distribution of $w$, we can also obtain the posterior predictive distribution of ATT by plugging $w^{(k)}$ into the estimator $\tilde{M}^{-1}\sum_{i=1}^{\tilde{M}}(\tilde{Y}^{(1)}_{i}-w^{\top}\tilde{X}_{(i)})$, where $\tilde{X}_{(i)}$ denotes the $i$-th row of $\tilde{X}$ (treated as a column vector).
3 A Metropolis-within-Gibbs Sampler for BVS-SS
For Bayesian variable selection involving a prior on the indicator vector $\gamma$ as we have considered, the standard approach to posterior sampling is to integrate out the regression coefficient vector and use add-delete-swap proposals to construct a Metropolis–Hastings algorithm targeting the marginal posterior distribution of $\gamma$ (or the joint distribution of $\gamma$ and some variance parameter); see, e.g., George and McCulloch (1997), Chipman et al. (2001), Guan and Stephens (2011), Zhou et al. (2022). However, this approach is not applicable to our model, since $\mu_{\gamma}$ cannot be integrated out in closed form from the joint posterior distribution of $(\gamma,\mu_{\gamma},\tau,\phi)$. To tackle this very unique challenge, we propose a novel Metropolis-within-Gibbs sampler targeting the joint posterior distribution of $(\mu,\tau,\phi)$.
3.1 Updating of $\tau$ and $\phi$
The updates for $\tau$ and $\phi$ are standard. We update $\phi$ by drawing it from the full conditional distribution given in (16), and we use a Metropolis–Hastings update for $\tau$ that is invariant with respect to $p(\tau\mid y,\mu,\phi)$. Explicitly, we propose a new value for $\tau^{*}$ by a Gaussian random walk on log-scale with variance $\eta>0$ and accept $\tau^{*}$ with probability
$\rho(\tau,\tau^{*})=\min\left\{1,\frac{p(y\mid\mu,\tau^{*},\phi)p(\tau^{*})}{p(y\mid\mu,\tau,\phi)p(\tau)}\frac{\log(\tau^{*})}{\log(\tau)}\right\},$ (21)
where the likelihood is given in (14) and $p(\tau)$ denotes the prior density. This step can be repeated multiple times so that $\tau$ is updated to some value with larger density (conditioned on $\mu$ and $\phi$).
3.2 Updating of $\mu$
For $\mu$, we propose a Gibbs-type updating scheme, which, to our knowledge, is quite different from any existing Gibbs scheme for similar problems. Let $\mu_{-j}$ denote the subvector of $\mu$ with $\mu_{j}$ removed. Since the prior of BVS-SS assigns probability one to $\mu\in\Delta^{N-1}$, given $\mu_{-j}$, we have $\mu_{j}=1-\sum_{i\neq j}\mu_{i}$ almost surely, and thus the full conditional distribution $p(\mu_{j}\mid y,\mu_{-j},\tau,\phi)$ is also degenerate. Hence, a Gibbs scheme that updates each $\mu_{j}$ from its full conditional posterior will not move at all. This motivates us to devise a more complicated updating scheme by updating $\mu_{i},\mu_{j}$ (with $i\neq j$) simultaneously from the full conditional posterior $p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$, where $\mu_{-(i,j)}$ denotes the subvector of $\mu$ with the $\mu_{i},\mu_{j}$ removed.
Let $\mu_{-(i,j)},\tau,\phi$ be given and define $s=1-\sum_{k\neq i,j}\mu_{k}$. If $s=0$, the simplex constraint implies that $\mu_{i}=\mu_{j}=0$ almost surely. If $s\in(0,1]$, $p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$ is a mixture of a discrete distribution on $\{(s,0),\,(0,s)\}$ and a degenerate continuous distribution on $\{(u,s-u)\colon 0<u<s\}$. Therefore, to find a closed-form expression for this conditional distribution, it suffices to find $p(\gamma_{i},\gamma_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$ for $(\gamma_{i},\gamma_{j})\in\{(1,0),(0,1),(1,1)\}$ and the conditional distribution $p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)$. To this end, we introduce the notation $\gamma^{0},\gamma^{i},\gamma^{j},\gamma^{ij}$ defined by
$\gamma^{0}_{k}=\mathbbm{1}(\mu_{k}\neq 0),\quad\gamma^{i}_{k}=\mathbbm{1}(\mu_{k}\neq 0\text{ or }k=i),\quad\gamma^{ij}_{k}=\mathbbm{1}(\mu_{k}\neq 0\text{ or }k\in\{i,j\}).$ (22)
That is, $\gamma^{i}$ is the model obtained by setting $\mu_{i}=s,\mu_{j}=0$, and $\gamma^{ij}$ is the one obtained by setting $\mu_{i},\mu_{j}>0$. The full conditional distribution of $(\mu_{i},\mu_{j})$ given other parameters and $\gamma_{i}=\gamma_{j}=1$ is characterized in the following lemma. For simplicity, we assume $\alpha=1$ in this section, and in Remark 1 we explain how to extend the results to other integer values of $\alpha$.
Lemma 1.
Assume $\alpha=1$. Fix $i\neq j$ such that $s=1-\sum_{k\neq i,j}\mu_{k}\in(0,1]$. The conditional distribution $p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)$ is degenerate with $\mu_{j}=s-\mu_{i}$ and
$\mu_{i}\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1\sim N_{(0,s)}\left(\beta_{i,j},\;\frac{1}{\phi\Lambda_{i,j}}\right).$ (23)
In the above expression, $N_{(a,b)}$ denotes the univariate truncated normal distribution restricted to the interval $(a,b)$,
$\beta_{i,j}=\;$ $\frac{1}{\Lambda_{i,j}}(X_{i}-X_{j})^{\top}\Sigma_{\gamma^{ij},\tau}\left(y-sX_{j}-\sum\nolimits_{k\neq i,j}\mu_{k}X_{k}\right),$ (24)
$\Lambda_{i,j}=\;$ $(X_{j}-X_{i})^{\top}\Sigma_{\gamma^{ij},\tau}(X_{j}-X_{i}),$ (25)
where $\gamma^{ij}$ is given by (22) and $\Sigma_{\gamma,\tau}$ is given by (15).
Proof.
See Section A in the Appendix. ∎
To find the posterior probabilities of $(\gamma_{i},\gamma_{j})=(1,0),(0,1),(1,1)$ given $y,\mu_{-(i,j)},\tau,\phi$, we first express the full conditional posterior of $(\mu_{i},\mu_{j})$ by
$p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi)=C_{1}\,p(y\mid\mu_{i},\mu_{j},\mu_{-(i,j)},\tau,\phi)p(\mu_{i},\mu_{j},\mu_{-(i,j)}),$ (26)
where $C_{1}$ is a constant that depends on $\mu_{-(i,j)}$ but not $(\mu_{i},\mu_{j})$. Since $\mu_{\gamma}\mid\gamma\sim\mathrm{Dirichlet}(\alpha)$, if $s=1-\sum_{k\neq i,j}\mu_{k}>0$, we have
$p(\mu_{i}=s,\mu_{j}=0,\mu_{-(i,j)})=\;$ $C_{2}\frac{\Gamma((\ell+1)\alpha)}{\Gamma(\alpha)^{\ell+1}}\,s^{\alpha-1}\,\theta^{\ell+1}(1-\theta)^{N-\ell-1},$ (27)
$p(\mu_{i}=u,\mu_{j}=s-u,\mu_{-(i,j)})=\;$ $C_{2}\frac{\Gamma((\ell+2)\alpha)}{\Gamma(\alpha)^{\ell+2}}\,u^{\alpha-1}(s-u)^{\alpha-1}\,\theta^{\ell+2}(1-\theta)^{N-\ell-2},\quad\forall u\in(0,s),$ (28)
where $\ell=\sum_{k\neq i,j}\mathbbm{1}_{\{\mu_{k}\neq 0\}}$ and $C_{2}$ is some constant that depends on $\mu_{-(i,j)}$ but not $(\mu_{i},\mu_{j})$. Integrating $p(y\mid\mu_{i}=u,\mu_{j}=s-u,\mu_{-(i,j)},\tau,\phi)p(\mu_{i}=u,\mu_{j}=s-u,\mu_{-(i,j)})$ with respect to $u\in(0,s)$, we obtain the marginal conditional posterior probabilities of $\gamma^{i},\gamma^{j},\gamma^{ij}$.
Lemma 2.
Define a function $A(\gamma,\tau)$ by
$A(\gamma,\tau)=\frac{\Gamma(|\gamma|\alpha)}{\Gamma(\alpha)^{|\gamma|}}\,\tau^{-|\gamma|/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\,\theta^{|\gamma|}(1-\theta)^{N-|\gamma|}.$ (29)
Fix $i\neq j$ such that $s=1-\sum_{k\neq i,j}\mu_{k}\in(0,1]$. Define
$\check{y}(u)=y-uX_{i}-(s-u)X_{j}-\sum\nolimits_{k\neq i,j}\mu_{k}X_{k}.$ (30)
Let $\gamma^{0},\gamma^{i},\gamma^{j},\gamma^{ij}$ be given by (22). Then,
$p(\gamma^{0}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $0,$ (31)
$p(\gamma^{i}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $C\,s^{\alpha-1}A(\gamma^{i},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(s)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(s)\right\},$ (32)
$p(\gamma^{j}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $C\,s^{\alpha-1}A(\gamma^{j},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(0)^{\top}\Sigma_{\gamma^{j},\tau}\check{y}(0)\right\},$ (33)
$p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $CA(\gamma^{ij},\tau)\int_{0}^{s}u^{\alpha-1}(s-u)^{\alpha-1}\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(u)\right\}\mathrm{d}u,$ (34)
where $C>0$ is a constant. In particular, when $\alpha=1$,
$p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)$ (35)
=\;= $\frac{CA(\gamma^{ij},\tau)\sqrt{2\pi}}{\sqrt{\phi\Lambda_{i,j}}}e^{-\frac{\phi}{2}\left\{\check{y}(0)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(0)-\beta_{i,j}^{2}\Lambda_{i,j}\right\}}\left\{\Phi\left((s-\beta_{i,j})\sqrt{\phi\Lambda_{i,j}}\right)-\Phi\left(-\beta_{i,j}\sqrt{\phi\Lambda_{i,j}}\right)\right\},$ (36)
where $\beta_{i,j},\Lambda_{i,j}$ are as given in Lemma 1 and $\Phi$ denotes the CDF of $N(0,1)$.
Proof.
See Section A in the Appendix. ∎
Remark 1.
For any positive integer $\alpha$, the integral involved in $p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)$ can always be expressed by using the CDF of the standard normal distribution, and thus the conditional posterior probabilities of $\gamma^{i},\gamma^{j},\gamma^{ij}$ can be computed very efficiently. For non-integer-valued $\alpha$, numerical integration is needed. The conditional posterior distribution of $(\mu_{i},\mu_{j})$ given $\gamma^{ij}$ is
$p(\mu_{i}=u,\mu_{j}=s-u\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)\,\propto\,u^{\alpha-1}(s-u)^{\alpha-1}e^{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(u)},$
for $u\in(0,s)$. When $\alpha=1$, $u$ follows the truncated normal distribution, as shown in Lemma 1, and this allows for straightforward sampling of $(\mu_{i},\mu_{j})$. For other integer values of $\alpha$, one can use rejection sampling to generate samples of $u$, where the truncated normal distribution can be used as a reference distribution.
3.3 Algorithm
We can now formally describe our Metropolis-within-Gibbs sampler targeting $p(\mu,\tau,\phi\mid y)$, which is given in Algorithm 1. To emphasize its application to synthetic control, we assume that we have access to another data set $(\tilde{X},\tilde{Y}^{(1)})$, and we take the approach outlined in Section 2.2 to estimating ATT defined in (20).
Input: training data set $(X,Y)$, hyperparameters $\kappa_{1},\kappa_{2},a_{1},a_{2},\alpha,\theta$, algorithm parameters $L\in\mathbb{N},\eta>0$ for updating $\tau$, number of iterations $n$, initial state $(\mu^{(0)},\tau^{(0)},\phi^{(0)})$, data set under treatment $(\tilde{X},\tilde{Y}^{(1)})$.
for $t=1,\dots,n$ do
       Set $\mu\leftarrow\mu^{(t-1)}$ ;
       for $i=1,\dots,N-1$ do
             for $j=i+1,\dots,N$ do
                   Calculate $s\leftarrow 1-\sum_{k\neq i,j}\mu_{k}$ ;
                   if $s=0$ then
                         Set $\mu_{i}\leftarrow 0$, $\mu_{j}\leftarrow 0$;
                  else
                         Calculate $p(\gamma\mid y,\mu_{-(i,j)},\tau,\phi)$ for $\gamma=\gamma^{i},\gamma^{j},\gamma^{ij}$ using Lemma 2;
                         Draw $\gamma\in\{\gamma^{i},\gamma^{j},\gamma^{ij}\}$ with probability $p(\gamma\mid y,\mu_{-(i,j)},\tau,\phi)$;
                         if $\gamma=\gamma^{i}$ then
                               Set $\mu_{i}\leftarrow s$, $\mu_{j}\leftarrow 0$;
                        else if $\gamma=\gamma^{j}$ then
                               Set $\mu_{i}\leftarrow 0$, $\mu_{j}\leftarrow s$;
                        else
                               Draw $u\sim N_{(0,s)}\left(\beta_{i,j},(\phi\Lambda_{i,j})^{-1}\right)$ by Lemma 1;
                               Set $\mu_{i}\leftarrow u$, $\mu_{j}\leftarrow s-u$;
      Set $\mu^{(t)}\leftarrow\mu$;
       Draw $\phi^{(t)}$ from $p(\phi\mid y,\mu^{(k)},\tau^{(k-1)})$ by (16);
       Set $\tau\leftarrow\tau^{(t-1)}$;
       for $l=1,\dots,L$ do
             Propose $\tau^{*}$ by $\log\tau^{*}=\log\tau+N(0,\eta)$;
             Calculate acceptance probability $\rho(\tau,\tau^{*})$ by (21);
             Set $\tau\leftarrow\tau^{*}$ with probability $\rho(\tau,\tau^{*})$;
      Set $\tau^{(t)}\leftarrow\tau$;
       Draw $w^{(t)}$ from the conditional distribution $p(w\mid y,\mu^{t},\tau^{t},\phi^{t})$ given in (17);
       Calculate $\tilde{Y}^{(0,t)}\leftarrow\tilde{X}w^{(t)}$ ;
       Set $\widehat{\mathrm{ATT}}^{(t)}\leftarrow\mathrm{mean}(\tilde{Y}^{(1)}-\tilde{Y}^{(0,t)})$;
Set $\widehat{\mathrm{ATT}}\leftarrow n^{-1}\sum\nolimits_{t=1}^{n}\widehat{\mathrm{ATT}}^{(t)}$;
Output: Samples $(\mu^{(t)},\tau^{(t)},\phi^{(t)},w^{(t)})_{t=1}^{n}$, ATT estimate $\widehat{\mathrm{ATT}}$.
Algorithm 1 Metropolis-within-Gibbs sampling for BVS-SS.
The most computationally intensive part of Algorithm 1 is the evaluation of expressions of the form $V_{\gamma,\tau}^{-1}z$ for a given vector $z$, which may seem highly time-consuming if $|\gamma|$ is large. But this can be implemented efficiently by computing and updating the Cholesky decomposition of $V_{\gamma,\tau}$, a technique commonly employed in Bayesian spike-and-slab variable selection (Smith and Kohn, 1996, George and McCulloch, 1997). More specifically, at the beginning of the $t$-th iteration, we compute the Cholesky decomposition of $V_{\gamma,\tau^{(t-1)}}$, where $\gamma$ is the model selected in the last iteration. Then, each time we draw $(\mu_{i},\mu_{j})$ from its full conditional posterior, we need to change one or two coordinates of $\gamma$, and the resulting Cholesky decomposition of $V_{\gamma,\tau^{(t-1)}}$ can be obtained by standard updating algorithms (Golub and Van Loan, 2013).
Certain variations of Algorithm 1 are also straightforward to implement. For example, one can replace the fixed-scan Gibbs updating of $\mu$ in Algorithm 1 by a random-scan update: in the $t$-th iteration, we uniformly sample a pair $i<j$ such that $\mu^{(t-1)}_{i}+\mu^{(t-1)}_{j}>0$ and update $(\mu^{(t)}_{i},\mu^{(t)}_{j})$ from its conditional posterior distribution (an acceptance-rejection step is needed to correct for the proposal bias), and then draw $\phi^{(t)}$ from the full conditional posterior and update $\tau^{(t)}$ by one Metropolis–Hastings step. In this case, the iterative complex factorization algorithm proposed by Zhou and Guan (2019) can be used to efficiently solve the linear system $V_{\gamma,\tau}^{-1}z$.
Remark 2.
In Algorithm 1, we draw $w^{(t)}$ from the conditional distribution $p(w\mid y,\mu^{t},\tau^{t},\phi^{t})$ so that the samples $(w^{(t)})_{t=1}^{n}$ can be used to approximate the posterior distribution of $w$. However, if the objective is to compute the posterior mean estimate for ATT, one can further reduce the variance of the estimator by directly setting $w^{(t)}$ to its conditional posterior mean as in (19).
4 Simulation Studies
4.1 Simulation Settings
We simulate the observed training data $X\in\mathbb{R}^{M\times N}$ and $Y\in\mathbb{R}^{M}$ by
$Y=\;$ $Xw^{*}+\epsilon,\text{ where }\epsilon_{i}\overset{\mathrm{i.i.d.}}{\sim}N(0,1/\phi^{*}),$ (37)
and $w^{*}\in\mathbb{R}^{N}$ is the true regression coefficient vector such that $w^{*}_{j}=0$ for $j=J+1,\dots,N$ (that is, only the first $J$ predictors have nonzero effects on $Y$). Similarly, we simulate the data under treatment $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ and $\tilde{Y}^{(1)}\in\mathbb{R}^{\tilde{M}}$ by
$\tilde{Y}^{(1)}=\;$ $\tilde{X}w^{*}+\delta+\tilde{\epsilon},\text{ where }\delta_{i}\overset{\mathrm{i.i.d.}}{\sim}N(\delta^{*},1/\nu^{*}),\;\tilde{\epsilon}_{i}\overset{\mathrm{i.i.d.}}{\sim}N(0,1/\phi^{*}).$ (38)
The vector $\delta\in\mathbb{R}^{\tilde{M}}$ represents the treatment effects, and $\delta^{*}\in\mathbb{R}$ denotes the true ATT. In our simulation, we use $M=\tilde{M}\in\{25,50,100,200\}$ and $N\in\{20,50\}$. When $N=20$, we set $J=5$, and when $N=50$, we set $J=10$. Given $J$, we define a vector $\mu^{*}\in\Delta^{N-1}$ by
$\mu^{*}_{j}=\left\{\begin{array}[]{cc}j/S_{J}&\text{ if }1\leq j\leq J,\\ 0,&\text{ if }J+1\leq j\leq N,\end{array}\right.\quad\text{ where }S_{J}=\frac{1}{2}J(J+1).$ (39)
Then, we define $w^{*}$ by $w^{*}=\lambda\mu^{*}$ where $\lambda>0$ is a scaling factor. When $\lambda=1$, the true data-generating model also satisfies the simplex constraint. Note that, assuming $\lambda>0$, we have $\lambda=\|w^{*}\|_{1}$. We use $\lambda\in\{1,2,3\}$. Given $w^{*}$, we set the variance parameters by $v^{*}=\phi^{*}=4/\|w^{*}\|_{2}^{2}$. For each choice of $(M,N,\lambda)$, we generate 100 replicates where entries of $X,\tilde{X}$ are also sampled independently from the standard normal distribution. We fix true ATT $\delta^{*}=0.5$ throughout our simulation studies.
We set the hyperparameters of BVS-SS by $\kappa_{1}=\kappa_{2}=1,a_{1}=0.01,a_{2}=0.1,\alpha=1,\theta=0.2.$ The choice of $(a_{1},a_{2})$ guarantees that $\tau$ has prior mean $0.1$ and prior variance $1$. Further, for numerical stability, we truncate the prior distribution of $\tau$ by assuming that $\tau\geq 10^{-6}$. For every simulated data set, we initialize Algorithm 1 at $\tau^{(0)}=\phi^{(0)}=1$ and $\mu^{(0)}$ sampled from its prior distribution, and we run Algorithm 1 for $1,000$ iterations and discard the first $500$ iterations as burn-in.
For comparison, we have also implemented the following methods for ATT estimation. First, we use the R package glmnet to perform variable selection and ATT estimation via Lasso; the parameter of Lasso is chosen by cross-validation. Second, we compute the OLS estimator for $w^{*}$ using the entire data set $(X,Y)$ without variable selection, denoted by $\hat{w}_{\mathrm{LS}}$. Then ATT is estimated by
$\widehat{\mathrm{ATT}}_{\mathrm{LS}}=\frac{1}{\tilde{M}}\sum\nolimits_{i=1}^{\tilde{M}}\left(\tilde{Y}^{(1)}_{i}-\hat{w}_{\mathrm{LS}}^{\top}\tilde{X}_{(i)}\right),$ (40)
where $\tilde{X}_{(i)}$ is the $i$-th row of $\tilde{X}$ treated as a column vector. We also compute the simplex-constrained estimator $\hat{w}_{\mathrm{QP}}$ (QP stands for “quadratic programming”) defined by
$\hat{w}_{\mathrm{QP}}=\operatorname*{argmin}_{w\in\Delta^{N-1}}\|Y-Xw\|_{2}^{2},$ (41)
which was introduced in Abadie and Gardeazabal (2003) and can be obtained by quadratic programming (Goldfarb and Idnani, 2006, Abadie et al., 2011). The corresponding ATT estimator $\widehat{\mathrm{ATT}}_{\mathrm{QP}}$ is obtained similarly by replacing $\hat{w}_{\mathrm{LS}}$ with $\hat{w}_{\mathrm{QP}}$ in (40). Finally, we consider two oracle estimators that are not implementable in practice. Denote by $\gamma^{*}=\gamma(w^{*})$ the true indicator vector corresponding to $w^{*}$. We compute the OLS and QP estimators using $(X_{\gamma^{*}},Y)$. That is, these two estimators have access to $\gamma^{*}$ (which is not possible in reality) and will be used for comparison in our simulation studies.
4.2 Estimation of Average Treatment Effect
For each method considered, we compute mean squared error of ATT estimation, i.e., the average of $(\widehat{\mathrm{ATT}}-\delta^{*})^{2}$ over 100 replicates; denote it by $\mathrm{MSE}(\widehat{\mathrm{ATT}})$. To measure the efficiency of an ATT estimator, we further compare its MSE with two reference estimators:
$\widehat{\mathrm{ATT}}^{*}=\tilde{M}^{-1}\sum\nolimits_{i=1}^{\tilde{M}}\left(\delta_{i}+\tilde{\epsilon}_{i}\right),\quad\widehat{\mathrm{ATT}}^{0}=\tilde{M}^{-1}\sum\nolimits_{i=1}^{\tilde{M}}\tilde{Y}^{(1)}_{i}.$ (42)
The first estimator $\widehat{\mathrm{ATT}}^{*}$ knows the true value of $w^{*}$, and thus its MSE can be interpreted as the optimal performance an ATT estimator can achieve. The second estimator $\widehat{\mathrm{ATT}}^{0}$ essentially estimates $w$ by the zero vector and can be considered as the most naive estimator (note that it is still unbiased since entries of $\tilde{X}$ are randomly generated with mean zero). Define the relative efficiency of an estimator $\widehat{\mathrm{ATT}}$ by
$\mathrm{RE}(\widehat{\mathrm{ATT}})=\frac{\mathrm{MSE}(\widehat{\mathrm{ATT}}^{0})-\mathrm{MSE}(\widehat{\mathrm{ATT}})}{\mathrm{MSE}(\widehat{\mathrm{ATT}}^{0})-\mathrm{MSE}(\widehat{\mathrm{ATT}}^{*})}.$ (43)
If $\mathrm{RE}(\widehat{\mathrm{ATT}})=100\%$, then this estimator is as efficient as the optimal estimator $\widehat{\mathrm{ATT}}^{*}$.
Table 1: Relative efficiency (%) when $\|w^{*}\|_{1}=1$.
Setting $M$ BVS-SS Lasso OLS QP OLS($\gamma^{*}$) QP($\gamma^{*}$)
$N=20$ $J=5$ 25 92 74 1 90 96 95
50 94 88 69 95 98 99
100 93 92 90 97 95 96
200 99 98 100 99 98 99
$N=50$ $J=10$ 25 43 4 NA NA 80 83
50 88 80 NA 87 98 98
100 98 92 48 95 100 101
200 103 98 93 100 101 102
We report the relative efficiency of the six estimators with $\lambda=\|w^{*}\|_{1}=1$ in Tables 1. Our method outperforms Lasso substantially when $M=25$ or $50$, which shows that some prior knowledge about the simplex constraint can be very helpful when the sample size is small or moderate. By comparing BVS-SS with the two QP methods, we see that performing variable selection also significantly improves ATT estimation when the sample size is small. Indeed, when $N=50,M=25$, QP with the entire data set is not feasible since the underlying optimization problem does not have a unique solution. We also note that as sample size increases, the performance of BVS-SS quickly approaches that of the two oracle estimators, OLS$(\gamma^{*})$ and QP$(\gamma^{*})$.
Tables 2 and 3 present the results for $\|w^{*}\|_{1}=2$ or $3$ (i.e., the simplex constraint is violated by the true model). Note that since the regression coefficient estimator of Lasso and OLS is scale-equivariant, the relative efficiency of Lasso and two OLS methods remain the same as in Table 1. The most interesting observation from Tables 2 and 3 is that, though BVS-SS utilizes the simplex constraint, its performance remains robust. When $M=25$, BVS-SS still has a significant advantage over Lasso. Indeed, Lasso almost fails to recover any signal contained in $w^{*}$ when $N=50$ and $M=25$. Once the sample size becomes large (say, $M\geq 100$), BVS-SS again achieves close-to-optimal performance, and it outperforms QP$(\gamma^{*})$ by a wide margin, which knows the true model but assumes the simplex constraint. The robust and superior performance of BVS-SS across the three tables is largely due to the use of the soft simplex constraint, which enables BVS-SS to adapt to the unknown level of $\|w^{*}\|_{1}$ in the given data.
Table 2: Relative efficiency (%) when $\|w^{*}\|_{1}=2$.
Setting $M$ BVS-SS Lasso OLS QP OLS($\gamma^{*}$) QP($\gamma^{*}$)
$N=20$ $J=5$ 25 75 74 1 66 96 67
50 91 88 69 79 98 79
100 92 92 90 75 95 75
200 98 98 100 73 98 73
$N=50$ $J=10$ 25 42 4 NA NA 80 52
50 82 80 NA 76 98 80
100 96 92 48 81 100 82
200 101 98 93 78 101 78
Table 3: Relative efficiency (%) when $\|w^{*}\|_{1}=3$.
Setting $M$ BVS-SS Lasso OLS QP OLS($\gamma^{*}$) QP($\gamma^{*}$)
$N=20$ $J=5$ 25 78 74 1 46 96 47
50 92 88 69 62 98 62
100 92 92 90 57 95 57
200 98 98 100 55 98 55
$N=50$ $J=10$ 25 23 4 NA NA 80 35
50 85 80 NA 60 98 62
100 97 92 48 63 100 63
200 102 98 93 57 101 57
4.3 Estimation of Variance Parameters and Variable Selection
We now examine the performance of BVS-SS in terms of variable selection and the estimation of the two variance parameters $\tau,\phi$. First, we compare the variable selection result of BVS-SS with that of Lasso in Table 4, where “$\ell^{1}$-loss” denotes the $\ell^{1}$-distance between the selected model $\hat{\gamma}$ and the true model $\gamma^{*}$, and “model size” refers to the cardinality of $\hat{\gamma}$ (i.e., the number of nonzero entries of $\hat{w}$). For BVS-SS, the two statistics are averaged over the collected MCMC samples. It is evident that BVS-SS has a much higher accuracy than Lasso across all settings. Especially when $M$ is large, BVS-SS can correctly identify most entries of $\gamma$, while the performance of Lasso remains similar to that observed with small $M$.
Table 4: Variable selection performance of BVS-SS and Lasso averaged over 100 replicates.
BVS-SS Lasso
$\|w^{*}\|_{1}$ Setting $M$ $\ell^{1}$-loss model size $\ell^{1}$-loss model size
$\|w^{*}\|_{1}$ = 1 $N=20$ $J=5$ 25 3.1 3.9 6.3 10.2
50 1.9 4.8 6.7 11.1
100 1.3 5.1 6 10.9
200 0.8 5.4 6.3 11.3
$N=50$ $J=10$ 25 10.9 6.2 13 14.2
50 7.8 9.1 15.4 22.2
100 5.5 10.9 14.4 22.9
200 4.1 11.4 13.4 22.5
$\|w^{*}\|_{1}$ = 3 $N=20$ $J=5$ 25 3.3 3.3 6.3 10.2
50 1.9 4.8 6.7 11.1
100 1.2 5 6 10.9
200 0.7 5.2 6.3 11.3
$N=50$ $J=10$ 25 10.2 4.5 13 14.2
50 8.1 9.2 15.4 22.2
100 5.6 11.2 14.4 22.9
200 3.7 10.8 13.4 22.5
Next, we visualize the distribution of the posterior mean estimate of $\phi$ (error precision) across 100 replicates in Figure 1. This statistic reflects how well the BVS-SS model fits the data. We only show the result for $N=50$, as the distribution for $N=20$ is similar. Observe that regardless of $\|w^{*}\|_{1}$, the posterior mean of $\phi$ increases with $M$, which is expected since a larger sample size enables BVS-SS to detect more signals in the data, resulting in a smaller estimate of the error variance. When $\|w^{*}\|_{1}=3$, BVS-SS appears to always underestimate $\phi$ even when the sample size is sufficiently large, which may be related to the fact that the simplex constraint is significantly violated in this case. Notably, when $\|w^{*}\|_{1}=2$, the posterior estimation of $\phi$ appears highly accurate for $M\geq 100$, indicating that our model achieves an optimal balance between model complexity and fitting.
Finally, we examine the posterior mean estimate of $\tau$ when $N=50$, for which the result is presented in Figure 2. As discussed in Section 2.1, $\tau$ can be considered as an indicator of whether the simplex constraint is satisfied by the given data, and we now explain why Figure 2 provides compelling empirical evidence supporting this. When $\|w^{*}\|_{1}=1$, we see that the posterior mean of $\tau$ decreases as $M$ increases, since a larger sample size offers stronger evidence that the simplex constraint is satisfied, which draws $\tau$ towards zero. In contrast, when $\|w^{*}\|_{1}=2$ or $3$, the trend reverses, since a larger sample size enables BVS-SS to detect deviations from the simplex constraint. When $M=200$, we find that the ratio of the posterior mean estimate of $\tau$ and that of $\phi$ equals $0.0003$ for $\|w^{*}\|_{1}=1$, $0.024$ for $\|w^{*}\|_{1}=2$, and $0.081$ for $\|w^{*}\|_{1}=3$. This actually aligns well with the true data-generating mechanism: since we use $J=10$ for $N=50$, we may estimate the “true” mean squared deviation of $w^{*}$ from the simplex constraint by $0.1\times\sum_{j=1}^{10}(w^{*}_{j}-0.1)^{2}$, which equals $0.021$ for $\|w^{*}\|_{1}=2$ and $0.065$ for $\|w^{*}\|_{1}=3$.
Figure 1: Distribution of the posterior mean of $\phi$ across $100$ replicates with $N=50,J=10$. The true value $\phi^{*}$ is indicated by the dotted line.
Figure 2: Distribution of the posterior mean of $\tau$ across $100$ replicates with $N=50,J=10$.
5 Empirical Examples
In this section, we revisit two empirical examples from the SCM literature. We will use the notation introduced in Section 1.1: $T_{0}$ is the number of pre-treatment time periods, $T_{1}$ is the number of post-treatment time periods, the total number of units is $N+1$ where the first one is used as the response and the other $N$ units as the explanatory variables. (For BVS-SS, $M=T_{0}$ and $\tilde{M}=T_{1}$.) The first example has $N<T_{0}$, while the second has $N\gg T_{0}$, and it will be shown that our method works well in both settings.
5.1 Nota Fiscal Paulista: Anti-tax Evasion
An anti-tax evasion program, Nota Fiscal Paulista (NFP), was implemented in São Paulo, Brazil in October 2007 with the goal of reducing tax evasion by incentivizing consumers to request electronic receipts in exchange for participation in monthly lotteries promoted by the government, as well as chances for receiving partly tax rebates. Thus, consumers are encouraged to participate in the auditing schemes to decrease the chances of tax evasion. However, those restaurant owners and retailers may pass on part of the tax costs to consumers by raising product prices. Although the NFP was implemented sequentially across various sectors, including restaurants, bakeries, bars, and food service retailers, the literature has suggested that the food away from home (FAH) index could be used as a suitable indicator for price levels of these sectors. Carvalho et al. (2018) proposed an artificial counterfactual (ArCo) estimation procedure, by connecting Lasso to SCM without simplex constraint, to explore how the NFP affects the inflation on FAH. They also included monthly GDP growth, retail sales growth and monthly credit growth. Their ArCo estimation indicated that the ATT of NFP is 0.4478%.
We obtain the data analyzed in Carvalho et al. (2018) from the R package ArCo. The pre-treatment periods span $T_{0}=33$ months, and the post-treatment periods span $T_{1}=23$ months. Slightly different from the original study, the data includes only $N=8$ control units, which is one fewer than that mentioned in their paper, and there is no data related to credit and retail sales. However, this does not affect the application of our method, as our framework currently does not involve covariates beyond the control units.
We run our Algorithm 1 for 1000 iterations to estimate the impact of NFP. Figure 3 shows the trajectories of the posterior samples of $\phi$ and $\log(\tau)$ (with burn-in included). Probably due to the limited number of observations, the MCMC samples of $\tau$ exhibit large variability, indicating that the simplex assumption is hard to confirm or reject. The trace plot suggests that our sampler has converged fast. Table 5 summarizes the posterior mean and 95% credible intervals of ATT, $\tau$, $\phi$ and the model size $|\gamma|$; see Figure 5 in the Appendix for histograms of the MCMC samples and Figure 7 in the Appendix for the counterfactual estimation. The first half of the samples have been dropped for burn-in. The ATT estimation is close to that of (Carvalho et al., 2018, Column 1, Table 5), but our credible interval is much smaller compared to their confidence set. While the analysis of Carvalho et al. (2018) selected all the predictors, our method only chooses around two areas for constructing the synthetic controls on average. This is consistent with the findings of our simulation study presented in Section 4, which shows that Lasso has the tendency of over-estimating the model size compared to BVS-SS.
Figure 3: The trace plot of $\log(\tau)$ and $\phi$ for the NFP data set.
Table 5: NFP Impact on Food Inflation
ATT $\tau$ $\phi$ $|\gamma|$
Mean 0.288 0.14 3.85 2.29
95% credible interval (0.141, 0.419) (0, 1.44) (1.94, 6.05) (1, 4)
5.2 China’s anti-corruption campaign
China launched an unprecedented anti-corruption campaign in November 2012, aimed at curbing corruption and power abuse within government and military departments. In the context of bribery, direct cash payments are often considered too blatant and are generally unpopular, and high-end imported luxury goods serve as a more discreet way to convey the value of a gift in line with cultural norms of subtlety. Empirical studies, such as Lan and Li (2018), have shown that there is a comovement between the importation of luxury watches in China and changes in the government leadership. An interesting study was conducted by Shi and Huang (2023), who investigated the effect of China’s anti-corruption campaign on the importation of luxury watches. They took the monthly growth rate of luxury watch imports in US dollars as the outcome of interest and selected the control group among the growth rates of $N=87$ commodity categories to construct the synthetic counterfactual for luxury watch imports. The study utilizes data from the United Nations Comtrade Database, focusing on the category “watches with case of, or clad with, precious metal.” January 2013 is considered the time of treatment, which was the month immediately following the announcement of the Eight-Point Policy (an anti-corruption policy). The pre-treatment period spans from February 2010 to December 2012 resulting in $T_{0}=35$ observations, and the post-treatment period covers January 2013 to December 2015, which yields $T_{1}=36$ observations. Their forward selection algorithm identified three control units: “knitted or crocheted fabric,” “cork and articles of cork,” and “salt, sulfur, earth, stone, plaster, lime, and cement.” Their estimated counterfactual predicted that without the anti-corruption campaign, luxury watch imports would have continued to increase. However, in reality, imports dropped by 42% in January 2013, while the counterfactual predicted a 1.7% increase. Their treatment effect estimation indicated a reduction of 3.09% in luxury watch imports per month over the post-treatment period. Accumulated over 36 months, the campaign reduced total luxury watch imports by approximately two-thirds.
We obtained the data of Shi and Huang (2023) from the R package fdPDA, which contains all the 88 commodity categories (including the response and $N=87$ control categories). Figure 4 shows the MCMC trajectories of the posterior samples of $\phi$ and $\log(\tau)$. Compared to the previous real-data example, the posterior distribution of $\tau$ is still dispersed, but the ratio $\tau/\phi$ becomes much smaller, indicating that the simplex constraint is more likely to be satisfied (at least approximately) in this data set. The estimation of $\phi$ appears to be much more accurate, suggesting that the error variance is around $0.05$. As detailed in Table 6, our ATT estimation is smaller than that of Shi and Huang (2023) in absolute value, but the 95% credible interval indicates that the anti-corruption has a very significant effect on the high-end watch imports. On average, our model size $|\gamma|$ is larger than that of Shi and Huang (2023) (who only selected 3 control categories), but we note that the posterior distribution of $|\gamma|$, which is provided in Figure 6 in the Appendix, still concentrates on small models with size $\leq 5$. Figure 8 in the Appendix depicts the counterfactual estimation, which indicates a dramatic fall in luxury watch importation right after the treatment.
Figure 4: The trace plot of $\log(\tau)$ and $\phi$ for the anti-corruption data set.
Table 6: Anti-corruption Campaign’s Impact on Luxury Watches Importation
ATT $\tau$ $\phi$ $|\gamma|$
Mean -0.021 0.069 20.86 5.09
95% credible interval (-0.032, -0.008) (0, 0.641) (12.22, 32.76) (1, 20)
6 Conclusion
We propose a novel Bayesian synthetic control method, BVS-SS, that integrates a relaxed simplex constraint with Bayesian spike-and-slab variable selection. Our approach introduces a hierarchical prior to determine how closely the data should adhere to the simplex constraint, a feature that is missing in existing frequentist or Bayesian synthetic control frameworks and helps mitigate the ongoing debate surrounding the use of simplex constraints. To efficiently compute the posterior distribution of BVS-SS, we develop a novel Metropolis-within-Gibbs sampler, which overcomes the hindrance caused by the simplex constraint by updating two regression coefficients simultaneously from the full conditional posterior. Simulation studies and real-data examples illustrate the advantages and usefulness of our proposed algorithm.
There are several interesting directions that future researchers may dive in. First, it is known that Gibbs sampling methods are susceptible to high collinearity among the explanatory variables, which is a common situation in SCM applications where factor models are often used as the underlying data-generating process. Though it is unclear if such issues affect the methodology proposed in this work, since our Algorithm 1 updates two coordinates at a time, it would be interesting to investigate the use of other MCMC techniques, such as pseudo-marginal sampling (Andrieu and Roberts, 2009). Second, to include time-invariant, time-variant predictors or characteristics, one can extend BVS-SS by considering two types of explanatory variables, with one type following the spike-and-slab prior and the other following a regular non-informative prior without variable selection. Third, more numerical studies or model extensions can be performed to study the performance of BVS-SS when the model is mis-specified (e.g., the data is non-stationary, heterogeneous error), which would be of great interest to economists. We hope our method will serve as an inspiration for future research and facilitate the widespread use of Bayesian synthetic control methods.
Acknowledgements
The work of QZ was supported in part by the NSF grant DMS-2311307 and DMS-2245591.
Appendices
Appendix A Proofs
Proof of Lemma 1.
We can express the conditional posterior of $(\mu_{i},\mu_{j})$ by
$p(\mu_{i},\mu_{j}\mid y,\phi,\tau,\mu_{-(i,j)})=C\,p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})p(\mu_{i},\mu_{j},\mu_{-(i,j)}),$ (44)
where $C$ is a constant that depends on $\mu_{-(i,j)}$ but not $\mu_{i},\mu_{j}$, and the marginal likelihood $p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})$ is shown in (14)
First, look at $\mu$’s prior $p(\mu_{i},\mu_{j},\mu_{-(i,j)})$. Recall that $\mu|\gamma\sim\text{sym-Dirichlet}(\alpha)$. Denote $p_{0}(|\gamma|)=\theta^{|\gamma|}(1-\theta)^{N-|\gamma|}$ for a short notation. Given $\alpha=1$, $\mu$ is uniform on the $\Delta^{|\gamma|-1}$. To satisfy the simplex constraint, given $\sum_{k\neq i,j}\mu_{k}\textless 1$, the density of $\mu_{i}=\mu_{j}=0$ must be zero. Noticing that $p(\mu)=p(\mu|\gamma)p(\gamma)$, we have:
$p(\mu_{i}=0,\mu_{j}=0,\mu_{-(i,j)})=\;$ $0,$ (45)
$p(\mu_{i}=s,\mu_{j}=0,\mu_{-(i,j)})=\;$ $(\ell)!\,p_{0}(\ell+1),$ (46)
$p(\mu_{i}=t,\mu_{j}=s-t,\mu_{-(i,j)})=\;$ $(\ell+1)!\,p_{0}(\ell+2),\quad\forall t\in(0,s),$ (47)
Integrating over $t$ in the last expression, we obtain that
[…]
$=s(\ell+1)!\,p_{0}(\ell+2)$ (49)
Recall the definition in Lemma 2:
$\check{y}(u)=y-uX_{i}-(s-u)X_{j}-\sum\nolimits_{k\neq i,j}\mu_{k}X_{k}.$ (50)
We can express the conditional posterior density of $(\mu_{i},\mu_{j})=(s,0)$ by
$p(\mu_{i}=s,\mu_{j}=0\mid y,\phi,\tau,\mu_{-(i,j)})$ $\propto A(\gamma^{i},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(s)^{\top}\Sigma(\gamma^{i},\tau)\check{y}(s)\right\},$ (51)
where
$A(\gamma,\tau)=\mathrm{det}(V_{\gamma,\tau})^{-1/2}(|\gamma|-1)!\,p_{0}(|\gamma|),\quad\Sigma(\gamma,\tau)=I-X_{\gamma}V_{\gamma,\tau}^{-1}X_{\gamma}^{\top}.$ (52)
The conditional posterior density of $(\mu_{i},\mu_{j})=(0,s)$ has the same expression with $\check{y}(s)$ replaced by $\check{y}(0)$ and $\gamma^{i}$ replaced by $\gamma^{j}$. The conditional posterior density of $(\mu_{i}>0,\mu_{j}>0)$ is:
$p(\mu_{i}=u,\mu_{j}=s-u\mid y,\phi,\tau,\mu_{-(i,j)})\propto A(\gamma^{i,j},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma(\gamma^{i,j},\tau)\check{y}(u)\right\}.$ (53)
Observe that the density is proportional to the structure of an exponential of $t$’s quadratic terms, and $t$ is bounded in $(0,s)$. $t$ must follow a truncated Gaussian distribution restricted to the interval of $(0,s)$. Rewrite the above exponential part as:
$\exp\left\{-\frac{\phi}{2}\left((X_{j}-X_{i})^{\top}\Sigma(\gamma^{i,j},\tau)(X_{j}-X_{i})t^{2}-2((X_{i}-X_{j})^{\top}\Sigma(\gamma^{i,j},\tau)y^{j})+C_{3}\right)\right\}.$ (54)
where $C_{3}$ is some constant not containing $t$. Then it is clear to see that $\mu_{i}$’s conditional posterior follows a truncated Gaussian distribution as stated in Lemma 1. ∎
Proof of Lemma 2.
For simplicity, we assume $\alpha=1$ as in Lemma 1. $A(\gamma,\tau)$ collapses to $\mathrm{det}(V_{\gamma,\tau})^{-1/2}(|\gamma|-1)!\,p_{0}(|\gamma|)$. The extension is almost trivial. Given $\mu_{-(i,j)}$, $\mu_{i}=s$ and $\mu_{j}=0$, $\gamma_{i}=1$ and $\gamma_{j}=0$ with probability equal to one. Thus,
$p(\gamma^{i}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $p(\gamma_{i}=1,\gamma_{j}=0\mid y,\mu_{-(i,j)},\tau,\phi)$ (55)
=\;= $p(\mu_{i}=s,\mu_{j}=0\mid y,\mu_{-(i,j)},\tau,\phi)$ (56)
=\;= $A(\gamma^{i},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(s)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(s)\right\}$ (57)
Similarly,
$p(\gamma^{0}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $0,$ (58)
$p(\gamma^{j}\mid y,\mu_{-(i,j)},\tau,\phi)\propto\;$ $A(\gamma^{j},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(0)^{\top}\Sigma_{\gamma^{j},\tau}\check{y}(0)\right\},$ (59)
To find the conditional posterior of $\gamma^{ij}$, we need to integrate out $u$:
[…]
=\;= $A(\gamma^{i,j},\tau)\exp\left\{-\frac{\phi}{2}\left(\check{y}(u)^{\top}\Sigma_{\gamma^{i,j},\tau}\check{y}(u)^{\top}-\beta_{i,j}^{2}\Lambda_{i,j}\right)\right\}$ (62)
$\frac{\sqrt{2\pi}}{\sqrt{\phi\Lambda_{i,j}}}\int_{0}^{s}\frac{\sqrt{\phi\Lambda_{i,j}}}{\sqrt{2\pi}}\exp\left\{-\frac{\phi\Lambda_{i,j}}{2}\left(u-\beta_{i,j}\right)^{2}\right\}\mathrm{d}u,$ (63)
Observe that the integration part equals to an integral of a truncated Gaussian variable, and thus can be considered as a normalizing constant:
$\int_{0}^{s}\frac{\sqrt{\phi\Lambda_{i,j}}}{\sqrt{2\pi}}\exp\left\{-\frac{\phi\Lambda_{i,j}}{2}\left(u-\beta_{i,j}\right)^{2}\right\}\mathrm{d}u=\;\Phi\left((s-\beta_{i,j})\sqrt{\phi\Lambda_{i,j}}\right)-\Phi\left(-\beta_{i,j}\sqrt{\phi\Lambda_{i,j}}\right),$ (64)
where $\Phi$ denotes the CDF of $N(0,1)$. ∎
Appendix B Supplementary Empirical Examples Results
In Figures 5 and 6 we visualize the posterior distributions of $\tau$, $\phi$, model size and ATT estimate for the two real data sets considered in Section 5.
Figure 5: Posterior distributions for the Nota Fiscal Paulista data set.
Figure 6: Counterfactual Estimations for the China’s Anti-corruption Campaign data set.
Figure 7: Counterfactual estimation for the Nota Fiscal Paulista data set.
Figure 8: Counterfactual estimation for the China’s Anti-corruption Campaign data set.