### v2

# Bayesian Synthetic Control with a Soft Simplex Constraint

Abstract
The challenges posed by high-dimensional data and use of the simplex constraint are two major concerns in the empirical application of the synthetic control method (SCM) in econometric studies. To address both issues simultaneously, we propose a Bayesian SCM that integrates a soft simplex constraint within spike-and-slab variable selection. The hierarchical prior structure captures the extent to which the data supports the simplex constraint, allowing for more efficient and data-adaptive counterfactual estimation. The intractable marginal likelihood induced by the soft simplex constraint presents a major computational challenge, which we resolve by developing a novel Metropolis-within-Gibbs algorithm that updates the regression coefficients of two predictors simultaneously. Our main theoretical contribution is a high-dimensional selection consistency result for the spike-and-slab variable selection under the simplex constraint, which significantly extends the current theory for high-dimensional Bayesian variable selection. Simulation studies demonstrate that our method performs well across diverse settings. To illustrate its practical values, we apply it to two empirical examples for estimating the effect of economic policies.
Keywords: average treatment effect; constrained linear regression; Metropolis-within-Gibbs sampler; spike-and-slab prior; strong selection consistency.
1 Introduction
1.1 Background on Synthetic Control Methods
In many fields of social science, researchers need to identify the causal effect of an intervention or treatment of interest using only observational data, since conducting case-control experiments is either infeasible or improper. The most commonly used causal inference method in economics is the Difference-in-Differences approach. Its validity, however, relies heavily on the parallel trends assumption: the control unit must be sufficiently similar to the treated unit so that it can mimic the pre-intervention dynamics of the treated. In many empirical settings, this assumption is overly restrictive, especially when the outcome of interest is measured at an aggregate level, such as a country’s GDP. The synthetic control method (SCM) was developed to overcome this limitation, which constructs a counterfactual outcome of the treated unit using a weighted average of the control units that closely tracks the pre-treatment trajectory of the treated. For example, in the seminal work of abadie2003economic, SCM was developed to study how terrorist conflicts affected the local economy in the Basque Country between the 1970s and 1990s.
Formally, SCM can be described as follows. Consider a panel with $T=T_{0}+T_{1}$ total observations over time for $N+1$ units, where only the first unit receives the treatment after period $T_{0}$. The remaining $N$ units form the control group, which are unaffected by the intervention. Let $Z_{i,t}$ denote the outcome of interest for unit $i$ at time $t$, where i=1,2,,N+1i=1,2,...,N+1 and t=1,2,,Tt=1,2,...,T. The objective of SCM is to estimate the average treatment effect on the treated (ATT) by estimating the counterfactual outcome $Z_{1,t}^{(0)}$ for the treated unit at times $t=T_{0}+1,\dots,T$, which represents what the outcome of unit 1 would have been in the absence of the treatment. The standard approach is to define $Z_{1,t}^{(0)}$ as a weighted combination of the outcomes of the $N$ untreated units:
$Z_{1,t}^{(0)}(w)=\sum_{i=2}^{N+1}w_{i-1}Z_{i,t},$ (1)
where the weights $(w_{i})_{i=1}^{N}$ are time-independent, non-negative and sum to one. A larger $w_{i}$ indicates that the $i$-th control unit better approximates the treated unit’s pre-intervention characteristics. To find the optimal set of weights $(w_{i})_{i=1}^{N}$ such that the synthetic control approximates the pre-treatment outcome of the treated unit as closely as possible, a common strategy is to minimize the mean squared error between the treated unit and the synthetic control in the pre-treatment period $t=1,\dots,T_{0}$:
$\hat{w}$ $=\operatorname*{argmin}_{w\in\Delta^{N-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\left(Z_{1,t}-\sum_{i=2}^{N+1}w_{i-1}Z_{i,t}\right)^{2},$ (2)
$\text{ where }\Delta^{N-1}$ $=\left\{u\in\mathbb{R}^{N}\colon u_{i}\geq 0\text{ for each }i,\text{ and }\sum\nolimits_{i=1}^{N}u_{i}=1\right\}.$ (3)
[…]
$\widehat{\mathrm{ATT}}=\frac{1}{T_{1}}\sum_{t=T_{0}+1}^{T}\left(Z_{1,t}-\sum_{i=2}^{N+1}\hat{w}_{i-1}Z_{i,t}\right).$ (4)
Since its introduction, SCM has gained widespread adoption in empirical research, and athey2017state described SCM as “arguably the most important innovation in the policy evaluation literature in the last 15 years.” However, the growing application of SCM has also revealed new challenges. The first is the high dimensionality: when the number of the control units is comparable to, or even larger than, the number of pre-treatment periods, conventional least-squares-based methods are no longer applicable. To address this issue, researchers have proposed alternatives for computing the weights in (1), such as penalized regression methods (doudchenko2016balancing, carvalho2018arco), factor-model-based approaches (xu_2017, pang2022bayesian), or variable selection (kim2020bayesian, shi2023forward). However, none of these methods utilizes the simplex constraint and thus the resulting estimator $\hat{w}$ loses its interpretation as weights. How to implement SCM in a high-dimensional setting while preserving the simplex constraint remains largely unclear.
This naturally leads to another major issue with SCM: whether the simplex constraint should be imposed. Proponents argue that the simplex constraint prevents extrapolation (king2006dangers, abadie2021syntheticcontrol). Critics, however, argue that the simplex constraint makes the extension of SCM to high-dimensional settings complicated and is often too restrictive in practice. A well-known example is the study of the economic impact of Hong Kong’s reunification with China in 1997 (hsiao2012panel). Since no other Asian regions faced a similar “handover” situation as Hong Kong, it is natural to apply SCM to estimate the impact. However, since Hong Kong was long known for its outstanding economic performance at that time (known as the leading region in the “Four Asian Tigers”), researchers found it unreasonable to use a convex combination of other regions to track the economic trajectory of Hong Kong. Allowing the sum of the coefficients to exceed one is in fact consistent with Hong Kong’s historically strong economic performance.
In this work, we develop a new Bayesian method for building the synthetic control in high-dimensional settings. Bayesian SCM has seen growing use in the recent literature for its flexibility in integrating with high-dimensional techniques. In variable selection context, Bayesian approach enables statistical inference by model averaging (kass1995bayes), which contrasts with penalized regression methods like Lasso that focus on identifying a single best model. Further, Bayesian SCM provides posterior distributions for both ATT and the counterfactual outcome, which are more informative than point estimates. A potential limitation, however, is that the posterior fitting often requires the use of Markov chain Monte Carlo (MCMC) sampling, which can be very time-consuming compared to optimization-based methods. One recent work by goh2022synthetic considered Bayesian linear regression where the prior distribution of the regression coefficients is uniform over a convex hull, corresponding to the simplex constraint proposed by abadie2003economic. They also discussed the Bayesian formulation of the method of li2020statistical, which involves using a prior over the shifted conical hull. A similar model was studied in martinez2022bayesian, and a Bernstein-von Mises result for the Bayes estimator was proved. kim2020bayesian proposed using either the horseshoe prior or the spike-and-slab prior, two well-known Bayesian approaches to variable selection, to construct synthetic control with selected control units in high-dimensional settings; the simplex constraint was dropped in their method. It was demonstrated in their paper that their method outperforms those frequentist SCM methods based on variable selection. Nevertheless, to our knowledge, there is no existing Bayesian variable selection method for constructing synthetic control that also incorporates the simplex constraint.
1.2 Overview of this Work
Motivated by the high-dimensional challenge and the mixed benefits and limitations of the simplex constraint, we propose a new SCM based on a Bayesian model called BVS-SS (Bayesian Variable Selection with Soft Simplex constraint), which captures the sparsity among control units and allows the data to determine whether the simplex constraint is appropriate on a case-by-case basis. BVS-SS aims to leverage the strengths of both simplex-constrained SCM and unconstrained regression-based methods, which serves as an effective solution to the debate surrounding the use of this constraint. There is no similar method in the frequentist SCM literature to our knowledge.
In Section 2 we formally introduce the BVS-SS model, which utilizes a standard spike-and-slab-type prior for selecting which control units should be used for constructing the synthetic control, a technique that has gained significant success and widespread popularity in the Bayesian community (george1993variable, george1997approaches, brown1998multivariate). Different from existing models, BVS-SS assumes that there is an underlying sparse “mean weight vector” $\mu$ that satisfies the simplex constraint, but the actual weights $w_{1},\dots,w_{N}$ involved in (1) may deviate from $\mu$. Next, we introduce a key variance parameter $\tau$ quantifying the amount of this deviation. By imposing a non-informative prior on $\tau$, we let BVS-SS learn its value from the data, revealing whether the data actually agrees with the simplex assumption. A similar model was utilized in the theoretical analysis of martinez2022bayesian, but they did not consider variable selection and the implementation of the model.
To generate samples from the posterior distribution of BVS-SS, in Section 3 we propose an original Metropolis-within-Gibbs sampler, where $\mu$ and error variance are updated by Gibbs schemes and $\tau$ is updated by Metropolis-Hastings steps. Our Gibbs-type updating for $\mu$ is different from existing Gibbs schemes for similar problems, as those methods typically only update one coordinate at a time, which is not feasible in our context due to the simplex constraint (i.e., given $\mu_{-j}$, we have $\mu_{j}=1-\sum_{i\neq j}\mu_{i}$ almost surely). We overcome this challenge by theoretically finding the joint conditional posterior distribution of two coordinates $(\mu_{i},\mu_{j})$ given all the other parameters, which enables us to update two coordinates simultaneously. This full conditional posterior distribution is complicated due to the spike-and-slab prior for variable selection, but sampling from it can be performed straightforwardly and efficiently. Using these MCMC samples, we can compute the posterior mean estimates and credible intervals of the counterfactual and ATT without extra computational cost.
In Section 4, we theoretically investigate the effect of the simplex constraint and establish high-dimensional consistency results for BVS-SS with $\tau$ fixed to zero (i.e., regression coefficients are forced to satisfy the simplex constraint). In Theorem 2, we show that the posterior inclusion probability of the true model converges to $1$ in probability under mild assumptions on the true data-generating process (DGP). While similar results have been obtained for various Bayesian variable selection models (yang2016computational), the presence of the simplex constraint introduces a substantial new challenge for the analysis: the marginal likelihood for each model in BVS-SS does not admit a closed-form expression and instead involves an intractable integral over the simplex. To overcome this difficulty, we leverage techniques from constrained least-squares estimation and Gaussian concentration inequalities to derive closed-form upper and lower bounds for the marginal likelihood. The signal size required for detection has roughly the same order as that in the unconstrained case (yang2016computational); see Assumption 20. To our knowledge, this is the first consistency result for high-dimensional Bayesian variable selection under geometric constraints on the regression coefficients. Moreover, building on this result, we establish the order of the posterior expected predictive loss in Theorem 3, providing theoretical guarantees on the ATT estimation. Our results complement the existing theoretical developments for frequentist SCM (abadie2010synthetic, firpo2018synthetic, carvalho2018arco, li2020statistical, chernozhukov2021exact).
Section 5 presents simulation studies which show that BVS-SS delivers robust and better performance than conventional methods, regardless of whether the simplex constraint is satisfied by the true DGP. Moreover, the posterior estimate of $\tau$ successfully captures the extent to which the true signals deviate from the simplex constraint. Section 6 presents two real-data applications of ATT estimation with BVS-SS, one from carvalho2018arco on the effectiveness of an anti-tax evasion program and the other from shi2023forward on the impact of an anti-corruption policy. Section 7 concludes the paper with a brief discussion. Details of the sampling algorithm, all proofs and additional numerical results are deferred to the supplementary materials.
2 The BVS-SS Model
2.1 Model and Prior
For generality and ease of notation, henceforth we use $Y\in\mathbb{R}^{M}$ to denote a response vector of $M$ observations and $X\in\mathbb{R}^{M\times N}$ to denote an arbitrary design matrix with $N$ explanatory variables. Assume that
$Y=Xw+\epsilon,\quad\epsilon\sim N(0,\phi^{-1}I),$ (5)
where $w\in\mathbb{R}^{N}$ is the vector of unknown regression coefficients, $\epsilon$ is the vector of i.i.d. Gaussian noise with variance $\phi^{-1}$, and $I$ denotes the identity matrix. For the synthetic control problem described in Section 1.1, we have $M=T_{0}$, $Y_{t}=Z_{1,t}$ and $X_{tj}=Z_{j+1,t}$ for $t=1,\dots,T_{0}$ and $j=1,\dots,N$.
We propose a hierarchical prior distribution on $w$ which induces sparsity and incorporates the simplex constraint. To describe it, introduce an indicator vector $\gamma\in\{0,1\}^{N}$ such that $\gamma_{i}=1$ if and only if $w_{i}\neq 0$. Let $|\gamma|=\sum_{j=1}^{N}\gamma_{j}$ denote the number of selected predictors. Given a vector $w$, we use $w_{\gamma}$ to denote the subvector of $w$ with entries indexed by $\{i\colon\gamma_{i}=1\}$. We consider the following prior on $(w,\phi)$:
$\phi\sim$ $\mathrm{Gamma}(\kappa_{1}/2,\kappa_{2}/2),$ (6)
$\gamma_{i}\overset{\mathrm{i.i.d.}}{\sim}$ $\mathrm{Bernoulli}(\theta),$
$\tau\sim$ $\mathrm{Gamma}(a_{1},a_{2}),$
$\mu_{\gamma}\mid\gamma\sim$ $\text{sym-Dirichlet}(\alpha),$
$w_{\gamma}\mid\gamma,\mu_{\gamma},\tau,\phi\sim$ $N(\mu_{\gamma},\,(\tau/\phi)I),$
and when $\gamma_{i}=0$, we let $\mu_{i}=w_{i}=0$ with probability one. In (6), $\kappa_{1},\kappa_{2},a_{1},a_{2},\alpha>0$ and $\theta\in(0,1)$ are fixed hyperparameters, $\text{sym-Dirichlet}(\alpha)$ denotes the symmetric Dirichlet distribution on the simplex with concentration parameter $\alpha$ (the dimension of the simplex is clear from context), and the Gamma distribution is in shape-rate parameterization. When $\alpha=1$ and $|\gamma|=\ell$, the prior of $\mu_{\gamma}$ is a uniform distribution on $\Delta^{|\gamma|-1}$ with density $p(\mu_{\gamma})=\Gamma(\ell)$ for each $\mu_{\gamma}\in\Delta^{|\gamma|-1}$, where $\Gamma(\ell)=(\ell-1)!$ is the inverse volume of the simplex.
In plain words, we assume that given $\gamma$ and $\phi$, $w_{\gamma}$ follows a normal distribution with mean vector $\mu_{\gamma}\in\Delta^{|\gamma|-1}$ and prior variance depending on the parameter $\tau$. This conditional prior for $w$ can be viewed as an interpolation between the simplex constraint and the unconstrained setting. If the data suggests that $w$ is likely to satisfy the simplex constraint, then the posterior distribution of $\tau$ should concentrate around zero, while if the simplex constraint is significantly violated, $\tau$ should stay away from zero in the posterior. As will be demonstrated in our simulation study, $\tau$ can be reasonably estimated even with a relatively small sample size, which tells the user if the simplex constraint is appropriate for the data set being analyzed. This approach effectively reconciles the different views on the use of the simplex constraint, and thus we call our model BVS-SS (Bayesian Variable Selection with Soft Simplex constraint). We note that our prior for $w$ is similar to that studied in martinez2022bayesian; however, they did not consider variable selection, and in their implementation they assumed that $w$ satisfies the hard simplex constraint (though the soft constraint was used in their theory). Regarding the hyperparameters, $a_{1},a_{2}$ should be chosen to reflect one’s prior belief on whether simplex constraint is likely to hold, and a larger value of $\alpha$ encourages a more even distribution of the weights $\{\mu_{i}\colon\gamma_{i}=1\}$. In practice, one often has limited information about the true weight vector $w$, but the effect of these hyperparameters quickly becomes negligible as sample size increases.
The other elements of our model are standard. The prior for $\phi$ is conjugate, and one can choose small values for $\kappa_{1},\kappa_{2}$ as a noninformative prior. The hyperparameter $\theta$ is the marginal prior probability of an explanatory variable being included in the regression model, and a smaller value of $\theta$ represents a heavier penalty on the model complexity.
2.2 Posterior Distribution and ATT Estimation
Let $X_{\gamma}$ denote the submatrix of $X$ with columns indexed by $\{i\colon\gamma_{i}=1\}$. Under the prior distribution of BVS-SS, we have
$Y\mid\gamma,\mu_{\gamma},\tau,\phi\sim N\left(X_{\gamma}\mu_{\gamma},\,\phi^{-1}(\tau X_{\gamma}X_{\gamma}^{\top}+I)\right).$ (7)
A routine calculation using Woodbury identity yields the marginal likelihood of $(\mu,\tau,\phi)$:
$p(y\mid\gamma,\mu_{\gamma},\tau,\phi)$ $\propto\frac{\phi^{M/2}}{\tau^{|\gamma|/2}\mathrm{det}(V_{\gamma,\tau})^{1/2}}\exp\left\{-\frac{\phi}{2}(y-X_{\gamma}\mu_{\gamma})^{\top}\Sigma_{\gamma,\tau}(y-X_{\gamma}\mu_{\gamma})\right\},$ (8)
$\text{ where }V_{\gamma,\tau}$ $=X_{\gamma}^{\top}X_{\gamma}+\tau^{-1}I,\quad\Sigma_{\gamma,\tau}=I-X_{\gamma}V_{\gamma,\tau}^{-1}X_{\gamma}^{\top}.$ (9)
Further, using the conjugacy of normal-gamma prior, we find that the conditional posterior distributions of $w_{\gamma}$ and $\phi$ are given by
$\phi\mid y,\gamma,\mu_{\gamma},\tau\sim\;$ $\mathrm{Gamma}\left(\frac{M+\kappa_{1}}{2},\,\frac{\kappa_{2}+(y-X_{\gamma}\mu_{\gamma})^{\top}\Sigma_{\gamma,\tau}(y-X_{\gamma}\mu_{\gamma})}{2}\right),$ (10)
$w_{\gamma}\mid y,\gamma,\mu_{\gamma},\tau,\phi\sim\;$ $N\left(V_{\gamma,\tau}^{-1}(X^{\top}_{\gamma}y+\tau^{-1}\mu_{\gamma}),\,\phi^{-1}V_{\gamma,\tau}^{-1}\right).$ (11)
Using the prior independence between $\mu,\tau,\phi$, we can express the joint posterior distribution by $p(\gamma,\mu_{\gamma},\tau,\phi\mid y)\propto p(y\mid\gamma,\mu_{\gamma},\tau,\phi)p(\mu_{\gamma}\mid\gamma)p(\gamma)p(\tau)p(\phi).$
Prediction with BVS-SS can be conducted as follows. Let $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ denote the design matrix of another $\tilde{M}$ observations, and suppose we are interested in estimating $\tilde{X}w$. Letting $\mathbb{E}_{y}$ denote the posterior distribution given $Y=y$ and using (11), we obtain that
$\mathbb{E}_{y}[\tilde{X}w]=\mathbb{E}_{y}\left[\mathbb{E}_{y}(\tilde{X}w\mid\gamma,\mu_{\gamma},\tau,\phi)\right]=\mathbb{E}_{y}\left[\tilde{X}V_{\gamma,\tau}^{-1}(X^{\top}_{\gamma}y+\tau^{-1}\mu_{\gamma})\right].$ (12)
Given $n$ MCMC samples $(\gamma^{(k)},\mu_{\gamma}^{(k)},\tau^{(k)},\phi^{(k)})_{k=1}^{n}$, we can approximate the above expectation using the sample average, which yields the posterior predictive mean of $\tilde{X}w$. The covariance matrix of the posterior predictive distribution of $\tilde{X}w$ can be calculated similarly.
For synthetic control, the goal is to estimate ATT among the $\tilde{M}$ observations. Denote their response under the treatment by $\tilde{Y}^{(1)}$, which is assumed observed, and denote the counterfactual response without treatment by $\tilde{Y}^{(0)}$. Let $\delta=\tilde{Y}^{(1)}-\tilde{Y}^{(0)}$, and define ATT by
$\mathrm{ATT}=\frac{1}{\tilde{M}}\sum_{i=1}^{\tilde{M}}\delta_{i}=\frac{1}{\tilde{M}}\sum_{i=1}^{\tilde{M}}\left(\tilde{Y}^{(1)}_{i}-\tilde{Y}^{(0)}_{i}\right)$, (13)
which is the main parameter of interest. By replacing $\tilde{Y}^{(0)}$ using the synthetic control $\mathbb{E}_{y}[\tilde{X}w]$, we obtain an estimator for $\bar{\delta}$. If we have samples $(w^{(k)})_{k=1}^{n}$ approximating the marginal posterior distribution of $w$, we can also obtain the posterior predictive distribution of ATT by plugging $w^{(k)}$ into the estimator $\tilde{M}^{-1}\sum_{i=1}^{\tilde{M}}(\tilde{Y}^{(1)}_{i}-w^{\top}\tilde{X}_{(i)})$, where $\tilde{X}_{(i)}$ denotes the $i$-th row of $\tilde{X}$ (treated as a column vector).
3 Metropolis-within-Gibbs Sampling for BVS-SS
The standard solution to posterior computation for the spike-and-slab variable selection is to run a Metropolis–Hasting sampler targeting the posterior distribution of $\gamma$ by add/delete/swap proposals; see george1997approaches, guan2011bayesian, chang2024dimension, among many others. However, this approach is not applicable in our setting since the likelihood $p(y\mid\gamma,\phi)$ in BVS-SS does not have a closed-form expression.
Instead of directly sampling from the marginal posterior of $\gamma$, we devise a Metropolis-within-Gibbs scheme targeting the full posterior distribution $p(\gamma,\mu,\tau,\phi\mid y)$. The Metropolis–Hastings update for $\tau$ and the Gibbs update for $\phi$ are standard, which we detail in Appendix S1.1. Since $\gamma_{i}=\mathbbm{1}_{\{\mu_{i}\neq 0\}}$, we update $(\gamma_{i},\mu_{i})$ simultaneously, which can be implemented straightforwardly as a Gibbs step in the unconstrained spike-and-slab variable selection. However, this is again infeasible for BVS-SS due to the simplex constraint: given $\mu_{-j}$, we have $\mu_{j}=1-\sum_{i\neq j}\mu_{i}$ almost surely. Therefore, we propose to a novel Gibbs scheme where we update two coordinates of $(\gamma,\mu)$ simultaneously.
Let $\mu_{-(i,j)}$ denote the subvector of $\mu$ with $(\mu_{i},\mu_{j})$ removed and define $\gamma_{-(i,j)}$ analogously. To implement the Gibbs update of $(\gamma_{i},\gamma_{j},\mu_{i},\mu_{j})$ for $i\neq j$, we need to find the full conditional posterior $p(\gamma_{i},\gamma_{j},\mu_{i},\mu_{j}\mid y,\gamma_{-(i,j)},\mu_{-(i,j)},\tau,\phi)$. If $\sum_{k\neq i,j}\mu_{k}=1$, neither $i$ nor $j$ can be selected due to the simplex constraint, and thus $\gamma_{i}=\gamma_{j}=\mu_{i}=\mu_{j}=0$. Now assume $\sum_{k\neq i,j}\mu_{k}<1$, in which case we must have $(\gamma_{i},\gamma_{j})=(1,0),(0,1)$ or $(1,1)$. The conditional posterior probability $p(\gamma_{i}=1,\gamma_{j}=0\mid y,\gamma_{-(i,j)},\mu_{-(i,j)},\tau,\phi)$ is easy to evaluate (up to a normalizing constant) since the simplex constraint implies $\mu_{i}=1-\sum_{k\neq i,j}\mu_{k}$. The case $\gamma_{i}=0,\gamma_{j}=1$ can be analyzed similarly. When $\gamma_{i}=\gamma_{j}=1$, we first find the full conditional posterior $p(\mu_{i},\mu_{j}\mid y,\gamma_{-(i,j)},\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)$, which is a degenerate truncated normal distribution due to the constraint $\mu_{i}>0,\mu_{j}>0$ and $\mu_{i}+\mu_{j}=1-\sum_{k\neq i,j}$. This is derived in Lemma S1. Integrating out $(\mu_{i},\mu_{j})$, we get the conditional posterior probability $p(\gamma_{i}=\gamma_{j}=1\mid y,\gamma_{-(i,j)},\mu_{-(i,j)},\tau,\phi)$ up to a normalizing constant. Normalizing over the three cases yields the posterior distribution $p(\gamma_{i},\gamma_{j}\mid y,\gamma_{-(i,j)},\mu_{-(i,j)},\tau,\phi)$, which has a closed-form expression in terms of the cumulative distribution function of the normal distribution; see Lemma S2. The resulting sampling algorithm is summarized in Algorithm 1, where $\beta_{i,j},\Lambda_{i,j}$ are defined in Lemma S1. More details are provided in Appendix S1.
Input: data set $(X,Y)$, hyperparameters $\kappa_{1},\kappa_{2},a_{1},a_{2},\alpha,\theta$, number of iterations $n$, algorithm parameters $n_{\tau}\in\mathbb{N},\eta>0$, initial state $(\mu^{(0)},\tau^{(0)},\phi^{(0)})$.
for $t=1,\dots,n$ do
   Set $\mu\leftarrow\mu^{(t-1)}$ ;
 for $i=1,\dots,N-1$ do
    for $j=i+1,\dots,N$ do
         Calculate $s\leftarrow 1-\sum_{k\neq i,j}\mu_{k}$ ;
       if $s=0$ then
            Set $\mu_{i}\leftarrow 0$, $\mu_{j}\leftarrow 0$;
       else
            Draw $(\gamma_{i},\gamma_{j})$ with probability $p(\gamma_{i},\gamma_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$ by Lemma S2;
          if $\gamma_{i}=1,\gamma_{j}=0$ then
               Set $\mu_{i}\leftarrow s$, $\mu_{j}\leftarrow 0$;
          else if $\gamma_{i}=0,\gamma_{j}=1$ then
               Set $\mu_{i}\leftarrow 0$, $\mu_{j}\leftarrow s$;
          else
               Draw $u\sim N_{(0,s)}\left(\beta_{i,j},(\phi\Lambda_{i,j})^{-1}\right)$ by Lemma S1;
               Set $\mu_{i}\leftarrow u$, $\mu_{j}\leftarrow s-u$;
  Set $\mu^{(t)}\leftarrow\mu$;
   Draw $\phi^{(t)}$ from $p(\phi\mid y,\mu^{(k)},\tau^{(k-1)})$ by (10);
   Set $\tau\leftarrow\tau^{(t-1)}$;
 for $i=1,\dots,n_{\tau}$ do
      Propose $\tau^{*}$ by $\log\tau^{*}=\log\tau+N(0,\eta)$;
      Set $\tau\leftarrow\tau^{*}$ with acceptance probability $\rho(\tau,\tau^{*})$ given by equation (S1);
  Set $\tau^{(t)}\leftarrow\tau$;
   Draw $w^{(t)}$ from $p(w\mid y,\mu^{(t)},\tau^{(t)},\phi^{(t)})$ by (11);
Output: samples $(\mu^{(t)},\tau^{(t)},\phi^{(t)},w^{(t)})_{t=1}^{n}$
Algorithm 1 Metropolis-within-Gibbs sampling for BVS-SS
4 Theoretical Results for BVS-SS
In this section, we study the theoretical properties of BVS-SS. For ease of analysis, we fix the hyperparameter $\tau$ and consider two complementary scenarios: $\tau=0$ and $\tau\rightarrow\infty$. Following the convention in the variable selection literature (yang2016computational), we interpret a model $\gamma$ as both a binary vector and a subset of $N]:=\{1,2,\dots,N\}$; that is, to indicate that the $i$-th covariate is selected, we can write either $\gamma_{i}=1$ or $i\in\gamma$. We use $|\cdot|$ to denote set cardinality, and thus $|\gamma|$ denotes the model size.
4.1 High-dimensional Selection Consistency with $\tau=0$
For the first scenario, we fix the prior parameter $\tau=0$ and assume that the true DGP satisfies the simplex constraint. Our goal is to establish, in high-dimensional regimes, the “strong selection consistency” property of BVS-SS which means that the posterior probability of the true model, denoted by $\gamma^{*}$, converges to one in probability under the true DGP; here, “strong” means that this property is stronger than other consistency criteria such as pairwise selection consistency (casella2009consistency). Since $\tau=0$ implies $w=\mu$ almost surely, we can rewrite the model as
$Y=X\mu+\epsilon,\quad\epsilon\sim N(0,\phi^{-1}I).$ (14)
For technical convenience, we slightly modify the prior specification as follows:
$p(\gamma)$ $\propto\Gamma(|\gamma|)^{-1}\left\{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}\right\}^{1/2}\det(X_{\gamma}^{\top}X_{\gamma})^{1/2}\theta^{|\gamma|}\,(1-\theta)^{N-|\gamma|}\mathbbm{1}_{\mathbb{S}_{L}}(\gamma),$ (15)
$p(\phi\mid\gamma)$ $\propto\phi^{(\kappa_{1}+|\gamma|-1)/2-1}e^{-\phi\kappa_{2}/2},$ (16)
$\mu_{\gamma}\mid\gamma$ $\sim\text{Dir}(\mathsf{1}),$ (17)
[…]
Lemma 1.
Let $\gamma^{\prime}=\gamma\cup\{j\}$ for some $j\notin\gamma$, and $p(\gamma)$ be given by (15). Then,
$\frac{\sqrt{M\underline{\lambda}}}{L}\leq\frac{p(\gamma^{\prime})/p(\gamma)}{\theta/(1-\theta)}\leq\sqrt{\frac{LM}{\underline{\lambda}}}.$
Proof.
See Appendix S3.1. ∎
By integrating out $\mu_{\gamma}$, we can express the posterior density of $(\gamma,\phi)$ as
$p(\gamma,\phi\mid y)$ $\propto p(\gamma)p(\phi\mid\gamma)\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\mathsf{1})}\left[p(y\mid\gamma,\mu_{\gamma},\phi)\right],$ (18)
$\text{ where }p(y\mid\gamma,\mu_{\gamma},\phi)$ $\propto\phi^{M/2}\,\exp\left\{-\frac{\phi}{2}\|y-X_{\gamma}\mu_{\gamma}\|_{2}^{2}\right\}.$ (19)
Unlike the unconstrained Bayesian linear regression, $p(\gamma,\phi\mid y)$ is not available in closed form due to the intractable integral involved in $p(y\mid\gamma,\phi)$. Consequently, the marginal posterior probability of $\gamma$, $p(\gamma\mid y)\propto p(\gamma)\int p(\phi\mid\gamma)p(y\mid\gamma,\phi)\mathrm{d}\phi$, has no closed-form expression either. To establish the high-dimensional consistency result, we make the following assumptions on the design matrix, prior parameters and true DGP, where all parameters involved may vary with the sample size $M$, except the universal constants denoted by $c_{\mu},c_{\theta},c_{M}$ and $c_{j}$ for integer $j$. In particular, the number of variables $N$, the maximum model size $L$, and the true model size $\ell^{*}=|\gamma^{*}|$ are all allowed to go to infinity with $M$.
  1. (A1)
    The design matrix $X$ satisfies that $\|X_{j}\|_{2}^{2}=M$ for each $j\in[N]$, and there exists a constant $\underline{\lambda}\in(0,1]$ such that $\lambda_{\mathrm{min}}(X^{\top}_{\gamma}X_{\gamma})\geq M\underline{\lambda}$ for all $\gamma\in\mathbb{S}_{L}$, where $\lambda_{\mathrm{min}}$ denotes the smallest eigenvalue.
  2. (A2)
    The prior distribution on $\gamma$ is given by (15) where $\theta$ satisfies $\theta/(1-\theta)=N^{-c_{\theta}\,L}$ for some universal constant $c_{\theta}>0$.
  3. (A3)
    The prior distribution on $\phi$ is given by (16) where hyperparameters $\kappa_{1},\kappa_{2}$ satisfy $0<\kappa_{1}\leq M$, and $0\leq\kappa_{2}\leq\sigma^{2}M/2$.
  4. (A4)
    The true DGP is given by $Y\mid X\sim\mathcal{N}(X_{\gamma^{*}}\mu^{*}_{\gamma^{*}},\,\sigma^{2}I_{M})$ where $\sigma>0$, $\gamma^{*}$ and $\mu^{*}_{\gamma^{*}}$ satisfy the following conditions: $\ell^{*}\coloneqq|\gamma^{*}|\leq L\wedge\sqrt{L\log N}$, $\mu^{*}_{\gamma^{*}}\in\Delta^{{\ell^{*}}-1}$, and
    $\min_{j\in\gamma^{*}}|\mu^{*}_{j}|\geq\frac{c_{\mu}\sigma\sqrt{L\log N}}{\underline{\lambda}\sqrt{M}},$ (20)
    for some universal constant $c_{\mu}>0$.
  5. (A5)
    $L\geq 3$ and $M\geq c_{M}\ell^{*}L\log N$ for some universal constant $c_{M}>0$.
Assumption (A1) requires that each column of $X$ has the same magnitude and that the minimum eigenvalue of the Gram matrix is well controlled for any model of size at most $L$. It is a mild condition, often known as “restricted eigenvalue,” commonly used in the literature (shang2011consistency, narisetty2014bayesian, yang2016computational). Assumption (A2) provides guidance for choosing the prior: the prior inclusion probability $\theta$ should be small enough in order to penalize large models and avoid overfitting. Similarly, Assumption (A3) requires that the choices of $\kappa_{1},\kappa_{2}$ are not too extreme. Assumption 20 assumes that the true DGP satisfies the simplex constraint, and (20) imposes a lower bound on the smallest nonzero coefficient of the true DGP. This is also a standard assumption and often known as the $\beta$-min condition.” Intuitively, this assumption is needed since coefficients close to zero cannot be consistently identified through variable selection. Note that the order of the detection threshold required by (20) is essentially the same as that in the unconstrained Bayesian variable selection (yang2016computational). Finally, Assumption (A5) specifies the sample size needed for consistent variable selection. We use $\mathbb{P}^{*}$ to denote the probability measure associated with the true DGP specified in Assumption 20.
Under the above assumptions, we prove that $p(\gamma^{*}\mid y)$ converges to one in probability with respect to $\mathbb{P}^{*}$ given sufficiently large $c_{\mu},c_{\theta}$ and $c_{M}$. A key step in our proof is to establish the following posterior probability ratio bound.
Theorem 1.
Suppose Assumptions (A1) to (A5) hold with $c_{M}\geq 4$ and $c_{\mu}\geq 6$. With probability at least $1-c_{1}N^{-c_{2}}$ for some universal constants $c_{1},c_{2}>0$, the following bound holds for all $\gamma\in\mathbb{S}_{L}$:
$\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq 3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{|\gamma|-\ell^{*}}\left(\frac{\kappa_{2}+\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}+\sigma^{2}\ell^{*}\log N}{\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}}\right)^{(\kappa_{1}+M)/2},$ (21)
where $\hat{\mu}_{\gamma}=(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}y$.
Proof sketch.
The complete proof of Theorem 1 is presented in Appendix S3.2. Here we provide a sketch of the proof to highlight the key challenges and proof techniques. As shown in (19), the marginal likelihood given $(\gamma,\phi)$ can be expressed as $p(y\mid\gamma,\phi)=\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\mathsf{1})}\left[p(y\mid\gamma,\mu_{\gamma},\phi)\right]$, where the expectation is an integral over the simplex. To analyze this intractable integral, we first show that, for any $\mu_{\gamma}\in\Delta^{|\gamma|-1}$, the likelihood $p(y\mid\gamma,\mu_{\gamma},\phi)$ can be expressed as
$p(y\mid\gamma,\mu_{\gamma},\phi)$ $\propto\phi^{M/2}\,\exp\left\{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}+\|X_{\gamma}(\mu_{\gamma}-\check{\mu}_{\gamma})\|_{2}^{2}\right)\right\},$ (22)
$\text{ where }\check{\mu}_{\gamma}$ $=\hat{\mu}_{\gamma}+\frac{1-\mathsf{1}^{\top}\hat{\mu}_{\gamma}}{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1},\quad b_{\gamma}=\frac{(1-\mathsf{1}^{\top}\hat{\mu}_{\gamma})^{2}}{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}}.$ (23)
That is, we decompose the residual sums of squares $\|y-X_{\gamma}\mu_{\gamma}\|_{2}^{2}$ into three parts. The first component, $\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}$, is the residual sum of squares for the least-squares estimator in the unconstrained setting. The second component, $b_{\gamma}\geq 0$, measures the deviation of $\hat{\mu}_{\gamma}$ from the simplex constraint. When $\hat{\mu}_{\gamma}\in\Delta^{|\gamma|-1}$, we have $b_{\gamma}=0$. Note that the first two components do not depend on $\mu_{\gamma}$. The last component, $\|X_{\gamma}(\mu_{\gamma}-\check{\mu}_{\gamma})\|_{2}^{2}$, measures the deviation of $\mu_{\gamma}$ from $\check{\mu}_{\gamma}$, where $\check{\mu}_{\gamma}$ is the least-squares estimator for model $\gamma$ under the linear constraint $\mathsf{1}^{\top}\check{\mu}_{\gamma}=1$. It has the following interpretation: $X_{\gamma}\check{\mu}_{\gamma}$ is the projection of $X_{\gamma}\hat{\mu}_{\gamma}$ onto the affine space $\{X_{\gamma}u\colon\mathsf{1}^{\top}u=1\}$. See Appendix S2 for a brief review on the constrained least-squares estimation.
In Proposition S1, we use (22) and the modified priors given in (15) and (16) to show that
$p(\gamma,\phi\mid y)\propto f(|\gamma|)\,\phi^{(M+\kappa_{1})/2}\,e^{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}+\kappa_{2}\right)}\,\mathbb{P}\left(\tilde{U}_{\gamma,\phi}\in{\tilde{\Delta}^{|\gamma|-1}}\right),$ (24)
where $f(|\gamma|)$ is a function only depending on $|\gamma|$ and prior parameter $\theta$, $\tilde{U}_{\gamma,\phi}$ is a normal random vector whose mean equals the subvector of $\check{\mu}_{\gamma}$ consisting of its first $|\gamma|-1$ entries, and $\tilde{\Delta}^{\ell}=\{\tilde{u}\in\mathbb{R}^{\ell}\colon\tilde{u}_{i}\geq 0\text{ for each $i$, and }\sum\nolimits_{i=1}^{\ell}\tilde{u}_{i}\leq 1\}.$ For any $\gamma\in\mathbb{S}_{L}$, we have $\mathbb{P}(\tilde{U}_{\gamma,\phi}\in{\tilde{\Delta}^{|\gamma|-1}})\leq 1$ and $b_{\gamma}\geq 0$. Hence, by integrating over $\phi$, we get an upper bound on $p(\gamma\mid y)$ that holds without any additional assumption; see Proposition S2.
To find a lower bound on $p(\gamma^{*},\phi\mid y)$, we need to find upper bounds on $\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}$ and $b_{\gamma^{*}}$ and a lower bound on $\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}})$. To this end, we construct some high-probability events in Lemma S5 to ensure that the error vector $\epsilon$ is well-behaved. This argument is standard in the high-dimensional literature and ensures that $\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}=O_{p}(\sigma^{2}M)$. For $b_{\gamma^{*}}$, we expect that it cannot be too large under our assumptions, as it measures how well $\hat{\mu}_{\gamma^{*}}$ satisfies the simplex constraint. In Lemma S6, we prove that $b_{\gamma^{*}}=O_{p}(\sigma^{2}\ell^{*}\log N)$. Deriving a lower bound on $\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}})$ is considerably more involved. We invoke the beta-min condition given in Assumption 20, which ensures that the true regression coefficients cannot be too small, to show that $\check{\mu}_{\gamma^{*}}$ lies in the simplex $\Delta^{\ell^{*}-1}$. Then, in Lemma S8, we apply standard Gaussian concentration inequalities to get
$1-\mathbb{P}\left(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}}\right)\leq\exp\left(-\frac{4\phi\,\sigma^{2}L\log N}{(\ell^{*})^{2}}+\frac{1}{2}\right).$ (25)
We visualize the intuition behind the proof of (25) in Figure 1. The form of the bound (25) enables us to integrate over $\phi$, and by using Assumptions (A3) and (A5), we obtain a lower bound on $p(\gamma^{*}\mid y)$; see Proposition S3. Theorem 1 then follows by combining Propositions S2 and S3. ∎
Figure 1: Intuition for inequality (25). Here $\ell^{*}=4$, the tetrahedron represents $\tilde{\Delta}^{3}$, and $\Pi\check{\mu}_{\gamma^{*}}$ denotes the mean of $\tilde{U}_{\gamma^{*},\phi}$. We show that $\Pi\check{\mu}_{\gamma^{*}}\in\tilde{\Delta}^{3}$ and use concentration inequality to bound the probability mass of a ball centered at $\Pi\check{\mu}_{\gamma^{*}}$ and contained in $\tilde{\Delta}^{3}$.
By Theorem 1, we can bound $p(\gamma\mid y)/p(\gamma^{*}\mid y)$ by analyzing the behavior of unconstrained least-squares estimators, which is well understood in the variable selection literature. This analysis then yields the desired strong selection consistency.
Theorem 2.
Suppose Assumptions (A1) to (A5) hold for sufficiently large $c_{\theta},c_{\mu},c_{M}$. Then, $\mathbb{P}^{*}\{p(\gamma^{*}\mid y)\geq 1-c_{3}N^{-c_{4}L}\}\geq 1-c_{1}N^{-c_{2}}$ for some universal constants $c_{1},c_{2},c_{3},c_{4}>0$.
Proof.
See Appendix S3.4. ∎
The strong selection consistency ensures that the true model $\gamma^{*}$ can be identified with high probability. Next, we consider the estimation of the regression coefficient vector $\mu$ and the task of prediction. The posterior expected $L^{2}$ loss for estimating $\mu$ can be expressed as
$\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]=\sum_{\gamma\in\mathbb{S}_{L}}p(\gamma\mid y)\left(\|\mu^{*}_{\gamma^{c}}\|_{2}^{2}+\int\|\mu_{\gamma}-\mu^{*}_{\gamma}\|_{2}^{2}\;p(\mu_{\gamma},\phi\mid y,\gamma)\mathrm{d}\mu_{\gamma}\mathrm{d}\phi\right),$ (26)
where we recall $\mathbb{E}_{y}$ denotes the expectation over the posterior distribution given $Y=y$. Similarly, given a new design matrix $\tilde{X}$, we can define the posterior expected predictive loss by $\mathbb{E}_{y}[\|\tilde{X}\mu-\tilde{X}\mu^{*}\|_{2}^{2}]$. In Theorem 3, we characterize the rate at which the two loss functions go to zero.
Theorem 3.
Under the setting of Theorem 2,
$\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]=O_{p}\left(\max\left\{\frac{\sigma^{2}L\log N}{\underline{\lambda}^{2}M},\;N^{-L/(\ell^{*})^{2}}\right\}\right).$ (27)
Let $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ be such that $\max_{\gamma\in\mathbb{S}_{L}}\lambda_{\mathrm{max}}(\tilde{X}^{\top}_{\gamma}\tilde{X}_{\gamma})\leq\tilde{M}\overline{\lambda}$ for some $\overline{\lambda}>0$. Then,
$\mathbb{E}_{y}\left[\|\tilde{X}\mu-\tilde{X}\mu^{*}\|_{2}^{2}\right]=O_{p}\left(\tilde{M}\overline{\lambda}\,\max\left\{\frac{\sigma^{2}L\log N}{\underline{\lambda}^{2}M},\;N^{-L/(\ell^{*})^{2}}\right\}\right).$ (28)
Proof.
See Appendix S3.5. ∎
The consistency of our ATT estimator follows from the the prediction loss bound in Theorem 3. We expect that this result can be extended to more general scenarios, including cases where $X^{\top}\;\tilde{X}^{\top}$] is a stationary process (li2020statistical).
4.2 Limiting Behavior when $\tau\rightarrow\infty$
We also investigate the performance of BVS-SS when $\tau\to\infty$. In this case, the prior places increasing mass on values of $w$ that do not satisfy the simplex constraint. We compare the behavior of the posterior distribution $p(\gamma\mid y)$ with the conventional unconstrained version, which we denote by $\tilde{p}(\gamma\mid y)$ (see Appendix S3.6 for details). The following theorem shows that the two posterior distributions become essentially the same as $\tau\rightarrow\infty$.
Theorem 4.
Let $\tilde{p}(\cdot\mid y,\tau)$ be the posterior distribution given $\tau$ under the unconstrained spike-and-slab prior detailed in Appendix S3.6. Then $\lim_{\tau\rightarrow\infty}p(\gamma\mid y,\tau)/\tilde{p}(\gamma\mid y,\tau)=1.$
Proof.
See Appendix S3.6. ∎
In Theorem 4, we treat the data $(X,y)$ as fixed and let $\tau\rightarrow\infty$, but we expect that our argument can be extended to a fully non-asymptotic analysis. Such an extension may be combined with existing variable selection consistency results to show that our model also achieves selection consistency when the true DGP significantly violates the simplex constraint but $\tau$ is sufficiently large.
5 Simulation Studies
5.1 Simulation Settings
We simulate the observed training data $X\in\mathbb{R}^{M\times N}$ and $Y\in\mathbb{R}^{M}$ by $Y=Xw^{*}+\epsilon,$ where $\epsilon_{i}\overset{\mathrm{i.i.d.}}{\sim}N(0,1/\phi^{*})$, and $w^{*}\in\mathbb{R}^{N}$ is the true regression coefficient vector such that $w^{*}_{j}=0$ for $j=J+1,\dots,N$ (that is, only the first $J$ predictors have nonzero effects on $Y$). Similarly, we simulate the data under treatment $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ and $\tilde{Y}^{(1)}\in\mathbb{R}^{\tilde{M}}$ by
$\tilde{Y}^{(1)}=\;$ $\tilde{X}w^{*}+\delta+\tilde{\epsilon},\text{ where }\delta_{i}\overset{\mathrm{i.i.d.}}{\sim}N(\delta^{*},1/\nu^{*}),\;\tilde{\epsilon}_{i}\overset{\mathrm{i.i.d.}}{\sim}N(0,1/\phi^{*}).$ (29)
The vector $\delta\in\mathbb{R}^{\tilde{M}}$ represents the treatment effects, and $\delta^{*}\in\mathbb{R}$ denotes the true ATT. In our simulation, we use $M=\tilde{M}\in\{25,50,100,200\}$ and $N\in\{20,50\}$. When $N=20$, we set $J=5$, and when $N=50$, we set $J=10$. Given $J$, we define a vector $\mu^{*}\in\Delta^{N-1}$ by
$\mu^{*}_{j}=\left\{\begin{array}[]{cc}j/S_{J}&\text{ if }1\leq j\leq J,\\ 0,&\text{ if }J+1\leq j\leq N,\end{array}\right.\quad\text{ where }S_{J}=\frac{1}{2}J(J+1).$ (30)
Then, we define $w^{*}$ by $w^{*}=\lambda\mu^{*}$ where $\lambda>0$ is a scaling factor. When $\lambda=1$, the true DGP satisfies the simplex constraint. Assuming $\lambda>0$, we have $\lambda=\|w^{*}\|_{1}$. We consider $\lambda\in\{1,2,3\}$ in our simulation studies. Given $w^{*}$, we set the variance parameters by $v^{*}=\phi^{*}=4/\|w^{*}\|_{2}^{2}$. For each choice of $(M,N,\lambda)$, we generate 100 replicates where entries of $X,\tilde{X}$ are also sampled independently from the standard normal distribution. See Appendix S4 for simulation results where the true DGP is driven by a factor model. We fix true ATT $\delta^{*}=0.5$ throughout our simulation studies.
We set the hyperparameters of BVS-SS by $\kappa_{1}=\kappa_{2}=1,a_{1}=0.01,a_{2}=0.1,\alpha=1,\theta=0.2.$ The choice of $(a_{1},a_{2})$ guarantees that $\tau$ has prior mean $0.1$ and prior variance $1$. Further, for numerical stability, we truncate the prior distribution of $\tau$ by assuming that $\tau\geq 10^{-6}$. For every simulated data set, we initialize Algorithm 1 at $\tau^{(0)}=\phi^{(0)}=1$ and $\mu^{(0)}$ sampled from its prior distribution, and we run Algorithm 1 for $1,000$ iterations and discard the first $500$ iterations as burn-in.
For comparison, we have also implemented the following methods for ATT estimation. First, we use the R package glmnet to perform variable selection and ATT estimation via Lasso; the parameter of Lasso is chosen by cross-validation. Second, we compute the OLS estimator for $w^{*}$ using the entire data set $(X,Y)$ without variable selection, denoted by $\hat{w}_{\mathrm{LS}}$. Then ATT is estimated by
$\widehat{\mathrm{ATT}}_{\mathrm{LS}}=\frac{1}{\tilde{M}}\sum\nolimits_{i=1}^{\tilde{M}}\left(\tilde{Y}^{(1)}_{i}-\hat{w}_{\mathrm{LS}}^{\top}\tilde{X}_{(i)}\right),$ (31)
where $\tilde{X}_{(i)}$ is the $i$-th row of $\tilde{X}$ treated as a column vector. We also compute the simplex-constrained estimator $\hat{w}_{\mathrm{QP}}$ (QP stands for “quadratic programming”) defined by $\hat{w}_{\mathrm{QP}}=\operatorname*{argmin}_{w\in\Delta^{N-1}}\|Y-Xw\|_{2}^{2},$ which was introduced in abadie2003economic and can be obtained by quadratic programming (goldfarb2006dual, abadie2011synth). The corresponding ATT estimator $\widehat{\mathrm{ATT}}_{\mathrm{QP}}$ is obtained similarly by replacing $\hat{w}_{\mathrm{LS}}$ with $\hat{w}_{\mathrm{QP}}$ in (31). Finally, we consider two oracle estimators that are not implementable in practice. Denote by $\gamma^{*}=\gamma(w^{*})$ the true indicator vector corresponding to $w^{*}$. We compute the OLS and QP estimators using $(X_{\gamma^{*}},Y)$. That is, these two estimators have access to $\gamma^{*}$ (which is not possible in reality) and will be used for comparison in our simulation studies. Note that since the regression coefficient estimator of Lasso and OLS are scale-equivariant, the ATT estimators of Lasso and two OLS methods does not depend on the signal size factor $\lambda$.
5.2 Estimation of Average Treatment Effect
[…]
$\mathrm{RE}(\widehat{\mathrm{ATT}})=\frac{\mathrm{MSE}(\widehat{\mathrm{ATT}}_{\mathrm{OLS}(\gamma^{*})})}{\mathrm{MSE}(\widehat{\mathrm{ATT}})}$. (32)
If $\mathrm{RE}(\widehat{\mathrm{ATT}})=100\%$, then this estimator is as efficient as the oracle OLS estimator.
[…]
Figure 2: Relative efficiency (%) vs. sample size for ATT estimation. BVS-$S$ and QP($\gamma^{*}$) are in black; Lasso, OLS and QP in gray. OLS($\gamma^{*}$) always has RE=100% (horizontal line).
[…]
5.3 Variable Selection and Estimation of Variance Parameters
We now examine the performance of BVS-SS in terms of variable selection and the estimation of the two variance parameters $\tau,\phi$. First, we compare the variable selection result of BVS-SS with that of Lasso in Table 1, where $\ell^{1}$-loss” denotes the $\ell^{1}$-distance between the selected model $\hat{\gamma}$ and the true model $\gamma^{*}$, and “model size” refers to the cardinality of $\hat{\gamma}$ (i.e., the number of nonzero entries of $\hat{w}$). For BVS-SS, the two statistics are averaged over the collected MCMC samples. It is evident that BVS-SS has a much higher accuracy than Lasso across all settings. Especially when $M$ is large, BVS-SS can correctly identify most entries of $\gamma$, while the performance of Lasso remains similar to that observed with small $M$.
Next, we visualize the distribution of the posterior mean estimate of $\phi$ (error precision) across 100 replicates in Figure 3. This statistic reflects how well the BVS-SS model fits the data. We only show the result for $N=50$, as the distribution for $N=20$ is similar. Observe that regardless of $\|w^{*}\|_{1}$, the posterior mean of $\phi$ increases with $M$, which is expected since a larger sample size enables BVS-SS to detect more signals in the data, resulting in a smaller estimate of the error variance. Notably, when $\|w^{*}\|_{1}=2$ or $3$, the posterior estimation of $\phi$ appears highly accurate for $M\geq 100$, indicating that our model achieves an optimal balance between model complexity and fitting.
BVS-$S$ Lasso
$\|w^{*}\|_{1}$ Setting $M$ $\ell^{1}$-loss model size $\ell^{1}$-loss model size
$\|w^{*}\|_{1}$ = 1
$N=20$
$J=5$
25 3.1 3.9 6.3 10.2
50 1.9 4.8 6.7 11.1
100 1.3 5.1 6 10.9
200 0.8 5.4 6.3 11.3
$N=50$
$J=10$
25 10.9 6.2 13 14.2
50 7.8 9.1 15.4 22.2
100 5.5 10.9 14.4 22.9
200 4.1 11.4 13.4 22.5
$\|w^{*}\|_{1}$ = 3
$N=20$
$J=5$
25 3.3 3.3 6.3 10.2
50 1.9 4.8 6.7 11.1
100 1.2 5 6 10.9
200 0.7 5.2 6.3 11.3
$N=50$
$J=10$
25 10.2 4.5 13 14.2
50 8.1 9.2 15.4 22.2
100 5.6 11.2 14.4 22.9
200 3.7 10.8 13.4 22.5
Table 1: Variable selection performance of BVS-SS and Lasso averaged over 100 replicates.
Figure 3: Distribution of the posterior mean of $\phi$ across $100$ replicates with $N=50,J=10$. The true value $\phi^{*}$ is indicated by the dotted line.
Finally, we examine the posterior mean estimate of $\tau$ when $N=50$, for which the result is presented in Figure 4. As discussed in Section 2.1, $\tau$ can be considered as an indicator of whether the simplex constraint is satisfied by the given data, and we now explain why Figure 4 provides compelling empirical evidence supporting this. When $\|w^{*}\|_{1}=1$, we see that the posterior mean of $\tau$ decreases as $M$ increases, since a larger sample size offers stronger evidence that the simplex constraint is satisfied, which draws $\tau$ towards zero. In contrast, when $\|w^{*}\|_{1}=2$ or $3$, the trend reverses, since a larger sample size enables BVS-SS to detect deviations from the simplex constraint. When $M=200$, we find that the ratio of the posterior mean estimate of $\tau$ and that of $\phi$ equals $0.0003$ for $\|w^{*}\|_{1}=1$, $0.024$ for $\|w^{*}\|_{1}=2$, and $0.081$ for $\|w^{*}\|_{1}=3$. This actually aligns well with the true data-generating mechanism: since we use $J=10$ for $N=50$, we may estimate the “true” mean squared deviation of $w^{*}$ from the simplex constraint by $0.1\times\sum_{j=1}^{10}(w^{*}_{j}-0.1)^{2}$, which equals $0.021$ for $\|w^{*}\|_{1}=2$ and $0.065$ for $\|w^{*}\|_{1}=3$.
Figure 4: Distribution of the posterior mean of $\tau$ across $100$ replicates with $N=50,J=10$.
6 Empirical Examples
In this section, we revisit two empirical examples from the SCM literature. We will use the notation introduced in Section 1.1: $T_{0}$ is the number of pre-treatment time periods, $T_{1}$ is the number of post-treatment time periods, the total number of units is $N+1$ where the first one is used as the response and the other $N$ units as the explanatory variables. (For BVS-SS, $M=T_{0}$ and $\tilde{M}=T_{1}$.) The first example has $N<T_{0}$, while the second has $N\gg T_{0}$, and it will be shown that our method works well in both settings.
6.1 Nota Fiscal Paulista: Anti-tax Evasion
An anti-tax evasion program, Nota Fiscal Paulista (NFP), was implemented in São Paulo, Brazil in October 2007 with the goal of reducing tax evasion by incentivizing consumers to request electronic receipts in exchange for potential tax rebates and participation in monthly lotteries promoted by the government. However, those restaurant owners and retailers may pass on part of the tax costs to consumers by raising product prices. Although the NFP was implemented sequentially across various sectors, including restaurants, bakeries, bars, and food service retailers, the literature has suggested that the food away from home (FAH) index could be used as a suitable indicator for price levels of these sectors. carvalho2018arco proposed an artificial counterfactual (ArCo) estimation procedure, by connecting Lasso to SCM without simplex constraint, to explore how the NFP affects the inflation on FAH. They also included monthly GDP growth, retail sales growth and monthly credit growth. Their ArCo estimation indicated that the ATT of NFP is 0.4478%.
We obtain the data analyzed in carvalho2018arco from the R package ArCo. The pre-treatment periods span $T_{0}=33$ months, and the post-treatment periods span $T_{1}=23$ months. Slightly different from the original study, the data includes only $N=8$ control units, which is one fewer than that mentioned in their paper, and there is no data related to credit and retail sales. We run our Algorithm 1 for 1000 iterations to estimate the impact of NFP, with the first half of the samples dropped for burn-in. We set the hyperparameters by $\kappa_{1}=\kappa_{2}=1,a_{1}=0.01,a_{2}=0.1,\alpha=1,\theta=0.25.$ Table 2 summarizes the posterior mean and 95% credible intervals of ATT, $\tau$, $\phi$ and the model size $|\gamma|$. The ATT estimation is close to that of carvalho2018arco, but our credible interval is much smaller compared to their confidence set. While the analysis of carvalho2018arco selected all the predictors to construct the synthetic control, our method only chooses about two predictors on average. This is consistent with the findings of our simulation study presented in Section 5, which shows that Lasso has the tendency of over-estimating the model size compared to BVS-SS. In Appendix S5.1, we provide the trace plots for $\phi$ and $\log(\tau)$ and visualize the posterior distributions of ATT, $\log(\tau)$, $\phi$, model size $|\gamma|$, and the counterfactual estimate are provided.
ATT $\tau$ $\phi$ $|\gamma|$
Mean 0.288 0.14 3.85 2.29
95% credible interval (0.141, 0.419) (0, 1.44) (1.94, 6.05) (1, 4)
Table 2: NFP Impact on Food Inflation
6.2 China’s Anti-corruption Campaign
shi2023forward investigated how China’s unprecedented anti-corruption campaign, launched in November 2012, affected the importation of luxury watches. This was motivated by the fact that high-end imported watches are often used as discreet bribery gifts in place of blatant cash payments, and empirical studies have shown that the imports of luxury watches in China move with government leadership changes lan2018swiss. shi2023forward used data from the United Nations Comtrade Database, taking the monthly growth rate of the category “watches with case of, or clad with, precious metal” as the outcome and constructing the synthetic counterfactual from $N=87$ other commodity categories. January 2013 is considered the time of treatment, which was the month immediately following the announcement of the Eight-Point anti-corruption policy in China. The pre-treatment period spans from February 2010 to December 2012, yielding $T_{0}=35$ observations, and the post-treatment period covers January 2013 to December 2015, yielding $T_{1}=36$ observations. Their forward selection algorithm identified three categories as control units: “knitted or crocheted fabric”, “cork and articles of cork”, and “salt, sulfur, earth, stone, plaster, lime, and cement”. Their estimated counterfactual predicted that without the anti-corruption campaign, luxury watch imports would have continued to increase by 1.7% in January 2013, while in reality imports dropped by 42%. Their treatment effect estimation indicated a reduction of 3.09% in luxury watch imports per month over the post-treatment period.
We obtain the data of shi2023forward from the R package fdPDA, which contains all the 88 commodity categories (including the response and $N=87$ control categories), and run our Algorithm 1 for 1000 iterations with the same hyperparameters as in the last example. Table 3 indicates that our ATT estimation (-2.1%) is smaller than that of shi2023forward in absolute value, but the 95% credible interval indicates that the anti-corruption has a significant effect on the high-end watch imports. On average, our model size $|\gamma|$ is larger than that of shi2023forward (who only selected 3 control categories), but Figure S11 shows that the posterior distribution of $|\gamma|$ still concentrates on small models with size $\leq 5$. Additional simulation results are provided in Appendix S5.2.
ATT $\tau$ $\phi$ $|\gamma|$
Mean -0.021 0.069 20.86 5.09
95% credible interval (-0.032, -0.008) (0, 0.641) (12.22, 32.76) (1, 20)
Table 3: Anti-corruption Campaign’s Impact on Luxury Watches Importation
It is worth noting that in our posterior samples, the categories “cork” and “salt” appear with negligible probabilities. Instead, the most frequently selected categories are those related to textiles—for example, “knitted or crocheted fabrics” (51.4%), “special woven fabrics such as lace, tapestries, trimmings, and embroidery” (16.6%), and “wool, fine or coarse animal hair” (12%). In addition, the category “optical, photographic, cinematographic, measuring, checking, medical or surgical instruments and apparatus; parts and accessories” also appears with a relatively high probability (10.2%). Compared with “cork” and “salt”, these categories seem more plausible to serve as control units. The textile products share decorative features similar to those of wristwatches, while the optical and photographic instruments also rely on precision mechanical engineering.
7 Conclusion and Future Directions
We propose a novel Bayesian synthetic control method, BVS-SS, that integrates a relaxed simplex constraint with Bayesian spike-and-slab variable selection. Our approach introduces a hierarchical prior to determine how closely the data should adhere to the simplex constraint, a feature that is missing in existing frequentist or Bayesian synthetic control frameworks and helps mitigate the ongoing debate surrounding the use of simplex constraints. To efficiently compute the posterior distribution of BVS-SS, we develop a novel Metropolis-within-Gibbs sampler, which overcomes the hindrance caused by the simplex constraint by updating two regression coefficients simultaneously from the full conditional posterior. We also show that in high-dimensional settings, when the true DGP satisfies the simplex constraint, BVS-SS with $\tau=0$ consistently selects the true control units and achieves consistency in terms of both regression coefficient estimation and prediction. Simulation studies and real-data examples illustrate the advantages and usefulness of our proposed algorithm.
There are several interesting directions that future researchers may dive in. First, it is known that Gibbs sampling methods are susceptible to high collinearity among the explanatory variables, which is a common situation in SCM applications where factor models are often used to model the DGP. Though it is unclear if such issues affect the methodology proposed in this work (as our algorithm updates two coordinates at a time), it would be interesting to investigate the use of other MCMC techniques, such as pseudo-marginal sampling (andrieu2009pseudo). Second, it will be interesting to further generalize BVS-SS to account for time-invariant characteristics, which are often used together with time-variant predictors in the SCM literature for constructing counterfactual outcomes. Third, more numerical studies or model extensions can be performed to study the performance of BVS-SS when the model is mis-specified (e.g., the data is non-stationary or contains heterogeneous errors), which would be of great interest to economists. We hope our method will serve as an inspiration for future research and facilitate the widespread use of Bayesian synthetic control methods.
Acknowledgements
Y.X. is grateful to Qi Li, Yonghong An and Angda Li for their helpful comments and suggestions. This work was supported in part by NSF via grants DMS-2311307 and DMS-2245591.
APPENDIX
S1 Details of the Metropolis-within-Gibbs Sampler
For Bayesian variable selection involving a prior on the indicator vector $\gamma$ as we have considered, the standard approach to posterior sampling is to integrate out the regression coefficient vector and use add-delete-swap proposals to construct a Metropolis–Hastings algorithm targeting the marginal posterior distribution of $\gamma$ (or the joint distribution of $\gamma$ and some variance parameter); see, e.g., george1997approaches, chipman2001practical, guan2011bayesian, zhou2022dimension. However, this approach is not applicable to our model, since $\mu_{\gamma}$ cannot be integrated out in closed form from the joint posterior distribution of $(\gamma,\mu_{\gamma},\tau,\phi)$. To tackle this very unique challenge, we propose a novel Metropolis-within-Gibbs sampler targeting the joint posterior distribution of $(\mu,\tau,\phi)$. Note that since $\gamma$ can be viewed as a function of $\mu$ with $\gamma_{i}=\mathbbm{1}_{\{\mu_{i}\neq 0\}}$, henceforth we will denote the conditional distribution in (7) by $p(y\mid\mu,\tau,\phi)$, and $\gamma$ should be understood as $\gamma=\gamma(\mu)$.
S1.1 Updating of $\tau$ and $\phi$
The updates for $\tau$ and $\phi$ are standard. We update $\phi$ by drawing it from the full conditional distribution given in (10), and we use a Metropolis–Hastings update for $\tau$ that is invariant with respect to $p(\tau\mid y,\mu,\phi)$. Explicitly, we propose a new value for $\tau^{*}$ by a Gaussian random walk on log-scale with variance $\eta>0$ and accept $\tau^{*}$ with probability
$\rho(\tau,\tau^{*})=\min\left\{1,\frac{p(y\mid\mu,\tau^{*},\phi)p(\tau^{*})}{p(y\mid\mu,\tau,\phi)p(\tau)}\frac{\tau^{*}}{\tau}\right\},$ (S1)
where the likelihood is given in (8) and $p(\tau)$ denotes the prior density. In Algorithm 1, we repeat this step $n_{\tau}$ times so that $\tau$ is more likely to be updated to some value with higher conditional posterior density given $\mu$ and $\phi$.
S1.2 Updating of $\mu$
For $\mu$, we propose a Gibbs-type updating scheme, which is quite different from any existing Gibbs scheme for similar problems. Let $\mu_{-j}$ denote the subvector of $\mu$ with $\mu_{j}$ removed. Since the prior of BVS-SS assigns probability one to $\mu\in\Delta^{N-1}$, given $\mu_{-j}$, we have $\mu_{j}=1-\sum_{i\neq j}\mu_{i}$ almost surely, and thus the full conditional distribution $p(\mu_{j}\mid y,\mu_{-j},\tau,\phi)$ is also degenerate. Hence, a Gibbs scheme that updates each $\mu_{j}$ from its full conditional posterior will not move at all. This motivates us to devise a more complicated updating scheme by updating $\mu_{i},\mu_{j}$ (with $i\neq j$) simultaneously from the full conditional posterior $p(\mu_{i},\mu_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$, where $\mu_{-(i,j)}$ denotes the subvector of $\mu$ with the $\mu_{i},\mu_{j}$ removed.
Let $\mu_{-(i,j)},\tau,\phi$ be given and define $s=1-\sum_{k\neq i,j}\mu_{k}$. If $s=0$, the simplex constraint implies that $\mu_{i}=\mu_{j}=0$ almost surely. If $s\in(0,1]$, the conditional posterior distribution of $\mu_{i},\mu_{j}$ falls into one of the following three cases:
  1. (i)
    given $\mu_{-(i,j)},\gamma_{i}=1,\gamma_{j}=0$, we have $\mu_{i}=s$ and $\mu_{j}=0$ a.s.;
  2. (ii)
    given $\mu_{-(i,j)},\gamma_{i}=0,\gamma_{j}=1$, we have $\mu_{i}=0$ and $\mu_{j}=s$ a.s.;
  3. (iii)
    given $\mu_{-(i,j)},\gamma_{i}=1,\gamma_{j}=1$, we have $\mu_{i}=u$ and $\mu_{j}=s-u$ for some $u\in(0,s)$ a.s.
Therefore, to find a closed-form expression for this conditional distribution, it suffices to find $p(\gamma_{i},\gamma_{j}\mid y,\mu_{-(i,j)},\tau,\phi)$ and the conditional distribution $p(\mu_{i}\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)$.
Let $\gamma_{-(i,j)}=\{k\notin\{i,j\}\colon\mu_{k}\neq 0\}\subset[N]$ denote the model corresponding to $\mu_{-(i,j)}$. We can define the four possible models under consideration by
$\gamma^{0}=\gamma_{-(i,j)},\quad\gamma^{i}=\gamma_{-(i,j)}\cup\{i\},\quad\gamma^{j}=\gamma_{-(i,j)}\cup\{j\},\quad\gamma^{ij}=\gamma_{-(i,j)}\cup\{i,j\}.$ (S2)
We will also interpret $\gamma^{0},\gamma^{i},\gamma^{j},\gamma^{ij}$ as indicator vectors in $\{0,1\}^{N}$. We first find the full conditional distribution of $(\mu_{i},\mu_{j})$ given other parameters and $\gamma_{i}=\gamma_{j}=1$ in the following lemma. For simplicity, we assume $\alpha=1$ in this section, and in Remark 1 we explain how to extend the results to other integer values of $\alpha$.
Lemma S1.
Assume $\alpha=1$. Fix $i\neq j$ such that $s=1-\sum_{k\neq i,j}\mu_{k}\in(0,1]$. The conditional distribution of $(\mu_{i},\mu_{j})\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1$ is degenerate with $\mu_{j}=s-\mu_{i}$ and
$\mu_{i}\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1\sim N_{(0,s)}\left(\beta_{i,j},\;\frac{1}{\phi\Lambda_{i,j}}\right).$ (S3)
In the above expression, $N_{(a,b)}$ denotes the univariate truncated normal distribution restricted to the interval $(a,b)$,
$\beta_{i,j}=\;$ $\frac{1}{\Lambda_{i,j}}(X_{i}-X_{j})^{\top}\Sigma_{\gamma^{ij},\tau}\left(y-sX_{j}-\sum\nolimits_{k\neq i,j}\mu_{k}X_{k}\right),$ (S4)
$\Lambda_{i,j}=\;$ $(X_{j}-X_{i})^{\top}\Sigma_{\gamma^{ij},\tau}(X_{j}-X_{i}),$ (S5)
where $\gamma^{ij}$ is given by (S2) and $\Sigma_{\gamma,\tau}$ is given by (9).
Proof.
We can rewrite the conditional posterior of $(\mu_{i},\mu_{j})$ as
$\quad p(\mu_{i},\mu_{j}\mid y,\phi,\tau,\mu_{-(i,j)},\gamma_{i}=\gamma_{j}=1)$ (S6)
$\propto p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})\,p(\mu_{i},\mu_{j}\mid\mu_{-(i,j)},\gamma_{i}=\gamma_{j}=1),$ (S7)
where the marginal likelihood $p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})$ is shown in (8). Since $s=1-\sum_{k\neq i,j}\mu_{k}$ and the symmetric Dirichlet distribution becomes a uniform distribution on the simplex $\Delta^{|\gamma|-1}$ when $\alpha=1$, we have
$p(\mu_{i}=u,\mu_{j}=s-u\mid\mu_{-(i,j)},\gamma_{i}=\gamma_{j}=1)=\;$ $(|\gamma^{ij}|-1)!,\quad\forall u\in(0,s).$ (S8)
Hence, omitting all constant terms that do not depend on $(\mu_{i},\mu_{j})$, we find that the conditional posterior density of $(\mu_{i},\mu_{j})$ is
$p(\mu_{i}=u,\mu_{j}=s-u\mid y,\phi,\tau,\mu_{-(i,j)})\propto\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{i,j},\tau}\check{y}(u)\right\}.$ (S9)
where
$\check{y}(u)=y-uX_{i}-(s-u)X_{j}-\sum\nolimits_{k\neq i,j}\mu_{k}X_{k}.$ (S10)
Using the notation $\beta_{i,j}$ and $\Lambda_{i,j}$ defined in (S4) and (S5), we obtain that
$\check{y}(u)^{\top}\Sigma_{\gamma^{i,j},\tau}\check{y}(u)$ $=\Lambda_{i,j}u^{2}-2\Lambda_{i,j}\beta_{i,j}u+\tilde{C}$
for some constant $\tilde{C}$ that does not depend on $(\mu_{i},\mu_{j})$. The claim thus follows. ∎
Next, we find the conditional posterior probabilities of $\gamma^{0},\gamma^{i},\gamma^{j}$, and $\gamma^{ij}$.
Lemma S2.
Fix $i\neq j$ such that $s=1-\sum_{k\neq i,j}\mu_{k}\in(0,1]$. Let $\check{y}(u)$ be given by (S10). The conditional posterior probabilities of $\gamma^{0},\gamma^{i},\gamma^{j},\gamma^{ij}$ defined by (S2) are given by
$p(\gamma^{0}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $0,$ (S11)
$p(\gamma^{i}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $C\,s^{\alpha-1}A_{\alpha}(\gamma^{i},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(s)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(s)\right\},$ (S12)
$p(\gamma^{j}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $C\,s^{\alpha-1}A_{\alpha}(\gamma^{j},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(0)^{\top}\Sigma_{\gamma^{j},\tau}\check{y}(0)\right\},$ (S13)
$p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)=\;$ $C\,A_{\alpha}(\gamma^{ij},\tau)\int_{0}^{s}u^{\alpha-1}(s-u)^{\alpha-1}\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(u)\right\}\mathrm{d}u,$ (S14)
where $C>0$ is a constant and
$A_{\alpha}(\gamma,\tau)=\frac{\Gamma(|\gamma|\alpha)}{\Gamma(\alpha)^{|\gamma|}}\,\tau^{-|\gamma|/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\,\theta^{|\gamma|}(1-\theta)^{N-|\gamma|}.$ (S15)
In particular, when $\alpha=1$,
$p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)$ $=\frac{CA_{\alpha}(\gamma^{ij},\tau)\sqrt{2\pi}}{\sqrt{\phi\Lambda_{i,j}}}\exp\left\{-\frac{\phi}{2}\left\{\check{y}(0)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(0)-\beta_{i,j}^{2}\Lambda_{i,j}\right\}\right\}$ (S16)
$\quad\times\left\{\Phi\left((s-\beta_{i,j})\sqrt{\phi\Lambda_{i,j}}\right)-\Phi\left(-\beta_{i,j}\sqrt{\phi\Lambda_{i,j}}\right)\right\},$ (S17)
where $\beta_{i,j},\Lambda_{i,j}$ are as given in Lemma S1 and $\Phi$ denotes the CDF of $N(0,1)$.
Proof.
We can express the conditional posterior of $(\gamma_{i},\gamma_{j},\mu_{i},\mu_{j})$ as
$\quad p(\gamma_{i},\gamma_{j},\mu_{i},\mu_{j}\mid y,\phi,\tau,\mu_{-(i,j)})$ (S18)
$=C^{\prime}p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})\,p(\gamma_{i},\gamma_{j},\mu_{i},\mu_{j}\mid\mu_{-(i,j)})\,p(\mu_{-(i,j)})$ (S19)
$=C^{\prime}p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})\,p(\gamma,\mu),$ (S20)
$=C^{\prime}p(y\mid\phi,\tau,\mu_{i},\mu_{j},\mu_{-(i,j)})\,p(\mu\mid\gamma)\,p(\gamma),$ (S21)
where $C^{\prime}$ denotes some unknown constant that does not depend on $(\mu_{i},\mu_{j})$, and in the second step we have used that $\gamma$ is determined by $\mu$. Note that $p(\gamma,\mu)$ can only be nonzero when $\gamma\in\{\gamma^{i},\gamma^{j},\gamma^{ij}\}$. In particular, $\gamma^{0}$ has zero conditional prior probability given $\mu_{-(i,j)}$ since it would violate the simplex constraint. By (6), $p(\gamma)=\theta^{|\gamma|}(1-\theta)^{N-|\gamma|}$ and
$p(\mu\mid\gamma)=\frac{\Gamma(|\gamma|\alpha)}{\Gamma(\alpha)^{|\gamma|}}\prod_{k\in\gamma}\mu_{k}^{\alpha-1}.$
Given $\mu_{-(i,j)}$ and $(\gamma_{i},\gamma_{j})=(1,0)$, we have $(\mu_{i},\mu_{j})=(s,0)$ with probability one. Hence, using the likelihood given by (8) and the expression for $A_{\alpha}(\gamma,\tau)$, we get
$\quad p(\gamma_{i}=1,\gamma_{j}=0\mid y,\phi,\tau,\mu_{-(i,j)})$
$=C^{\prime}p(y\mid\phi,\tau,\mu_{i}=s,\mu_{j}=0,\mu_{-(i,j)})\,p(\mu_{i}=s,\mu_{j}=0,\mu_{-(i,j)}\mid\gamma^{i})\,p(\gamma^{i})$
$=Cs^{\alpha-1}A_{\alpha}(\gamma^{i},\tau)\exp\left\{-\frac{\phi}{2}\check{y}(s)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(s)\right\},$
$\text{ where }C=C^{\prime}\phi^{M/2}\prod_{k\in\gamma\setminus\{i,j\}}\mu_{k}^{\alpha-1}.$
This proves the claimed expression for $p(\gamma^{i}\mid y,\phi,\tau,\mu_{-(i,j)})$. An analogous calculation yields the claimed expression for $p(\gamma^{j}\mid y,\phi,\tau,\mu_{-(i,j)})$.
To find the conditional posterior probability of $\gamma^{ij}$, we use the conditional distribution of $\mu_{i}$ given $\gamma_{i}=\gamma_{j}=1$ in Lemma S1 to get
$\quad p(\gamma_{i}=1,\gamma_{j}=1\mid y,\phi,\tau,\mu_{-(i,j)})$ (S22)
$=\int_{0}^{s}p(\gamma_{i}=1,\gamma_{j}=1,\mu_{i}=u,\mu_{j}=s-u\mid y,\phi,\tau,\mu_{-(i,j)})\mathrm{d}u$ (S23)
$=Cp(\gamma^{ij})\int_{0}^{s}p(y\mid\phi,\tau,\mu_{i}=s,\mu_{j}=s-u,\mu_{-(i,j)})\,p(\mu_{i}=u,\mu_{j}=s-u,\mu_{-(i,j)}\mid\gamma^{ij})\mathrm{d}u$ (S24)
$=C^{\prime}A_{\alpha}(\gamma^{ij},\tau)\int_{0}^{s}\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(u)\right\}u^{\alpha-1}(s-u)^{\alpha-1}\mathrm{d}u,$ (S25)
which proves the claimed expression for $p(\gamma^{ij}\mid y,\phi,\tau,\mu_{-(i,j)})$.
When $\alpha=1$, a routine calculation shows that the integral in (S25) can be computed as
$\quad\int_{0}^{s}\exp\left\{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{i},\tau}\check{y}(u)\right\}\mathrm{d}u$
$=\exp\left\{-\frac{\phi}{2}\left(\check{y}(0)^{\top}\Sigma_{\gamma^{i,j},\tau}\check{y}(0)^{\top}-\beta_{i,j}^{2}\Lambda_{i,j}\right)\right\}\frac{\sqrt{2\pi}}{\sqrt{\phi\Lambda_{i,j}}}\int_{0}^{s}\frac{\sqrt{\phi\Lambda_{i,j}}}{\sqrt{2\pi}}\exp\left\{-\frac{\phi\Lambda_{i,j}}{2}\left(u-\beta_{i,j}\right)^{2}\right\}\mathrm{d}u,$
where the integral on the right-hand side involves the Gaussian density with mean $\beta_{i,j}$ and variance $(\phi\Lambda_{i,j})^{-1}$. Hence,
$\int_{0}^{s}\frac{\sqrt{\phi\Lambda_{i,j}}}{\sqrt{2\pi}}\exp\left\{-\frac{\phi\Lambda_{i,j}}{2}\left(u-\beta_{i,j}\right)^{2}\right\}\mathrm{d}u=\;\Phi\left((s-\beta_{i,j})\sqrt{\phi\Lambda_{i,j}}\right)-\Phi\left(-\beta_{i,j}\sqrt{\phi\Lambda_{i,j}}\right),$ (S26)
proving the claim. ∎
Remark 1.
For any positive integer $\alpha$, the integral involved in $p(\gamma^{ij}\mid y,\mu_{-(i,j)},\tau,\phi)$ can always be expressed by using the PDF and CDF of the standard normal distribution, and thus the conditional posterior probabilities of $\gamma^{i},\gamma^{j},\gamma^{ij}$ can be computed very efficiently. For non-integer-valued $\alpha$, numerical integration is needed. The conditional posterior distribution of $(\mu_{i},\mu_{j})$ given $\gamma^{ij}$ is
$p(\mu_{i}=u,\mu_{j}=s-u\mid y,\mu_{-(i,j)},\tau,\phi,\gamma_{i}=\gamma_{j}=1)\,\propto\,u^{\alpha-1}(s-u)^{\alpha-1}e^{-\frac{\phi}{2}\check{y}(u)^{\top}\Sigma_{\gamma^{ij},\tau}\check{y}(u)},$
for $u\in(0,s)$. When $\alpha=1$, $u$ follows the truncated normal distribution, as shown in Lemma S1, and this allows for straightforward sampling of $(\mu_{i},\mu_{j})$. For other integer values of $\alpha$, one can use rejection sampling to generate samples of $u$, where the truncated normal distribution can be used as a reference distribution.
S1.3 Algorithm
The pseudocode for our proposed Metropolis-within-Gibbs sampler targeting $p(\mu,\tau,\phi\mid y)$ is given in Algorithm 1 in the main text. The most computationally intensive part of Algorithm 1 is the evaluation of expressions of the form $V_{\gamma,\tau}^{-1}z$ for a given vector $z$, which may seem highly time-consuming if $|\gamma|$ is large. But this can be implemented efficiently by computing and updating the Cholesky decomposition of $V_{\gamma,\tau}$, a technique commonly employed in Bayesian spike-and-slab variable selection (smith1996nonparametric, george1997approaches). More specifically, at the beginning of the $t$-th iteration, we compute the Cholesky decomposition of $V_{\gamma,\tau^{(t-1)}}$, where $\gamma$ is the model selected in the last iteration. Then, each time we draw $(\mu_{i},\mu_{j})$ from its full conditional posterior, we need to change one or two coordinates of $\gamma$, and the resulting Cholesky decomposition of $V_{\gamma,\tau^{(t-1)}}$ can be obtained by standard updating algorithms (golub2013matrix).
Certain variations of Algorithm 1 are also straightforward to implement. For example, one can replace the fixed-scan Gibbs updating of $\mu$ in Algorithm 1 by a random-scan update: in the $t$-th iteration, we uniformly sample a pair $i<j$ such that $\mu^{(t-1)}_{i}+\mu^{(t-1)}_{j}>0$ and update $(\mu^{(t)}_{i},\mu^{(t)}_{j})$ from its conditional posterior distribution (an acceptance-rejection step is needed to correct for the proposal bias), and then draw $\phi^{(t)}$ from the full conditional posterior and update $\tau^{(t)}$ by one Metropolis–Hastings step. In this case, the iterative complex factorization algorithm proposed by zhou2019fast can be used to efficiently solve the linear system $V_{\gamma,\tau}^{-1}z$.
Remark 2.
Consider the application of BVS-SS to SCM. Let $(\tilde{X},\tilde{Y}^{(1)})$ denote another data set with $\tilde{M}$ observations under the treatment. Since Algorithm 1 outputs samples $(\mu^{(t)},\tau^{(t)},\phi^{(t)},w^{(t)})_{t=1}^{n}$ from the joint posterior distribution, we can simply compute an ATT estimate for each sample by
$\tilde{Y}^{(0,t)}=\tilde{X}w^{(t)},\quad\widehat{\mathrm{ATT}}^{(t)}=\frac{1}{\tilde{M}}\sum_{i=1}^{\tilde{M}}(\tilde{Y}_{i}^{(1)}-\tilde{Y}_{i}^{(0,t)}).$
Then, $(\widehat{\mathrm{ATT}}^{(t)})_{t=1}^{n}$ approximates the posterior distribution of the ATT estimate, and their average yields the final ATT estimator. However, if the objective is to compute the posterior mean estimate for ATT, one can further reduce the variance of the estimator by integrating out $w^{(t)}$ as in (12) and compute the counterfactual $\tilde{Y}^{(0,t)}$ by
$\tilde{Y}^{(0,t)}=\mathbb{E}_{y}[\tilde{X}w\mid\gamma^{(t)},\mu^{(t)},\tau^{(t)},\phi^{(t)}]=\tilde{X}V_{\gamma^{(t)},\tau^{(t)}}^{-1}\left(X^{\top}_{\gamma^{(t)}}y+(1/\tau^{(t)})\mu^{(t)}_{\gamma^{(t)}}\right).$
S2 Constrained Least-squares Estimators
In this section, we review some known results about the least-squares estimator under a linear constraint (amemiya1985advanced). We present the results in the general form although for our analysis, we only need to consider the constraint $\mathsf{1}^{\top}\mu=1$, which is part of the simplex constraint.
Theorem S1.
Let $y\in\mathbb{R}^{n}$, $Z\in\mathbb{R}^{n\times p}$, $Q\in\mathbb{R}^{p\times k}$ for some $k<p$, and $v\in\mathbb{R}^{k}$. Assume that $Z^{\top}Z$ is invertible. Let $\mathcal{U}=\{\mu\in\mathbb{R}^{p}\colon Q^{\top}\mu=v\}$. Define the unconstrained least-squares estimator $\hat{\mu}$ and the linearly constrained least-squares estimator $\check{\mu}\in\mathcal{U}$ by
$\hat{\mu}=\operatorname*{argmin}_{\mu\in\mathbb{R}^{p}}\|y-Z\mu\|^{2}_{2},\quad\check{\mu}=\operatorname*{argmin}_{\mu\in\mathcal{U}}\|y-Z\mu\|^{2}_{2}.$
Let $R\in\mathbb{R}^{p\times(p-k)}$ be any matrix such that $R^{\top}Q=0$ and [QR][Q\;R] is invertible. Then, the following results hold.
  1. (i)
    $\check{\mu}$ can be expressed by
    $\check{\mu}$ $=\hat{\mu}-(Z^{\top}Z)^{-1}Q\left[Q^{\top}(Z^{\top}Z)^{-1}Q\right]^{-1}(Q^{\top}\hat{\mu}-v)$ (S27)
    $=R(R^{\top}Z^{\top}ZR)^{-1}R^{\top}Z^{\top}y+\left[I-R(R^{\top}Z^{\top}ZR)^{-1}R^{\top}Z^{\top}Z\right]Q(Q^{\top}Q)^{-1}v.$ (S28)
  2. (ii)
    For any $\mu\in\mathcal{U}$, we have $\|Z(\hat{\mu}-\mu)\|^{2}=\|Z(\hat{\mu}-\check{\mu})\|^{2}+\|Z(\check{\mu}-\mu)\|^{2}.$ Hence,
    $\|y-Z\mu\|^{2}_{2}=\|y-Z\hat{\mu}\|^{2}_{2}+\|Z(\hat{\mu}-\check{\mu})\|^{2}_{2}+\|Z(\check{\mu}-\mu)\|^{2}_{2}.$
  3. (iii)
    If $y=Z\mu^{*}+\epsilon$ for some $\mu^{*}\in\mathcal{U}$, then
    $\check{\mu}=\mu^{*}+R(R^{\top}Z^{\top}ZR)^{-1}R^{\top}Z^{\top}\epsilon.$ (S29)
Proof.
The claimed expressions for $\check{\mu}$ given in part (i) are proved in Chapters 1.4.1 and 1.4.2 of amemiya1985advanced. Part (iii) is proved in Chapter 1.4.3 of amemiya1985advanced. To prove part (ii), we use part (i) to get that, for any $w$ such that $Q^{\top}w=0$,
$w^{\top}Z^{\top}Z(\check{\mu}-\hat{\mu})=-w^{\top}Q\left[Q^{\top}(Z^{\top}Z)^{-1}Q\right]^{-1}(Q^{\top}\hat{\mu}-v)=0.$
Since $Q^{\top}(\check{\mu}-\mu)=v-v=0$ for any $\mu\in\mathcal{U}$, this implies that $\|Z(\hat{\mu}-\mu)\|^{2}=\|Z(\hat{\mu}-\check{\mu})\|^{2}+\|Z(\check{\mu}-\mu)\|^{2}$. The claimed decomposition for $\|y-Z\mu\|^{2}_{2}$ then follows from that $Z\hat{\mu}$ is the orthogonal projection of $y$ onto the column space of $Z$. ∎
Next, we prove some linear algebra results that will be useful for analyzing the constrained least-squares estimator.
Lemma S3.
Let $Q\in\mathbb{R}^{p\times k}$ and $R\in\mathbb{R}^{p\times(p-k)}$ be such that $R^{\top}Q=0$ and [QR][Q\;R] is invertible. For any positive definite matrix $A\in\mathbb{R}^{p\times p}$,
$Q(Q^{\top}A^{-1}Q)^{-1}Q^{\top}$ $=A-AR(R^{\top}AR)^{-1}R^{\top}A,$ (S30)
$\lambda_{\max}(R(R^{\top}AR)^{-1}R^{\top})$ $\leq\lambda_{\max}(A^{-1}).$ (S31)
Proof.
The first identity is well known in the statistical literature; see, e.g., smyth1996conditional. For completeness, we provide a proof here. Since $A$ is positive definite, it has a real-valued square root $A^{1/2}$. Consider the block matrix $B=[A^{-1/2}Q\quad A^{1/2}R]\in\mathbb{R}^{p\times p}$, where the two component matrices are orthogonal due to the assumption $R^{\top}Q=0$. Hence, the orthogonal projection on the column space of $B$ can be written as $P_{1}+P_{2}$ where
$P_{1}$ $=A^{-1/2}Q\left(Q^{\top}A^{-1}Q\right)^{-1}QA^{-1/2},$
$P_{2}$ $=A^{1/2}R\left(R^{\top}AR\right)^{-1}RA^{1/2}.$
Since $\mathrm{rank}(B)=\mathrm{rank}(A^{-1/2}Q)+\mathrm{rank}(A^{1/2}R)$ and [QR][Q\;R] is invertible, the matrix $B$ also has full rank. It follows that $P_{1}+P_{2}=I$, which yields (S30).
To prove (S31), it suffices to note that for any vector $z\in\mathbb{R}^{p}$,
$z^{\top}R(R^{\top}AR)^{-1}R^{\top}z=z^{\top}A^{-1/2}P_{2}A^{-1/2}z\leq\|A^{-1/2}z\|^{2}\leq\lambda_{\max}(A^{-1})\|z\|^{2},$
where the first equality follows from the fact that $P_{2}$ is an orthogonal projection matrix. ∎
S3 Proofs for Section 4
As in Section S1, we use $|\gamma|$ and $\ell$ interchangeably. All corresponding indices and symbols (e.g., star, tilde, etc.) are defined equivalently.
S3.1 Proof of Lemma 1
Proof.
It suffices to prove that under Assumption (A1), for any $\gamma,\gamma^{\prime}\in\mathbb{S}_{L}$ such that $\emptyset\neq\gamma\subseteq\gamma^{\prime}$, we have
$\frac{|\gamma^{\prime}|}{M\underline{\lambda}}\geq\mathsf{1}^{\top}(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})^{-1}\mathsf{1}\geq\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}\geq\frac{1}{M},$ (S32)
$(M\underline{\lambda})^{|\gamma^{\prime}|-|\gamma|}\leq\frac{\det(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})}{\det(X_{\gamma}^{\top}X_{\gamma})}\leq M^{|\gamma^{\prime}|-|\gamma|}.$ (S33)
Consider (S32) first. Let $S=X_{\gamma^{\prime}\setminus\gamma}^{\top}(I-H_{\gamma})X_{\gamma^{\prime}\setminus\gamma}$. Since $X_{\gamma^{\prime}}=[X_{\gamma}\;X_{\gamma^{\prime}\setminus\gamma}]$, applying the block matrix inversion formula shows that for any $z^{\top}=[z_{1}^{\top}\;z_{2}^{\top}]$,
$z^{\top}(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})^{-1}z$ $=z_{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}z_{1}+\tilde{z}^{\top}S^{-1}\tilde{z}\geq z_{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}z_{1},$
where $\tilde{z}=z_{2}-X_{\gamma^{\prime}\setminus\gamma}^{\top}X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}z_{1}$. In particular, $\mathsf{1}^{\top}(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})^{-1}\mathsf{1}\geq\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}.$ When $\gamma$ only contains one covariate, Assumption (A1) implies that $X_{\gamma}^{\top}X_{\gamma}=M$, which yields the asserted lower bound. For the asserted upper bound, by the min-max theorem for eigenvalues,
$\mathsf{1}^{\top}(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})^{-1}\mathsf{1}\leq\|\mathsf{1}\|_{2}^{2}\;\lambda_{\max}\left((X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})^{-1}\right)=\frac{|\gamma^{\prime}|}{\lambda_{\min}(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})}\leq\frac{|\gamma^{\prime}|}{M\underline{\lambda}},$ (S34)
where the last step follows from Assumption (A1).
To prove (S33), we first assume $|\gamma^{\prime}|-|\gamma|=1$ and let $\gamma^{\prime}\setminus\gamma=\{j\}$. Applying the block matrix inversion formula, we obtain that
$\det(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})=\det(X_{\gamma}^{\top}X_{\gamma})\det(S).$
Since $S=X_{j}^{\top}(I-H_{\gamma})X_{j}$ and $I-H_{\gamma}$ is a projection matrix, we have $S\leq M$ under Assumption (A1). Moreover, by Lemma S4 below, we have $S\geq M\underline{\lambda}$. Hence,
$M\underline{\lambda}\leq\frac{\det(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})}{\det(X_{\gamma}^{\top}X_{\gamma})}\leq M.$ (S35)
To conclude the proof, observe that for any $\gamma\subseteq\gamma^{\prime}$, $\det(X_{\gamma^{\prime}}^{\top}X_{\gamma^{\prime}})/\det(X_{\gamma}^{\top}X_{\gamma})$ can be written as the product of $(|\gamma^{\prime}|-|\gamma|)$ ratios, where the $k$-th ratio is $\det(X_{\gamma_{k+1}}^{\top}X_{\gamma_{k+1}})/\det(X_{\gamma_{k}}^{\top}X_{\gamma_{k}})$ for some $\gamma_{k},\gamma_{k+1}$ such that $\gamma_{k+1}\setminus\gamma_{k}=\{j\}$ for some $j$. ∎
Lemma S4.
Under Assumption (A1), for any $\gamma,\gamma^{\prime}\in\mathbb{S}_{L}$ such that $\gamma\subseteq\gamma^{\prime}$,
$\lambda_{\min}(X_{\gamma^{\prime}\setminus\gamma}^{\top}(I-H_{\gamma})X_{\gamma^{\prime}\setminus\gamma})\geq M\underline{\lambda}.$ (S36)
Proof.
See yang2016computational. ∎
S3.2 Proof of Theorem 1
We divide the proof into multiple parts. First, in Section S3.2.1, we construct the events that happen with probability at least $1-c_{1}N^{-c_{2}}$ under the true data-generating probability measure $\mathbb{P}^{*}$, where $c_{1},c_{2}>0$ are some universal constants. In Section S3.2.2, we derive an explicit expression for $p(\gamma,\phi\mid y)$ by using the property of the constrained least-squares estimators. In Section S3.2.3, we find a lower bound on $p(\gamma^{*},\phi\mid y)$ by analyzing both the unconstrained and constrained least-squares estimators under $\gamma^{*}$, In Section S3.2.4, we prove Theorem 1 by integrating over $\phi$. Throughout our proof, we use
$H_{\gamma}\coloneqq X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}$ (S37)
to denote the projection matrix onto the column space of $X_{\gamma}$.
S3.2.1 High-probability Events
We begin by constructing the events that are essential to our non-asymptotic analysis of the posterior distribution. They enable us to control the behavior of the error vector $\epsilon$. The proof is based on standard concentration inequalities for normal distributions.
Lemma S5.
Suppose Assumption (A1) holds, $M\geq\log N$, $N\geq 2$ and $L\geq 3$. Then,
$\mathbb{P}^{*}(E_{1}\cap E_{2}\cap E_{3}\cap E_{4})\geq 1-c_{1}N^{-c_{2}},$
for some universal constants $c_{1},c_{2}>0$, where
$E_{1}$ $=\left\{\max_{\begin{subarray}{c}(\gamma_{1},\gamma_{2})\in\mathbb{S}_{L}\times\mathbb{S}_{L}\\ \gamma_{2}\subset\gamma_{1}\end{subarray}}\frac{\epsilon^{\top}(H_{\gamma_{1}}-H_{\gamma_{2}})\epsilon}{\ell_{1}-\ell_{2}}\leq 3\sigma^{2}L\log N\right\},$ (S38)
$E_{2}$ $=\left\{\frac{1}{M}\max_{1\leq j\leq N}|\epsilon^{\top}X_{j}|\leq\frac{3\sigma\sqrt{\log N}}{\sqrt{M}}\right\},$ (S39)
$E_{3}$ $=\left\{\epsilon^{\top}H_{\gamma^{*}}\epsilon\leq\sigma^{2}\ell^{*}\log N\right\},$ (S40)
$E_{4}$ $=\left\{\frac{7}{8}\leq\frac{\epsilon^{\top}\epsilon}{\sigma^{2}M}\leq\frac{5}{4}\right\}.$ (S41)
Proof.
Given any $\gamma_{2}\subset\gamma_{1}$, we can write $H_{\gamma_{1}}-H_{\gamma_{2}}=\sum_{k}(H_{\gamma_{k}}-H_{\gamma_{k+1}})$ where $\gamma_{k}=\gamma_{k+1}\cup\{j\}$ for some $j\notin\gamma_{k+1}$. Hence, $E_{1}=\tilde{E}_{1}$, where
$\tilde{E}_{1}$ $=\left\{\max_{\begin{subarray}{c}\gamma\in\mathbb{S}_{L-1}\\ j\in[N],j\notin\gamma\end{subarray}}\epsilon^{\top}(H_{\gamma\cup\{j\}}-H_{\gamma})\epsilon\leq 3\sigma^{2}L\log N\right\}.$ (S42)
The rank of $H_{\gamma\cup\{j\}}-H_{\gamma}$ is 1, and thus $\sigma^{-2}\epsilon^{\top}(H_{\gamma\cup\{j\}}-H_{\gamma})\epsilon\sim\chi^{2}_{1}$. The concentration inequality for normal distribution yields
$\mathbb{P}^{*}\left\{\epsilon^{\top}(H_{\gamma\cup\{j\}}-H_{\gamma})\epsilon\geq 2\sigma^{2}t\right\}\leq 2e^{-t},\quad\forall\,t\geq 0.$ (S43)
Choose $t=(3/2)L\log N$. Applying the union bound and using $L\geq 3$, we get
$\mathbb{P}^{*}(E_{1})=\mathbb{P}^{*}(\tilde{E}_{1})$ $\geq 1-\sum_{\begin{subarray}{c}\gamma\in\mathbb{S}_{L-1}\\ j\in[N],j\notin\gamma\end{subarray}}\mathbb{P}^{*}\left\{\epsilon^{\top}(H_{\gamma\cup\{j\}}-H_{\gamma})\epsilon\geq 3\sigma^{2}L\log N\right\}$ (S44)
$\geq 1-2N^{L+1}\cdot N^{-\frac{3L}{2}}$ (S45)
$\geq 1-2N^{-\frac{1}{2}}.$ (S46)
Consider $E_{2}$. Define
$d(\epsilon)=\max_{1\leq j\leq N}|\epsilon^{\top}X_{j}|.$ (S47)
A routine argument using (A1) shows that $d(\epsilon)$ is Lipschitz continuous in $\epsilon$ with Lipschitz constant $\sqrt{M}$. Since $\epsilon\sim N(0,\sigma^{2}I_{M})$ under $\mathbb{P}^{*}$, by the concentration inequality for Lipschitz functions of Gaussian vectors, we have
$\mathbb{P}^{*}\left(|d(\epsilon)-\mathbb{E}^{*}[d(\epsilon)]|\geq t\sigma\sqrt{M}\right)\leq 2e^{-t^{2}/2}.$ (S48)
Since each $X_{j}^{\top}\epsilon\sim N(0,\sigma^{2}M)$, a standard result for the maximum of (possibly dependent) Gaussian random variables yields that
$\mathbb{E}^{*}\left[d(\epsilon)\right]\leq\sigma\sqrt{2M\log(2N)}\leq 2\sigma\sqrt{2M\log N},$ (S49)
whenever $N>1$. Letting $t=\sqrt{\log N}$ in (S48) and using (S49), we obtain that
$\mathbb{P}^{*}(E_{2})$ $=1-\mathbb{P}^{*}\left(d(\epsilon)>3\sigma\sqrt{M\log N}\right)$
$\geq 1-\mathbb{P}^{*}\left(|d(\epsilon)-\mathbb{E}^{*}[d(\epsilon)]|>\sigma\sqrt{M\log N}\right)$
$\geq 1-2N^{-1/2}$
An application of the union bound concludes the proof.
For $E_{3}$, the standard concentration inequality for chi-squared distributions yields that $\mathbb{P}^{*}(E_{3})\geq 1-c_{1}N^{-c_{2}\ell^{*}}$ for some universal constants $c_{1},c_{2}>0$ (laurent2000adaptive). Similarly, one can also verify that $\mathbb{P}^{*}(E_{4})\geq 1-c_{1}e^{-c_{2}M}$ for some $c_{1},c_{2}>0$. ∎
S3.2.2 Posterior Density of $(\gamma,\phi)$
The presence of the simplex constraint makes the technical analysis substantially different from that for unconstrained variable selection models. The main challenge is to bound the marginal likelihood
$p(y\mid\gamma,\phi)=\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\mathsf{1})}\left[p(y\mid\gamma,\mu_{\gamma},\phi)\right].$ (S50)
To address this, we begin by deriving an equivalent expression for $p(y\mid\gamma,\mu_{\gamma},\phi)$ using Theorem S1. For each $\gamma\in\mathbb{S}_{L}$, define the OLS estimator for the model $\gamma$ by
$\hat{\mu}_{\gamma}=(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}y.$ (S51)
Let $\check{\mu}_{\gamma}$ denote the least-squares estimator under the constraint $\mathsf{1}^{\top}\mu_{\gamma}=1$. By Theorem S1,
$\check{\mu}_{\gamma}=\hat{\mu}_{\gamma}+\frac{1-\mathsf{1}^{\top}\hat{\mu}_{\gamma}}{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}.$ (S52)
Further, for any $u\in\Delta^{|\gamma|-1}$, we have
$\big\|y-X_{\gamma}u\big\|^{2}_{2}=\big\|y-X_{\gamma}\hat{\mu}_{\gamma}\big\|_{2}^{2}+\big\|X_{\gamma}(u-\check{\mu}_{\gamma})\big\|_{2}^{2}+b_{\gamma},$ (S53)
where
$b_{\gamma}=\big\|X_{\gamma}(\hat{\mu}_{\gamma}-\check{\mu}_{\gamma})\big\|_{2}^{2}=\frac{(1-\mathsf{1}^{\top}\hat{\mu}_{\gamma})^{2}}{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}}.$ (S54)
Equation (S53) decomposes the residual sum of squares for any $u$ satisfying the simplex constraint into three parts. The first term, $\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}$, is the residual sum of squares for the least-squares estimator in the unconstrained setting. The second term, $\|X_{\gamma}(u-\check{\mu}_{\gamma})\|_{2}^{2}$, measures the deviation of $u$ from the constrained least-squares estimator $\check{\mu}_{\gamma}$. Note that $X_{\gamma}\check{\mu}_{\gamma}$ is the projection of $X_{\gamma}\hat{\mu}_{\gamma}$ onto the affine space $\{X_{\gamma}u\colon\mathsf{1}^{\top}u=1\}$; in particular, we always have $\mathsf{1}^{\top}\check{\mu}_{\gamma}=1$. The last quantity, $b_{\gamma}\geq 0$, measures the deviation of $\check{\mu}_{\gamma}$ from $\hat{\mu}_{\gamma}$, which captures how well the model satisfies the simplex constraint. When $\hat{\mu}_{\gamma}\in\Delta^{|\gamma|-1}$, we have $b_{\gamma}=0$.
For each $\gamma$, we also define two matrices $D_{\gamma}\in\mathbb{R}^{|\gamma|\times(|\gamma|-1)}$ and $\Pi_{\gamma}\in\mathbb{R}^{(|\gamma|-1)\times|\gamma|}$ by
$D_{\gamma}=\begin{bmatrix}I_{|\gamma|-1}\quad(-\mathsf{1})\end{bmatrix}^{\top},\quad\Pi_{\gamma}=\begin{bmatrix}I_{|\gamma|-1}&0\end{bmatrix}$. (S55)
Clearly $D_{\gamma}$ is a full-rank matrix whose column space is orthogonal to the one vector $\mathsf{1}$, and $\Pi_{\gamma}u$ yields the subvector of $u$ containing the first $|\gamma|-1$ entries. Since the dimension of $D_{\gamma}$ or $\Pi_{\gamma}$ is always clear from context, we will omit the subscript and simply write $D$ and $\Pi$.
In Proposition S1, we show that the decomposition given in (S53) enables us to compare the marginal likelihood $p(y\mid\gamma,\phi)$ with that in the unconstrained setting. The expectation with respect to $u\sim\text{Dir}(\mathsf{1})$ is converted to the probability mass (up to a normalizing constant) of a $(|\gamma|-1)$-dimensional Gaussian distribution centered at $\Pi\check{\mu}_{\gamma}$.
Proposition S1.
Let $D,\Pi$ beas defined in (S55). Define
$f(\ell)=\theta^{\ell}(1-\theta)^{N-\ell}(2\pi)^{(\ell-1)/2}.$ (S56)
For any $\gamma\in\mathbb{S}_{L}$ and $\phi>0$, the posterior density $p(\gamma,\phi\mid y)$ given in (19) can be expressed as
$p(\gamma,\phi\mid y)=Cf(|\gamma|)\,\phi^{(M+\kappa_{1})/2}\,e^{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}+\kappa_{2}\right)}\,\mathbb{P}\left(\tilde{U}_{\gamma,\phi}\in{\tilde{\Delta}^{|\gamma|-1}}\right),$ (S57)
where $C>0$ is the unknown normalizing constant, $\tilde{U}_{\gamma,\phi}\sim\mathcal{N}(\Pi\check{\mu}_{\gamma},\,(\phi D^{\top}X_{\gamma}^{\top}X_{\gamma}D)^{-1})$, and
$\tilde{\Delta}^{\ell}=\left\{\tilde{u}\in\mathbb{R}^{\ell}\colon\tilde{u}_{i}\geq 0\text{ for each $i$, and }\sum\nolimits_{i=1}^{\ell}\tilde{u}_{i}\leq 1\right\}.$ (S58)
Proof.
Observe that for any vector $w$ with $\mathsf{1}^{\top}w=0$, we have $w=D\Pi w$. Hence, for any $u\in\Delta^{|\gamma|-1}$,
$\|X_{\gamma}(u-\check{\mu}_{\gamma})\|_{2}^{2}=\|X_{\gamma}D\,\Pi(u-\check{\mu}_{\gamma})\|_{2}^{2}=\|X_{\gamma}D(\tilde{u}-\Pi\,\check{\mu}_{\gamma})\|_{2}^{2}.$ (S59)
where $\tilde{u}=\Pi u\in\tilde{\Delta}^{|\gamma|-1}$. By equations (S53) and (S59) and letting $u=\mu_{\gamma}$, we find that
$\quad p(y\mid\gamma,u,\phi)$ (S60)
$\propto\phi^{M/2}\,\exp\left\{-\frac{\phi}{2}\|y-X_{\gamma}u\|_{2}^{2}\right\}$ (S61)
$=\phi^{M/2}\,\exp\left\{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}+\|X_{\gamma}D(\tilde{u}-\Pi\check{\mu}_{\gamma})\|_{2}^{2}\right)\right\}$ (S62)
$=\frac{\phi^{(M-|\gamma|+1)/2}(2\pi)^{(|\gamma|-1)/2}}{\sqrt{\det(D^{\top}X_{\gamma}^{\top}X_{\gamma}D)}}e^{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}\right)}\,\mathcal{N}(\tilde{u};\,\Pi\check{\mu}_{\gamma},\,(\phi D^{\top}X_{\gamma}^{\top}X_{\gamma}D)^{-1}),$ (S63)
where $\tilde{u}=\Pi u$ and $\mathcal{N}(\cdot;m,V)$ denotes the density function of the normal distribution with mean $v$ and covariance matrix $V$. We claim that
$\det(D^{\top}X_{\gamma}^{\top}X_{\gamma}D)=\det(X_{\gamma}^{\top}X_{\gamma})\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}.$ (S64)
To prove this, define $\tilde{D}=[D\;\mathsf{1}]$. A routine calculation yields $\det(\tilde{D}\tilde{D}^{\top})=|\gamma|^{2}$. Hence,
[…]
$=\frac{1}{|\gamma|^{2}}\det(D^{\top}X_{\gamma}^{\top}X_{\gamma}D)$ $\det\left(\mathsf{1}^{\top}X_{\gamma}^{\top}X_{\gamma}\mathsf{1}-\mathsf{1}^{\top}X_{\gamma}^{\top}X_{\gamma}D\left[D^{\top}X_{\gamma}^{\top}X_{\gamma}D\right]^{-1}D^{\top}X_{\gamma}^{\top}X_{\gamma}\mathsf{1}\right),$ (S66)
where the second step follows from the standard block matrix determinant formula. By the identity (S30) in Lemma S3, the matrix involved in the second determinant term can be simplified to
$\mathsf{1}^{\top}X_{\gamma}^{\top}X_{\gamma}\mathsf{1}-\mathsf{1}^{\top}X_{\gamma}^{\top}X_{\gamma}D\left[D^{\top}X_{\gamma}^{\top}X_{\gamma}D\right]^{-1}D^{\top}X_{\gamma}^{\top}X_{\gamma}\mathsf{1}$ $=\frac{|\gamma|^{2}}{\mathsf{1}^{\top}(X_{\gamma}^{\top}X_{\gamma})^{-1}\mathsf{1}},$ (S67)
which yields (S64).
Using (S63), (S64) and the prior specification given in (15) and (16), we get
$p(\gamma,u,\phi\mid y)$ $=Cp(\gamma)p(\phi\mid\gamma)p(u\mid\gamma)p(y\mid\gamma,u,\phi)$ (S68)
$=Cf(|\gamma|)\,\phi^{(M+\kappa_{1})/2}\,e^{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+b_{\gamma}+\kappa_{2}\right)}\,$ (S69)
$\quad\times\mathcal{N}(\tilde{u};\,\Pi\check{\mu}_{\gamma},\,(\phi D^{\top}X_{\gamma}^{\top}X_{\gamma}D)^{-1}).$ (S70)
Since integrating in $u$ over $\Delta^{|\gamma|-1}$ is the same as integrating in $\tilde{u}=\Pi u$ over $\tilde{\Delta}^{|\gamma|-1}$, we get
$p(\gamma,\phi\mid y)=\int_{\tilde{\Delta}^{|\gamma|-1}}p(\gamma,u,\phi\mid y)\mathrm{d}\tilde{u}$,
which yields the claimed identity. ∎
S3.2.3 Lower Bound on the Posterior Density of $(\gamma^{*},\phi)$
Next, we consider the true model $\gamma^{*}$. By Proposition S1, to lower bound $p(\gamma^{*},\phi\mid y)$, we need to find an upper bound on $b_{\gamma}^{*}$ and a lower bound on $\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}})$ where $\tilde{U}_{\gamma^{*},\phi}\sim\mathcal{N}(\Pi\,\check{\mu}_{\gamma^{*}},(\phi D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1})$. Consider $b_{\gamma}^{*}$ first.
Lemma S6.
Suppose $\mathsf{1}^{\top}\mu^{*}_{\gamma^{*}}=1$. On the event $E_{3}$ defined in Lemma S5, $b_{\gamma^{*}}\leq\sigma^{2}\ell^{*}\log N$.
Proof.
For the true model $\gamma^{*}$, $\hat{\mu}_{\gamma^{*}}$ can be written as
$\hat{\mu}_{\gamma^{*}}=(X_{\gamma^{*}}^{\top}X_{\gamma^{*}})^{-1}X_{\gamma^{*}}^{\top}(X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}+\epsilon)=\mu^{*}_{\gamma^{*}}+(X_{\gamma^{*}}^{\top}X_{\gamma^{*}})^{-1}X_{\gamma^{*}}^{\top}\epsilon.$ (S71)
Using $\mathsf{1}^{\top}\mu^{*}_{\gamma^{*}}=1$ we find that
$1-\mathsf{1}^{\top}\hat{\mu}_{\gamma^{*}}=-\mathsf{1}^{\top}(X_{\gamma^{*}}^{\top}X_{\gamma^{*}})^{-1}X_{\gamma^{*}}^{\top}\epsilon.$ (S72)
By Theorem S1 and the definition of $b_{\gamma}$ given in (S54),
$b_{\gamma^{*}}$ $=\big\|X_{\gamma^{*}}(\hat{\mu}_{\gamma^{*}}-\check{\mu}_{\gamma^{*}})\big\|_{2}^{2}\leq\big\|X_{\gamma^{*}}(\hat{\mu}_{\gamma^{*}}-\mu^{*}_{\gamma^{*}})\big\|_{2}^{2}=\epsilon^{\top}H_{\gamma^{*}}\epsilon\leq\sigma^{2}\ell^{*}\log N,$ (S73)
on the event $E_{3}$. ∎
Bounding $\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}})$ is more difficult, and we divide the argument into two lemmas.
Lemma S7.
Suppose Assumptions (A1) and 20 hold. On the event $E_{2}$ defined in Lemma S5, $\check{\mu}_{\gamma^{*}}\in\Delta^{\ell^{*}-1}$, and the smallest element of $\check{\mu}_{\gamma^{*}}$, denoted by $\check{\mu}^{*}_{\min}$, satisfies
$\check{\mu}^{*}_{\min}\geq(c_{\mu}-3)\frac{\sigma\sqrt{L\log N}}{\underline{\lambda}\sqrt{M}}.$ (S74)
Proof.
Since $\mathsf{1}^{\top}\check{\mu}_{\gamma}=1$, we only need to prove the lower bound (S74). By part (iii) of Theorem S1,
$\check{\mu}_{\gamma^{*}}=\mu^{*}_{\gamma^{*}}+D(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1}D^{\top}X_{\gamma^{*}}^{\top}\epsilon.$ (S75)
Let $e_{j}$ denote the vector whose $j$-th entry is one and other entries are all zero. By the inequality (S31) given in Lemma S3, on the event $E_{2}$,
$\|\check{\mu}_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{\infty}$ $=\left\|D(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1}D^{\top}X_{\gamma^{*}}^{\top}\epsilon\right\|_{\infty}$ (S76)
$=\max_{j}\left|e_{j}^{\top}D(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1}D^{\top}X_{\gamma^{*}}^{\top}\epsilon\right|$ (S77)
$\leq\lambda_{\max}\left(D(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1}D^{\top}\right)\left(\max_{j}\|e_{j}\|_{2}\right)\|X_{\gamma^{*}}^{\top}\epsilon\|_{2}$ (S78)
$\leq\lambda_{\max}\left((X_{\gamma^{*}}^{\top}X_{\gamma^{*}})^{-1}\right)\,\sqrt{\ell^{*}}\|X_{\gamma^{*}}^{\top}\epsilon\|_{\infty}$ (S79)
$\leq\frac{3\sigma\sqrt{\ell^{*}\log N}}{\underline{\lambda}\sqrt{M}},$ (S80)
where the last step follows from Assumption (A1) and the definition of $E_{2}$. By Assumption 20,
$\|\check{\mu}_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{\infty}\leq\frac{3}{c_{\mu}}\min_{j\in\gamma^{*}}|\mu^{*}_{j}|,$ (S81)
Hence,
$\check{\mu}^{*}_{\min}\geq\min_{j\in\gamma^{*}}|\mu^{*}_{j}|-\|\check{\mu}_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{\infty}\geq\left(1-\frac{3}{c_{\mu}}\right)\min_{j\in\gamma^{*}}|\mu^{*}_{j}|,$ (S82)
from which the claimed bound follows. ∎
Lemma S8.
Suppose Assumptions (A1) and 20 hold with $c_{\mu}>6$. Let
$r=\frac{c_{\mu}-3}{\sqrt{\ell^{*}-1}}\frac{\sigma\sqrt{L\log N}}{\underline{\lambda}\sqrt{M}}.$ (S83)
On the event $E_{2}$ defined in Lemma S5, we have
$\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})\coloneqq\{\tilde{u}\in\mathbb{R}^{\ell^{*}-1}\colon\|\tilde{u}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}\leq r\}\subset\tilde{\Delta}^{\ell^{*}-1}.$
Further, for $\tilde{U}_{\gamma^{*},\phi}\sim\mathcal{N}(\Pi\,\check{\mu}_{\gamma^{*}},(\phi D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1})$, we have
$\mathbb{P}\left(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{\ell^{*}-1}}\right)\geq\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}}))\geq 1-e^{1/2}\exp\left(-\frac{4\phi\,\sigma^{2}L\log N}{(\ell^{*})^{2}}\right).$ (S84)
Proof.
Since $\Pi\,\check{\mu}_{\gamma^{*}}$ is the subvector of $\check{\mu}_{\gamma^{*}}$, Lemma S7 implies that $\Pi\,\check{\mu}_{\gamma^{*}}\in\tilde{\Delta}^{\ell^{*}-1}$ when $c_{\mu}>3$. We first find the minimum distance from $\Pi\,\check{\mu}_{\gamma^{*}}$ to the hyperplanes of $\tilde{\Delta}^{\ell^{*}-1}$, which are determined by $\{\tilde{u}\colon\tilde{u}_{i}=0\}$ for $i=1,\dots,\ell^{*}-1$ and $\{\tilde{u}\colon\sum_{i=1}^{\ell^{*}-1}\tilde{u}_{i}=1\}$. The distance from $\Pi\,\check{\mu}_{\gamma^{*}}$ to $\{\tilde{u}:\tilde{u}_{i}=0\}$ equals the $i$-th entry of $\Pi\,\check{\mu}_{\gamma^{*}}$, and the distance to $\{\tilde{u}\colon\sum_{i=1}^{\ell^{*}-1}\tilde{u}_{i}=1\}$ is $(1-\mathsf{1}^{\top}\Pi\,\check{\mu}_{\gamma^{*}})/\sqrt{\ell^{*}-1}$. Since $\mathsf{1}^{\top}\check{\mu}_{\gamma^{*}}=1$, $1-\mathsf{1}^{\top}\Pi\,\check{\mu}_{\gamma^{*}}$ equals the last entry of $\check{\mu}_{\gamma^{*}}$, which proves that the ball $\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})$ is contained in the simplex $\tilde{\Delta}^{\ell^{*}-1}$ whenever $r\leq\check{\mu}^{*}_{\mathrm{min}}/\sqrt{\ell^{*}-1}$. By Lemma S7, this condition is satisfied for the given choice of $r$. Clearly, this implies that $\mathbb{P}(\tilde{U}\in{\tilde{\Delta}^{\ell^{*}-1}})\geq\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})).$
By a well-known result for Gaussian concentration (vershynin2018high, Theorem 5.1.4), for the random vector $\tilde{U}=\tilde{U}_{\gamma^{*},\phi}$, we have
$\mathbb{P}\left(\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}-\mathbb{E}\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}>t\right)$ $\leq\exp\left\{-\frac{\phi t^{2}}{2}\lambda_{\min}(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)\right\},$ (S85)
for every $t\geq 0$. A direct calculation using $D^{\top}D=I_{\ell^{*}}+\mathsf{1}\mathsf{1}^{\top}$ shows that $\lambda_{\max}(D^{\top}D)=\ell^{*}$ and $\lambda_{\min}(D^{\top}D)=1$. Hence, under Assumption (A1),
$\lambda_{\min}(D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)\geq\lambda_{\min}(X_{\gamma^{*}}^{\top}X_{\gamma^{*}})\lambda_{\min}(D^{\top}D)\geq M\underline{\lambda}.$ (S86)
The expectation term on the left-hand side of (S85) then can be bounded by
$\mathbb{E}\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}\leq\sqrt{\lambda_{\max}((\phi D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1})}\,\mathbb{E}\chi_{\ell^{*}-1}\leq\frac{\sqrt{\ell^{*}}}{\sqrt{\phi M\underline{\lambda}}},$ (S87)
where $\chi_{\ell^{*}-1}$ denotes a random variable following the $\chi$-distribution with $(\ell^{*}-1)$ degrees of freedom. Applying (S85) with (S86) and (S87) yields
$\mathbb{P}\left(\tilde{U}\in\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})\right)$ $=1-\mathbb{P}\left(\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}-\mathbb{E}\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}>r-\mathbb{E}\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}\right)$ (S88)
$\geq 1-\mathbb{P}\left(\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}-\mathbb{E}\|\tilde{U}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}>r-\frac{\sqrt{\ell^{*}}}{\sqrt{\phi M\underline{\lambda}}}\right)$ (S89)
$\geq 1-\exp\left\{-\frac{\phi M\underline{\lambda}}{2}\left(r-\frac{\sqrt{\ell^{*}}}{\sqrt{\phi M\underline{\lambda}}}\right)^{2}\right\}.$ (S90)
Using $2ab\leq ca^{2}+c^{-1}b^{2}$ with $a=r$, $b=\sqrt{\ell^{*}}/\sqrt{\phi M\underline{\lambda}}$ and $c=\ell^{*}/(\ell^{*}+1)$, we get
$\left(r-\frac{\sqrt{\ell^{*}}}{\sqrt{\phi M\underline{\lambda}}}\right)^{2}$ $\geq r^{2}+\frac{\ell^{*}}{\phi M\underline{\lambda}}-\frac{\ell^{*}}{\ell^{*}+1}r^{2}-\frac{\ell^{*}+1}{\ell^{*}}\frac{\ell^{*}}{\phi M\underline{\lambda}}$
$=\frac{r^{2}}{\ell^{*}+1}-\frac{1}{\phi M\underline{\lambda}}$
$=\frac{(\check{\mu}^{*}_{\mathrm{min}})^{2}}{(\ell^{*}-1)(\ell^{*}+1)}-\frac{1}{\phi M\underline{\lambda}}$
$\geq(c_{\mu}-3)^{2}\frac{\sigma^{2}L\log N}{(\ell^{*})^{2}M\underline{\lambda}^{2}}-\frac{1}{\phi M\underline{\lambda}}$
$\geq\frac{9\sigma^{2}L\log N}{(\ell^{*})^{2}M\underline{\lambda}}-\frac{1}{\phi M\underline{\lambda}},$
where the last step follows from $c_{\mu}\geq 6$ and $\underline{\lambda}\leq 1$. Therefore,
$\mathbb{P}\left(\tilde{U}\in\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})\right)\geq 1-\exp\left\{-\frac{9\phi\sigma^{2}L\log N}{2(\ell^{*})^{2}}+\frac{1}{2}\right\}.$
The stated bound thus follows. ∎
S3.2.4 Proof of Theorem 1
Proof of Theorem 1.
This follows from Propositions S2 and S3 that we prove below, where we further bound $b_{\gamma^{*}}$ using Lemma S6. ∎
Proposition S2.
For any $\gamma\in\mathbb{S}_{L}$ and $\phi>0$,
$p(\gamma,\phi\mid y)\leq Cf(|\gamma|)\,\Gamma\left(\frac{M+\kappa_{1}}{2}\right)\left\{\frac{\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+\kappa_{2}}{2}\right\}^{-(\kappa_{1}+M)/2},$ (S91)
where the constant $C$ and function $f$ are as given in Proposition S1.
Proof.
Since $\mathbb{P}(\tilde{U}_{\gamma,\phi}\in{\tilde{\Delta}^{|\gamma|-1}})\in[0,1]$ and $b_{\gamma}\geq 0$, by Proposition S1,
$p(\gamma,\phi\mid y)\leq Cf(|\gamma|)\phi^{(M+\kappa_{1})/2}e^{-\frac{\phi}{2}\left(\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}+\kappa_{2}\right)}.$
Integrating over $\phi$ yields the claimed bound. ∎
Proposition S3.
Suppose Assumptions (A1) and (A5) hold with $c_{\mu}>6$ and $c_{M}\geq 4$. On the event $E_{2}\cap E_{3}\cap E_{4}$ defined in Lemma S5, we have
$p(\gamma^{*}\mid y)\geq\frac{C}{3}f(\ell^{*})\Gamma\left(\frac{M+\kappa_{1}}{2}\right)\left\{\frac{\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}+b_{\gamma^{*}}+\kappa_{2}}{2}\right\}^{-(M+\kappa_{1})/2},$ (S92)
where the constant $C$ and function $f$ are as given in Proposition S1.
Proof.
Define
$K=\frac{1}{2}\left(\kappa_{2}+\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}+b_{\gamma^{*}}\right),\quad\tilde{K}=\frac{4\sigma^{2}L\log N}{(\ell^{*})^{2}}.$ (S93)
By Proposition S1 and Lemma S8, we have
$p(\gamma^{*},\phi\mid y)\geq Cf(\ell^{*})\phi^{(M+\kappa_{1})/2}e^{-K\phi}\left(1-e^{1/2}e^{-\tilde{K}\phi}\right).$
Integrating over $\phi$ yields
$p(\gamma^{*}\mid y)$ $\geq Cf(\ell^{*})\int\phi^{(M+\kappa_{1})/2}e^{-K\phi}\left(1-e^{1/2}e^{-\tilde{K}\phi}\right)\mathrm{d}\phi$ (S94)
$=Cf(\ell^{*})\left[\int\phi^{(M+\kappa_{1})/2}e^{-K\phi}\mathrm{d}\phi-e^{1/2}\int\phi^{(M+\kappa_{1})/2}e^{-(K+\tilde{K})\phi}\mathrm{d}\phi\right]$ (S95)
$=Cf(\ell^{*})\Gamma\left(\frac{M+\kappa_{1}}{2}\right)\left[K^{-(M+\kappa_{1})/2}-e^{1/2}(K+\tilde{K})^{-(M+\kappa_{1})/2}\right].$ (S96)
It remains to show that the term involving $K+\tilde{K}$ in (S96) is negligible. First, we bound $K$ and $\tilde{K}$. On the event $E_{4}$,
$\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}=\epsilon^{\top}(I-H_{\gamma^{*}})\epsilon\leq\epsilon^{\top}\epsilon\leq\frac{5}{4}\sigma^{2}M.$ (S97)
Using Assumption (A3), Lemma S6 and $M\geq 4L\log N$, we get
$2K=\kappa_{2}+\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}+b_{\gamma^{*}}\leq\frac{1}{2}\sigma^{2}M+\frac{5}{4}\sigma^{2}M+\sigma^{2}\ell^{*}\log N\leq 2\sigma^{2}M.$ (S98)
Meanwhile, the conditions $M\geq 4L\log N$ and $\ell^{*}\leq\sqrt{L\log N}$ imply that
$4\sigma^{2}\leq\tilde{K}\leq\sigma^{2}M.$ (S99)
Hence, we can use the inequality $\log(1+x)\geq x/(1+x)$ for x>1x>-1 to get
[…]
Plugging this bound into (S96) yields
$p(\gamma^{*}\mid y)$ $\geq Cf(\ell^{*})\Gamma\left(\frac{M+\kappa_{1}}{2}\right)K^{-(M+\kappa_{1})/2}\left\{1-e^{1/2}\left(\frac{K+\tilde{K}}{K}\right)^{-(M+\kappa_{1})/2}\right\}$ (S102)
$\geq(1-e^{-1/2})Cf(\ell^{*})\Gamma\left(\frac{M+\kappa_{1}}{2}\right)K^{-(M+\kappa_{1})/2}.$ (S103)
The claimed bound then follows by using $1-e^{-1/2}\geq 1/3$. ∎
S3.3 Proof of Theorem 2
To prove Theorem 2, we first use Theorem 1 to identify the rate at which $p(\gamma\mid y)/p(\gamma^{*}\mid y)$ goes to zero for any single $\gamma\neq\gamma^{*}$. We consider two different cases separately: $\gamma^{*}\subset\gamma$ (overfitted) and $\gamma^{*}\not\subset\gamma$ (underfitted).
S3.3.1 Bounds for Overfitted Models
Proposition S4.
Suppose Assumptions (A1) to (A5) hold and $c_{\theta}$ is sufficiently large. On the event $E_{1}\cap E_{2}\cap E_{3}\cap E_{4}$ defined in Lemma S5,
$\sup_{\gamma\in\mathbb{S}_{L}\colon\gamma^{*}\subsetneq\gamma}\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq 3N^{-|\gamma\setminus\gamma^{*}|L}.$ (S104)
Proof.
Fix an overfitted model $\gamma\supsetneq\gamma^{*}$ and let $\ell=|\gamma|>\ell^{*}$. Applying the inequality $(1+x)^{n}\leq e^{nx}$ for $n>0$ and using
$\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}-\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}=y^{\top}(H_{\gamma^{*}}-H_{\gamma})y,$ (S105)
we obtain from Theorem 1 that
$\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq 3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\ell^{*}}\exp\left\{\frac{\kappa_{1}+M}{2}\frac{y^{\top}(H_{\gamma}-H_{\gamma^{*}})y+\sigma^{2}\ell^{*}\log N}{\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}}\right\}.$ (S106)
Since $\gamma^{*}\subset\gamma$, we have $H_{\gamma^{*}}X_{\gamma^{*}}=H_{\gamma}X_{\gamma^{*}}=X_{\gamma^{*}}$, which implies that, on the event $E_{1}$,
$y^{\top}(H_{\gamma}-H_{\gamma^{*}})y=\epsilon^{\top}(H_{\gamma}-H_{\gamma^{*}})\epsilon\leq 3(\ell-\ell^{*})\sigma^{2}L\log N.$ (S107)
This further yields that $\sigma^{2}\ell^{*}\log N+y^{\top}(H_{\gamma}-H_{\gamma^{*}})y\leq 4(\ell-\ell^{*})\sigma^{2}L\log N$, as $\ell-\ell^{*}\geq 1$ and $L\geq\ell$. Similarly, on the event $E_{1}\cap E_{4}$, we can use $M\geq 4L\log N$ to get
$\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}=\epsilon^{\top}(I-H_{\gamma})\epsilon\geq\frac{7}{8}\sigma^{2}M-3\sigma^{2}L\log N\geq\frac{1}{8}\sigma^{2}M.$ (S108)
It thus follows from Theorem 1 and Assumptions (A2) and (A3) that
$\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}$ $\leq 3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\ell^{*}}\exp\left\{\frac{\kappa_{1}+M}{2}\frac{y^{\top}(H_{\gamma}-H_{\gamma^{*}})y+\sigma^{2}\ell^{*}\log N}{\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}}\right\}$
$\leq 3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\ell^{*}}\exp\left\{\frac{4(\ell-\ell^{*})\sigma^{2}L\log N}{(\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2})/M}\right\}$
$\leq 3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\ell^{*}}\exp\left\{32(\ell-\ell^{*})L\log N\right\}$
$=3(2\pi)^{(\ell-\ell^{*})/2}\exp\left\{(32-c_{\theta})(\ell-\ell^{*})L\log N\right\}.$
Letting $c_{\theta}$ be sufficiently large, we obtain the claimed bound. ∎
S3.3.2 Bounds for Underfitted Models
Proposition S5.
Suppose Assumptions (A1) to (A5) hold for sufficiently large $c_{\theta},c_{\mu},c_{M}$. On the event $E_{1}\cap E_{2}\cap E_{3}\cap E_{4}$ defined in Lemma S5,
$\sup_{\gamma\in\mathbb{S}_{L}\colon\gamma^{*}\not\subset\gamma}\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq 3N^{-(|\gamma\setminus\gamma^{*}|+|\gamma^{*}\setminus\gamma|)L}.$ (S109)
Proof.
[…]
$\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq R_{\rm{over}}R_{\rm{under}},$ (S110)
where
$R_{\rm{over}}$ $=3\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\tilde{\ell}-\ell^{*}}\left(\frac{\kappa_{2}+\|y-X_{\gamma^{*}}\hat{\mu}_{\gamma^{*}}\|_{2}^{2}+\sigma^{2}\ell^{*}\log N}{\kappa_{2}+\|y-X_{\tilde{\gamma}}\hat{\mu}_{\tilde{\gamma}}\|_{2}^{2}}\right)^{(\kappa_{1}+M)/2},$
$R_{\rm{under}}$ $=\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\tilde{\ell}}\left(\frac{\kappa_{2}+\|y-X_{\tilde{\gamma}}\hat{\mu}_{\tilde{\gamma}}\|_{2}^{2}}{\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}}\right)^{(\kappa_{1}+M)/2}.$
Since $\gamma^{*}\subset\tilde{\gamma}$, Proposition S4 can be applied to bound $R_{\rm{over}}$ by $R_{\rm{over}}\leq 2N^{-L(\tilde{\ell}-\ell^{*})}$.
It only remains to bound $R_{\rm{under}}$. As in the overfitted case, applying the inequality $(1+x)^{n}\leq e^{nx}$ and using $\kappa_{1}\leq M$ in Assumption (A3), we get
$R_{\rm{under}}\leq\left(\frac{\theta\sqrt{2\pi}}{1-\theta}\right)^{\ell-\tilde{\ell}}\exp\left\{-\frac{y^{\top}(H_{\tilde{\gamma}}-H_{\gamma})y}{(\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2})/M}\right\}.$ (S111)
The difference in residual sum of squares can be bounded as
$y^{\top}(H_{\tilde{\gamma}}-H_{\gamma})y$ $=\|(H_{\tilde{\gamma}}-H_{\gamma})(X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}+\epsilon)\|_{2}^{2}$
$=\|(I-H_{\gamma})X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}+(H_{\tilde{\gamma}}-H_{\gamma})\epsilon\|_{2}^{2}$
$\geq\left(\|(I-H_{\gamma})X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}\|_{2}-\|(H_{\tilde{\gamma}}-H_{\gamma})\epsilon\|_{2}\right)^{2},$
where on the second line we have used $H_{\tilde{\gamma}}X_{\gamma^{*}}=X_{\gamma^{*}}$ since $\tilde{\gamma}$ is overfitted, and the last step follows from the reverse triangle inequality. On the event $E_{1}$, $\|(H_{\tilde{\gamma}}-H_{\gamma})\epsilon\|_{2}^{2}\leq 3(\tilde{\ell}-\ell)\sigma^{2}L\log N$. Under Assumptions (A1) and 20, we apply Lemma S4 to get
$\|(I-H_{\gamma})X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}\|_{2}^{2}$ $=\|(I-H_{\gamma})X_{\gamma^{*}\setminus\gamma}\mu^{*}_{\gamma^{*}\setminus\gamma}\|_{2}^{2}$ (S112)
$\geq M\underline{\lambda}\|\mu^{*}_{\gamma^{*}\setminus\gamma}\|_{2}^{2}$ (S113)
$\geq(\tilde{\ell}-\ell)c_{\mu}^{2}\sigma^{2}L\log N.$ (S114)
For sufficiently large $c_{\mu}$, we thus find that
$y^{\top}(H_{\tilde{\gamma}}-H_{\gamma})y\geq\frac{1}{2}\|(I-H_{\gamma})X_{\gamma^{*}}\mu^{*}_{\gamma^{*}}\|_{2}^{2}.$ (S115)
To bound the denominator in the exponent, we apply the triangle inequality to get
$\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2}\leq\frac{1}{2}\sigma^{2}M+2\epsilon^{\top}\epsilon+2\|(I-H_{\gamma})X_{\gamma^{*}}\mu_{\gamma^{*}}\|_{2}^{2}\leq 3\sigma^{2}M+2\|(I-H_{\gamma})X_{\gamma^{*}}\mu_{\gamma^{*}}\|_{2}^{2},$ (S116)
on the event $E_{4}$. Define $S=\|(I-H_{\gamma})X_{\gamma^{*}}\mu_{\gamma^{*}}\|_{2}^{2}$. Using the elementary inequality $(x_{1}+x_{2})^{-1}\geq(2x_{1})^{-1}\wedge(2x_{2})^{-1}$, we can bound the exponent in (S111) as
$\frac{y^{\top}(H_{\tilde{\gamma}}-H_{\gamma})y}{(\kappa_{2}+\|y-X_{\gamma}\hat{\mu}_{\gamma}\|_{2}^{2})/M}$ $\geq\frac{S/2}{3\sigma^{2}+2S/M}$
$\geq\frac{S}{12\sigma^{2}}\wedge\frac{M}{8}$
$\geq\left(\frac{c_{\mu}^{2}}{12}\wedge\frac{c_{M}}{8}\right)(\tilde{\ell}-\ell)L\log N,$
where in the last step we have used Assumptions 20 and (A5) and the fact $\tilde{\ell}-\ell\leq\ell^{*}$. It is now clear from (S111) that as long as $c_{\mu}$ and $c_{M}$ are sufficiently large relative to $c_{\theta}$, $R_{\rm{under}}\leq N^{-L(\tilde{\ell}-\ell)}$. Hence, we conclude that
$\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq R_{\rm{over}}R_{\rm{under}}\leq 3N^{-L(2\tilde{\ell}-\ell^{*}-\ell)}$.
This yields the claimed bound upon since $\tilde{\ell}-\ell^{*}=|\gamma\setminus\gamma^{*}|$ and $\tilde{\ell}-\ell=|\gamma^{*}\setminus\gamma|$. ∎
S3.4 Proof of Theorem 2
Proof.
In Appendix S3.3, we prove that under our assumptions, with high probability,
$\sup_{\gamma\in\mathbb{S}_{L}\setminus\{\gamma^{*}\}}\frac{p(\gamma\mid y)}{p(\gamma^{*}\mid y)}\leq 3N^{-(|\gamma\setminus\gamma^{*}|+|\gamma^{*}\setminus\gamma|)L}.$ (S117)
See Proposition S4 for the overfitted case (i.e., $\gamma^{*}\subset\gamma$) and Proposition S5 for the underfitted case (i.e., $\gamma^{*}\not\subset\gamma$). A routine calculation using $|\{\gamma\in\mathbb{S}_{L}\colon|\gamma\setminus\gamma^{*}|+|\gamma^{*}\setminus\gamma|=k\}|\leq N^{k}$ yields the result. ∎
S3.5 Proof of Theorem 3
Proof.
Consider $\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]$ first. For any $\mu$ satisfying the simplex constraint, it is easy to show that $\|\mu-\mu^{*}\|_{2}^{2}\leq 2$. Hence, for any $\gamma\in\mathbb{S}_{L}$, $\mathbb{E}_{y}\left[\|\mu_{\gamma}-\mu^{*}_{\gamma}\|_{2}^{2}\right]\leq 2$ since $p(\mu_{\gamma}\mid y,\gamma)$ has support $\Delta^{|\gamma|-1}$. Applying Theorem 2, we get
$\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]$ $=\sum_{\gamma\in\mathbb{S}_{L}}p(\gamma\mid y)\mathbb{E}_{y}\left[\|\mu_{\gamma}-\mu^{*}_{\gamma}\|_{2}^{2}\right]$ (S118)
$\leq p(\gamma^{*}\mid y)\mathbb{E}_{y}\left[\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\right]+2\sum_{\gamma\in\mathbb{S}_{L}\setminus\{\gamma^{*}\}}p(\gamma\mid y)$ (S119)
$\leq\mathbb{E}_{y}\left[\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\right]+2c_{3}N^{-c_{4}L},$ (S120)
where $c_{3},c_{4}>0$ are universal constants introduced in Theorem 2.
It remains to bound
$\mathbb{E}_{y}\left[\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\right]$ $=\int\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\;p(\mu_{\gamma^{*}},\phi\mid y,\gamma^{*})\mathrm{d}\mu_{\gamma^{*}}\mathrm{d}\phi$ (S121)
$=\frac{1}{p(\gamma^{*}\mid y)}\int\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\;p(\gamma^{*},\mu_{\gamma^{*}},\phi\mid y)\mathrm{d}\mu_{\gamma^{*}}\mathrm{d}\phi.$ (S122)
Since when $\ell^{*}=1$, $\mathbb{E}_{y}\left[\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\right]=0$ due to the simplex constraint, we assume $\ell^{*}\geq 2$ henceforth. We use the notation introduced in Proposition S1 and write $u=\mu_{\gamma}^{*}$, $\tilde{u}=\Pi u$. The integral in (S122) is understood as an integral over $(\tilde{u},\phi)\in\tilde{\Delta}^{\ell^{*}-1}\times[0,\infty)$. Let $\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})$ be as defined in Lemma S8, and define
$B_{1}=\{(\tilde{u},\phi)\colon\tilde{u}\in\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}}),\,\phi>0\},\quad B_{2}=\{(\tilde{u},\phi)\colon\tilde{u}\in\tilde{\Delta}^{\ell^{*}-1}\setminus\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}}),\,\phi>0\}.$
Since $\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})\subset\tilde{\Delta}^{\ell^{*}-1}$ by Lemma S8, we can split the integral in (S122) into integrals over $B_{1}$ and $B_{2}$. We claim that for some univeral constant $c_{5}>0$,
$\|u-\mu^{*}_{\gamma^{*}}\|\leq\frac{c_{5}\sigma\sqrt{L\log N}}{\underline{\lambda}\sqrt{M}}\text{ on }B_{1},\text{ and }\|u-\mu^{*}_{\gamma^{*}}\|\leq 2\text{ on }B_{2}.$ (S123)
To prove the claim for $B_{1}$, we note that by the definition of $\mathcal{B}_{r}(\Pi\,\check{\mu}_{\gamma^{*}})$, if $(\tilde{u},\phi)\in B_{1}$,
$\|u-\check{\mu}_{\gamma^{*}}\|_{2}^{2}$ $=\|\tilde{\mu}-\Pi\,\check{\mu}_{\gamma^{*}}\|_{2}^{2}+\|1-\mathsf{1}^{\top}\tilde{u}-(1-\mathsf{1}^{\top}\Pi\,\check{\mu}_{\gamma^{*}})\|^{2}$ (S124)
$\leq r^{2}+\|\mathsf{1}^{\top}(\tilde{u}-\Pi\,\check{\mu}_{\gamma^{*}})\|^{2}\leq\ell^{*}r^{2},$ (S125)
in which the last step follows from Cauchy-Schwarz inequality. Using the definition of $r$ given in Lemma S7 and the assumption $\ell^{*}\geq 2$, we find that
$\|u-\check{\mu}_{\gamma^{*}}\|_{2}\leq\sqrt{2}(c_{\mu}-3)\frac{\sigma\sqrt{L\log N}}{\underline{\lambda}\sqrt{M}},$ (S126)
Meanwhile, repeating the argument for (S80), we get
$\|\mu^{*}_{\gamma^{*}}-\check{\mu}_{\gamma^{*}}\|_{2}\leq\frac{3\sigma\sqrt{\ell^{*}\log N}}{\underline{\lambda}\sqrt{M}},$ (S127)
on the event $E_{2}$. Combining (S126) and (S127) proves (S123). Plugging the bounds in (S123) back into (S122), we get
$\mathbb{E}_{y}\left[\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|_{2}^{2}\right]\leq\frac{c_{5}^{2}\sigma^{2}L\log N}{\underline{\lambda}^{2}M}+\frac{2}{p(\gamma^{*}\mid y)}\int_{B_{2}}p(\gamma^{*},u,\phi\mid y)\mathrm{d}\tilde{u}\,\mathrm{d}\phi.$ (S128)
By (S70) and using the notation introduced in equation (S93), we have
$p(\gamma^{*},u,\phi\mid y)=Cf(\ell^{*})\,\phi^{(M+\kappa_{1})/2}\,e^{-K\phi}\,\mathcal{N}(\tilde{u};\,\Pi\check{\mu}_{\gamma^{*}},\,(\phi D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1}).$ (S129)
Applying Proposition S3, we get
$\frac{\int_{B_{2}}p(\gamma^{*},u,\phi\mid y)\mathrm{d}\tilde{u}\,\mathrm{d}\phi}{p(\gamma^{*}\mid y)}$ $\leq\frac{3\int_{B_{2}}\phi^{(M+\kappa_{1})/2}\,e^{-K\phi}\,\mathcal{N}(\tilde{u};\,\Pi\check{\mu}_{\gamma^{*}},\,(\phi D^{\top}X_{\gamma^{*}}^{\top}X_{\gamma^{*}}D)^{-1})\mathrm{d}u\,\mathrm{d}\phi}{\Gamma\left(\frac{M+\kappa_{1}}{2}\right)K^{-(M+\kappa_{1})/2}}$ (S130)
$=\frac{3\int\phi^{(M+\kappa_{1})/2}\,e^{-K\phi}[1-\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{|\gamma|-1}})]\mathrm{d}\phi}{\Gamma\left(\frac{M+\kappa_{1}}{2}\right)K^{-(M+\kappa_{1})/2}},$ (S131)
where $\tilde{U}_{\gamma^{*},\phi}\sim\mathcal{N}(\Pi\check{\mu}_{\gamma},\,(\phi D^{\top}X_{\gamma}^{\top}X_{\gamma}D)^{-1})$ is as defined in Proposition S1. Using Lemma S8 and the notation $\tilde{K}$ introduced in equation (S93),
$\int\phi^{(M+\kappa_{1})/2}\,e^{-K\phi}[1$- $\mathbb{P}(\tilde{U}_{\gamma^{*},\phi}\in{\tilde{\Delta}^{|\gamma|-1}})]\mathrm{d}\phi\leq e^{1/2}\int\phi^{(M+\kappa_{1})/2}\,e^{-(K+\tilde{K})\phi}\mathrm{d}\phi$
$=e^{1/2}\Gamma\left(\frac{M+\kappa_{1}}{2}\right)(K+\tilde{K})^{-(M+\kappa_{1})/2}.$
By (S101),
$\left(\frac{K+\tilde{K}}{K}\right)^{-(M+\kappa_{1})/2}\leq\exp\left(-\frac{\tilde{K}}{4\sigma^{2}}\right)=\exp\left(-\frac{L\log N}{(\ell^{*})^{2}}\right).$
Since $e^{1/2}\leq 2$, it follows from (S131) that
$\frac{\int_{B_{2}}p(\gamma^{*},u,\phi\mid y)\mathrm{d}\tilde{u}\,\mathrm{d}\phi}{p(\gamma^{*}\mid y)}\leq 6\exp\left(-\frac{L\log N}{(\ell^{*})^{2}}\right).$ (S132)
Combining (S120), (S128) and (S132) yields the claimed bound for $\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]$.
For the posterior expected prediction loss, observe that given any model $\gamma$ and $\mu_{\gamma}\in\Delta^{|\gamma|-1}$, we can write $\tilde{X}\mu=\tilde{X}_{\gamma^{*}}\mu_{\gamma^{*}}+\tilde{X}_{\gamma\setminus\gamma^{*}}\mu_{\gamma\setminus\gamma^{*}}$. Hence,
$\|\tilde{X}\mu-\tilde{X}\mu^{*}\|^{2}_{2}$ $=\|\tilde{X}_{\gamma^{*}}(\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}})+\tilde{X}_{\gamma\setminus\gamma^{*}}\mu_{\gamma\setminus\gamma^{*}}\|_{2}^{2}\leq 2\|\tilde{X}_{\gamma^{*}}(\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}})\|^{2}_{2}+2\|\tilde{X}_{\gamma\setminus\gamma^{*}}\mu_{\gamma\setminus\gamma^{*}}\|_{2}^{2}$
$\leq 2\tilde{M}\overline{\lambda}\|\mu_{\gamma^{*}}-\mu^{*}_{\gamma^{*}}\|^{2}_{2}+2\tilde{M}\overline{\lambda}\|\mu_{\gamma\setminus\gamma^{*}}\|^{2}_{2}=2\tilde{M}\overline{\lambda}\|\mu-\mu^{*}\|_{2}^{2}.$
Therefore, the claimed bound on $\mathbb{E}_{y}\left[\|\tilde{X}\mu-\tilde{X}\mu^{*}\|_{2}^{2}\right]$ immediately follows from the bound on $\mathbb{E}_{y}\left[\|\mu-\mu^{*}\|_{2}^{2}\right]$. ∎
S3.6 Proof of Theorem 4
Consider the model $Y=Xw+\epsilon$ with $\epsilon\sim N(0,\phi^{-1}I)$. We use $\tilde{p}(\cdot\mid y)$ to denote the posterior distribution under the following prior:
$\phi\sim\;$ $\mathrm{Gamma}(\kappa_{1}/2,\kappa_{2}/2),$ (S133)
$\gamma_{i}\overset{\mathrm{i.i.d.}}{\sim}\;$ $\mathrm{Bernoulli}(\theta),$ (S134)
$\tau\sim\;$ $\mathrm{Gamma}(a_{1},a_{2}),$ (S135)
$w_{\gamma}\mid\gamma,\tau,\phi\sim\;$ $N(0,\,(\tau/\phi)I),$ (S136)
$w_{i}\mid\gamma_{i}=0\sim\;$ $\delta_{0}.$ (S137)
This is the classical setup for the spike-and-slab Bayesian variable selection, except that we place an additional prior distribution on $\tau$ so that it aligns with our BVS-SS model.
Proof of Theorem 4.
The posterior probability of $\gamma$ given $\tau$ under the two models can be expressed as
$p(\gamma\mid y,\tau)=C_{\tau}\,p_{0}(\gamma\mid y,\tau),\quad\tilde{p}(\gamma\mid y,\tau)=\tilde{C}_{\tau}\,\tilde{p}_{0}(\gamma\mid y,\tau),$ (S138)
where $C_{\tau},\tilde{C}_{\tau}$ denote the normalizing constants, and $p_{0},\tilde{p}_{0}$ denote the un-normalized posterior distributions given by
$p_{0}(\gamma\mid y,\tau)$ $=p(\gamma)\int p(y\mid\gamma,\mu_{\gamma},\phi,\tau)p(\mu_{\gamma}\mid\gamma)p(\phi)\mathrm{d}\mu_{\gamma}\,\mathrm{d}\phi,$ (S139)
$\tilde{p}_{0}(\gamma\mid y,\tau)$ $=\tilde{p}(\gamma)\int\tilde{p}(y\mid\gamma,\phi,\tau)\tilde{p}(\phi)\mathrm{d}\phi,$ (S140)
Note that the two models share the same prior distribution on $(\gamma,\phi)$: $p(\gamma)=\tilde{p}(\gamma)$ and $p(\phi)=\tilde{p}(\phi)$. For the unconstrained spike-and-slab model, we have
$\tilde{p}(y\mid\gamma,\phi,\tau)\;$ $=\left(\frac{\phi}{2\pi}\right)^{M/2}\tau^{-\ell/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\exp\left\{-\frac{\phi}{2}y^{\top}\Sigma_{\gamma,\tau}y\right\},$ (S141)
where $\ell=|\gamma|$, $V_{\gamma,\tau}$ and $\Sigma_{\gamma,\tau}$ are as defined in (9). Hence,
$\tilde{p}_{0}(\gamma\mid y,\tau)=G(\gamma,\tau)\left(y^{\top}\Sigma_{\gamma,\tau}y+\kappa_{2}\right)^{-(M+\kappa_{1})/2},$ (S142)
where $G(\gamma,\tau)=p(\gamma)\frac{\Gamma((M+\kappa_{1})/2)\,(\kappa_{2}/2)^{\kappa_{1}/2}}{\Gamma(\kappa_{1}/2)\,(2\pi)^{M/2}}\,\tau^{-\ell/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\,2^{(M+\kappa_{1})/2}.$ (S143)
Recall our likelihood given by (8):
$p(y\mid\mu_{\gamma},\gamma,\phi,\tau)\;$ $=\left(\frac{\phi}{2\pi}\right)^{M/2}\tau^{-\ell/2}\mathrm{det}(V_{\gamma,\tau})^{-1/2}\exp\left\{-\frac{\phi}{2}(y-X_{\gamma}\mu_{\gamma})^{\top}\Sigma_{\gamma,\tau}(y-X_{\gamma}\mu_{\gamma})\right\}.$ (S144)
By integrating out $\mu_{\gamma}$ and $\phi$, we get
$p_{0}(\gamma\mid y,\tau)=G(\gamma,\tau)\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\alpha)}\left[\left\{(X_{\gamma}\mu_{\gamma}-y)^{\top}\Sigma_{\gamma,\tau}(X_{\gamma}\mu_{\gamma}-y)+\kappa_{2}\right\}^{-(M+\kappa_{1})/2}\right],$ (S145)
where $\mathrm{Dir}(\alpha)$ denotes the Dirichlet distribution with parameter vector $\alpha\mathsf{1}$ (i.e., symmetric Dirichlet distribution with concentration parameter $\alpha$).
Combining (S138), (S142) and (S145), we get
$\frac{p(\gamma\mid y,\tau)}{\tilde{p}(\gamma\mid y,\tau)}=\frac{C_{\tau}}{\tilde{C}_{\tau}}\frac{p_{0}(y\mid\gamma,\tau)}{\tilde{p}_{0}(y\mid\gamma,\tau)}=\frac{C_{\tau}}{\tilde{C}_{\tau}}\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\alpha)}\left[F(\mu_{\gamma};\gamma,\tau)\right],$ (S146)
where
$F(\mu_{\gamma};\gamma,\tau)=\left(\frac{(X_{\gamma}\mu_{\gamma}-y)^{\top}\Sigma_{\gamma,\tau}(X_{\gamma}\mu_{\gamma}-y)+\kappa_{2}}{y^{\top}\Sigma_{\gamma,\tau}y+\kappa_{2}}\right)^{-\frac{M+\kappa_{1}}{2}}\leq\left(\frac{\kappa_{2}}{y^{\top}\Sigma_{\gamma,\tau}y+\kappa_{2}}\right)^{-\frac{M+\kappa_{1}}{2}}.$ (S147)
The inequality follows from that $\Sigma_{\gamma,\tau}=(I+\tau X_{\gamma}X_{\gamma}^{\top})^{-1}$ is always non-negative definite. Hence, we can apply dominated convergence theorem to get
$\lim_{\tau\rightarrow\infty}\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\alpha)}\left[F(\mu_{\gamma};\gamma,\tau)\right]=\mathbb{E}_{\mu_{\gamma}\sim\mathrm{Dir}(\alpha)}\left[\lim_{\tau\rightarrow\infty}F(\mu_{\gamma};\gamma,\tau)\right]=1.$ (S148)
To see that $F(\mu_{\gamma};\gamma,\tau)$ converges to $1$ as $\tau\rightarrow\infty$, we note that
$\quad\lim_{\tau\rightarrow\infty}\lambda_{\max}\left(\Sigma_{\gamma,\tau}-(I-X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top})\right)$ (S149)
$=\lim_{\tau\rightarrow\infty}\lambda_{\max}\left(X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}-X_{\gamma}(X_{\gamma}^{\top}X_{\gamma}+\tau^{-1}I)^{-1}X_{\gamma}^{\top}\right)=0,$ (S150)
which can be proved by using the singular value decomposition (SVD) of $X_{\gamma}$. Assume the rank of $X_{\gamma}$ is $\ell<M$. Let the SVD of $X_{\gamma}$ be $X_{\gamma}=U[S\;0]^{\top}V^{\top}$, where $U\in\mathbb{R}^{M\times M}$, $V\in\mathbb{R}^{\ell\times\ell}$ are orthogonal, and $S\in\mathbb{R}^{\ell\times\ell}$ is diagonal. Denote the diagonal elements of $S$ (i.e, singular values) as $s_{i}$ for $i\in[\ell]$. A routine calculation shows that the eigenvalues of $X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}-X_{\gamma}(X_{\gamma}^{\top}X_{\gamma}+\tau^{-1}I)^{-1}X_{\gamma}^{\top}$ are $1/(\tau s_{i}^{2}+1)$ for $i\in[\ell]$, which converge to 0 as $\tau\rightarrow\infty$. When $X_{\gamma}$ does not have full rank, we only need to change the limit to $I-X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{+}X_{\gamma}^{\top}$, where $A^{+}$ denotes the Moore-Penrose pseudoinverse, and then apply the same argument. Hence, $\Sigma_{\gamma,\tau}$ converges to the projection matrix $I-X_{\gamma}(X_{\gamma}^{\top}X_{\gamma})^{-1}X_{\gamma}^{\top}$ in operator norm, which implies that $X_{\gamma}^{\top}\Sigma_{\gamma,\tau}\rightarrow 0$ in $L^{2}$-norm and thus $F(\mu_{\gamma};\gamma,\tau)\rightarrow 1$.
By (S148) and (S146), we can prove the claim provided that $\lim_{\tau\rightarrow\infty}C_{\tau}/\tilde{C}_{\tau}=1$. To this end, recall that $C_{\tau},\tilde{C}_{\tau}$ are normalizing constants and thus can be expressed as
$C_{\tau}^{-1}=\sum_{\gamma\in\mathbb{S}_{L}}p_{0}(\gamma\mid y,\tau),\quad\tilde{C}_{\tau}^{-1}=\sum_{\gamma\in\mathbb{S}_{L}}\tilde{p}_{0}(\gamma\mid y,\tau).$ (S151)
Since the set $\mathbb{S}_{L}$ is finite and we have shown that $p_{0}(y\mid\gamma,\tau)/\tilde{p}_{0}(y\mid\gamma,\tau)\rightarrow 1$ for each $\gamma$, we have $C_{\tau}/\tilde{C}_{\tau}\rightarrow 1$ as well. ∎
S4 Additional Simulation Studies
We present in the main text the simulation results for a sparse regression model, where each entry of $X$ is generated independently from a standard normal distribution. In this section, we provide additional simulations to evaluate the performance of BVS-SS in estimating the ATT, selecting relevant variables, and estimating variance components under two alternative data-generating processes.
First, in Section S4.1, we consider a non-sparse setting where the data $(Y,X)$ are generated from a factor model. Second, in Section S4.2, we examine a sparse model in which the covariates in $X$ are correlated. These simulation designs follow the factor-model setups commonly used in the literature (e.g., hsiao2012panel, shi2023forward).
S4.1 Non-sparse Factor Models
We first assume that the training data $X\in\mathbb{R}^{M\times N}$ and $Y\in\mathbb{R}^{M}$ are generated from a non-sparse factor model,
$(Y,\,X)=F^{\top}\Lambda^{\top}+e,$ (S152)
where $(Y,X)\in\mathbb{R}^{M\times(N+1)}$ denotes the block matrix obtained by concatenating $Y$ and $X$, $e$ is an $M\times(N+1)$ error matrix with each entry $e_{ij}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,0.5^{2})$, $F\in\mathbb{R}^{4\times M}$ is a factor matrix with 4 factors, and $\Lambda\in\mathbb{R}^{(N+1)\times 4}$ is the corresponding loading matrix. For the matrix $F$, we let the four common factors evolve according to
$f_{1,i}$ $\sim\mathcal{N}(0,1),$ (S153)
$f_{2,i}$ $=0.9f_{2,i-1}+u_{1,i},$ (S154)
$f_{3,i}$ $=0.5f_{3,i-1}+u_{2,i}+0.5u_{2,i-1},$ (S155)
$f_{4,i}$ $=u_{3,i}+0.8u_{3,i-1}+0.4u_{3,i-2},\qquad i=1,\dots,M,$ (S156)
in which $u_{1,i},u_{2,i},u_{3,i}$ are i.i.d. standard normal. For $\Lambda$, we specify each loading $\lambda_{j,l}$ as
$\lambda_{j,l}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\begin{cases}\mathrm{Unif}[1,2],&j=1,\dots,J+1,\\[4.26773pt] \mathrm{Unif}\!\left[-\frac{2}{M+\tilde{M}},-\frac{2}{M+\tilde{M}}\right],&j>J+1,\end{cases}$ (S157)
where $J$ is a parameter whose value will be specified in the simulation.
Then we simulate the data under treatment $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$ and $\tilde{Y}^{(1)}\in\mathbb{R}^{\tilde{M}}$ by
[…]
$\tilde{Y}^{(1)}$ $=\tilde{Y}^{(0)}+\delta,$ (S159)
where $\delta_{i}$ is defined as in Section 5.1, and $\tilde{e}\in\mathbb{R}^{\tilde{M}\times(N+1)}$ are defined similarly. $\tilde{F}\in\mathbb{R}^{4\times\tilde{M}}$ follows the same factor structure as $F$; that is, we generate (F,F~)($\tilde{F}$) by (S156) for $i=1,\dots,M+\tilde{M}$. All control units are correlated with the treated unit through the common factors. However, the dependence persists primarily for the first $J$ units, while the correlation for the remaining units diminish as the sample size grows.
[…]
Figure S1: Relative efficiency (%) against sample size for ATT estimation for factor models. Left: $N=20,J=5$; right: $N=50,J=10$.
Meanwhile, the posterior mean of $\phi$ does not exhibit a clear increasing trend. This is due to our simulation setting, in which the true DGP is a factor model while our Bayesian estimation is based on a sparse regression model. The trend of $\tau$ remains relatively stable, exhibiting only a very slight decrease as $M$ increases. Moreover, the magnitude of the posterior mean of $\tau$ is very close to that in the sparse regression model with $w^{*}||_{1}=2$ presented in Section 5. This further supports the notion that the factor model is likely to violate the simplex constraint.
Figure S2: Distribution of the posterior mean of $\tau$ and $\phi$ across $100$ replicates with $N=20,J=5$.
Figure S3: Distribution of the posterior mean of $\tau$ and $\phi$ across $100$ replicates with $N=50,J=10$.
S4.2 Sparse Factor Models
Next, we consider a sparse model in which $X\in\mathbb{R}^{M\times N}$ is generated from a factor model with factor matrix $F\in\mathbb{R}^{4\times M}$ and loading matrix $\Lambda\in\mathbb{R}^{N\times 4}$, and the untreated potential outcome for $Y\in\mathbb{R}^{M\times 1}$ is a linear combination of $X$,
$X$ $=F^{\top}\Lambda^{\top}+e,$ (S160)
$Y$ $=Xw^{*}+\epsilon.$ (S161)
The treatment and sparse coefficients $w^{*}$ are specified in the same way as in Section 5.1. The sparsity still remains through $w^{*}$, but all units in the control group are linked by some factors that do not directly affect the treated unit. Analogously, we generate the following post-treatment data,
$\tilde{X}$ $=\tilde{F}^{\top}\Lambda^{\top}+\tilde{e}$, (S162)
$\tilde{Y}^{(1)}$ $=\tilde{X}w^{*}+\delta+\tilde{\epsilon}$, (S163)
where $\tilde{Y}^{(1)}\in\mathbb{R}^{\tilde{M}\times 1}$, $\tilde{X}\in\mathbb{R}^{\tilde{M}\times N}$, $\tilde{F}\in\mathbb{R}^{4\times\tilde{M}}$. $\delta\in\mathbb{R}^{\tilde{M}}$ and $\tilde{\epsilon}\in\mathbb{R}^{\tilde{M}}$ are defined as in Section 5.1. The specifications of (F,F~)($\tilde{F}$) and $\Lambda$ are the same as in Section S4.1, except that $\Lambda$ contains $N$ rows rather than $N+1$.
[…]
Figure S4: Relative efficiency (%) against sample size for ATT estimation for $X$ with correlations. First row: $N=20,J=5$; second row: $N=50,J=10$; Top row: $\|w^{*}\|_{1}=1$; middle row: $\|w^{*}\|_{1}=2$; bottom row: $\|w^{*}\|_{1}=3$.
Regarding variable selection, BVS-SS generally attains higher accuracy than Lasso, particularly when the sample size $M$ is large. In smaller samples, however, its model selection performance deteriorates. For example, when $N=50,J=10,M\leq 50$, BVS-SS fails to select any true control unit on average, although Lasso performs nearly as poorly under these small-sample scenarios. This pattern is understandable: because our sampler is a Gibbs sampler that updates coefficients pairwise, strong correlations among covariates can lead to highly correlated variables being selected in place of the true signals. Importantly, despite the low accuracy of variable selection, BVS-SS delivers substantially better ATT estimation than Lasso when $M\leq 50$, as shown in Figure S4, highlighting its robustness even when exact model recovery is difficult. This observation also underscores the need for more advanced MCMC algorithms which might improve selection accuracy in such correlated designs.
BVS-$S$ Lasso
$\|w^{*}\|_{1}$ Setting $M$ $\ell^{1}$-loss model size $\ell^{1}$-loss model size
$\|w^{*}\|_{1}$ = 1
$N=20$
$J=5$
25 5.6 7.2 5.7 8.9
50 3.8 6.4 5.2 8.9
100 2.2 5.8 5.5 9.9
200 1.2 5.4 5.1 9.8
$N=50$
$J=10$
25 19.5 19.1 12.7 10.9
50 15.5 16.7 12.5 13.3
100 10 13.3 13.9 20.3
200 6.1 11.2 14.1 21.2
$\|w^{*}\|_{1}$ = 3
$N=20$
$J=5$
25 3.8 3.2 5.7 8.9
50 2.4 3.8 5.2 8.9
100 1.6 4.1 5.5 9.9
200 1 4.4 5.1 9.8
$N=50$
$J=10$
25 10.7 4.6 12.7 10.9
50 8.4 5.1 12.5 13.3
100 6 6.1 13.9 20.3
200 3.9 7 14.1 21.2
Table S1: Variable selection performance of BVS-SS and Lasso averaged over 100 replicates.
Finally, the posterior means of $\tau$ and $\phi$ exhibit patterns very similar to those reported in the main text, and we visualize their distributions in Figures S5 and S6.
Figure S5: Distribution of the posterior mean of $\phi$ across $100$ replicates with $N=50,J=10$. The true value $\phi^{*}$ is indicated by the dotted line.
Figure S6: Distribution of the posterior mean of $\tau$ across $100$ replicates with $N=50,J=10$.
S5 Additional Results for Empirical Examples
This section presents supplementary details for the empirical examples analyzed in the main text.
S5.1 Additional Results for NFP Anti-tax Evasion
We present more results for the NFP anti-tax evasion study in this section. Figure S7 shows the trajectories of the posterior samples of $\phi$ and $\log(\tau)$ (with burn-in included). Probably due to the limited number of observations ($T_{0}=33$), the MCMC samples of $\tau$ exhibit large variability, indicating that the simplex assumption is hard to confirm or reject, though the ratio of the posterior means of $\tau$ and $\phi$ given in Table 2 is relatively small. The trace plot suggests that our sampler has converged fast.
Figure S7: The trace plot of $\log(\tau)$ and $\phi$ for the NFP data set.
The posterior distributions of ATT, $\log(\tau)$, $\phi$ and model size $|\gamma|$ are visualized in the histograms in Figure S8. The posterior mean of the ATT exhibits a clear unimodal and bell-shaped distribution, concentrated in the interval $[0.26,0.34]$. The distribution of $\phi$ is symmetric with a mode between $[3.5,4.0]$, consistent with the posterior mean in Table 2, while the distribution of $\log(\tau)$ is right-skewed, indicating that a large portion of the posterior draws for $\tau$ are not concentrated near zero. This pattern is also reflected in the posterior mean of $\tau$ reported in Table 2. Taken together, the evidence suggests that the data in this empirical example do not favor the simplex constraint. The posterior distribution of $|\gamma|$ is left-skewed, with more than half of the posterior mass supporting a relatively small model that includes only two control units.
Figure S8: Posterior distributions for the Nota Fiscal Paulista data set.
Finally, Figure S9 plots the observed FHA inflation index together with the estimated counterfactual path. Before the intervention (marked by the vertical green line), the estimated pre-treatment counterfactual (blue dashed line) tracks the observed FHA index well, indicating a good in-sample fit. Following the intervention, the post-treatment counterfactual (red dashed line) diverges from the observed outcomes, capturing the estimated treatment effect over time.
Figure S9: Counterfactual estimation for the Nota Fiscal Paulista data set.
S5.2 Additional Results for China’s Anti-corruption Campaign
The section provides more details on the China’s anti-corruption example. From Figure S10, one can observe that the posterior distribution of $\tau$ is still dispersed, but that of $\phi$ is more concentrated and much larger on average, compared to the previous example. The estimation of $\phi$ appears to be more accurate, suggesting that the error variance is around 0.05 (1/20.86). The ratio $\tau/\phi$ becomes much smaller, indicating that the simplex constraint is more likely to be satisfied (at least approximately) in this data set.
Figure S10: The trace plot of $\log(\tau)$ and $\phi$ for the anti-corruption data set.
Figure S11 presents the posterior distributions. The ATT distribution is unimodal but a little right-skewed. The distribution of $\log(\tau)$ is in line with its trace plot, with a large probability mass for $\log(\tau)<-2$, but the distribution appears relatively diffuse since the bars in the histogram are fairly evenly distributed. More than 80% samples of $\phi$ concentrate in the interval of $[15,30]$, providing evidence that our BVS-SS has detected abundant signals in the data. BVS-SS still suggests a small model size for building the counterfactual, as around 70% MCMC samples have model size smaller than six.
Figure S11: Counterfactual Estimations for the China’s Anti-corruption Campaign data set.
Finally, we provide a counterfactual plot for the growth rate of the luxury watches import in Figure S12. Compared with the previous example, the in-sample fit is noticeably better, likely due to the availability of a richer set of control units. A pronounced initial decline in the import growth rate is detected in the post-intervention periods, which implies a substantial immediate policy impact. However, the effect attenuates over time, with the gap between the observed and counterfactual paths gradually narrowing. Overall, the figure illustrates that our BVS-SS reproduces the pre-treatment dynamics effectively and provides a coherent counterfactual trajectory to assess the policy impact.
Figure S12: Counterfactual estimation for the China’s Anti-corruption Campaign data set.